A Complete Bibliography of Publications in Annals of Probability (1990–1999)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 March 2019 Version 1.05

Title word cross-reference

(2, p) [Hou91]. {0, 1}N [BK95b]. 1 [AD98, BGM93, ST91, Tal90a]. 1/r^{2} [Fon93]. 2 [OR93, vdB97]. 2 + ϵ [HM98]. 3/2 [PWM90, PWM92]. α [AT92b, Elo90, Xia95]. B(3, n) [Sto95b]. d̄ [Mar96]. β [Nol88, Nol92]. C [Bun97]. d [BS97]. Δ_{2} [KN97]. d ≥ 3 [Jan98, Jan02]. D ⊆ R^{d} [EP99]. ϵ [Fel96]. G [DD99]. Γ [Ken95]. ≥ 3 [Tal94b]. H [Alb96]. H^{p} [Leu95]. I [Sch93b]. ∞ [AL99]. J [HP93]. K [Sep99, BGM93]. L [LPP95, MS92]. L^{2} [Deu94, RS97, dS92, IMM^{+}90]. L^{α} [Hou92]. L^{P} [CK91, GM95, CS93b]. L_{2} [Lig91]. l_{p} [Lin91a]. l_{p}, p ∈ [1, ∞[ [Fer92]. [−1, 1] [dlC90]. m [De 93, Wen90]. R^{2} [Roy90]. R^{d} [KR91, RHHQ90]. R^{n} [Bor93]. Z [Tot95]. Z^{1} [Tót96]. Z^{d} [Geo93, Lan91]. R^{d} [Tan93]. E [CKS98a]. n [Bár99, Voi97]. p [Dud92, MR92a, Qia98]. p = 1 [Jac91]. ϕ [Che91]. R^{2} [Duh98]. R^{3} [Duh98]. ρ [Sha95]. S^{3} × H^{3} [Hug92]. σ [DFST96, FT96, Van93]. t [GGM97]. U [AG93, BGZ94, BG99, BD94b, CdlPT93, FR93, GZ94, KN97, KW90, LdlP97, Maj94, OR93, Sch93a, Stu91, Stu94, Zha99, dlP92, dlPMSS94, dlPMS95].

X_{n+1} = a_{n}X_{n} + b_{n} (mod p) [Hil93]. X_{n} = A_{n}X_{n−1} + B_{n} [BBE97]. Y_{t} = B_{t} + α sup{Y_{s}: s ≤ t} + β inf{Y_{s}: s ≤ t} [Dav96]. Z [PV99]. Z^{2} [Che92, GG96]. Z^{d} [Lan93]. Z^{d}, d ≥ 2 [Che94a].

-Block [BGM93]. -Bounded [Hou91]. -Boundedness [GM95]. -close [Fel96]. -Congruence [Elo90]. -D [vdB97]. -decomposable [Bun97]. -Dependent [BGM93, De 93]. -Differentiability [dS92]. -dimensional [BS97]. -distance [Mar96]. -Exclusion [Sep99]. -Fields [Van93, DFST96, FT96]. -functions [KN97]. -Martingale [Ken95]. -martingales [CKS98a]. -Mixing [Che91, Sha95]. -Processes [AG93, Sch93a]. -Projection [Sch93b]. -projections [RS97]. -Regularity [DD99]. -spheres [Voi97]. -Stable [AL99, AT92b, PWM90, PWM92, ST91, Tal90a, AD98]. -statistic [GGM97, Stu94]. -Statistics [BGZ94, BG99, CdlPT93, FR93, GZ94, KW90, Maj94, MS92, OR93, Stu91, Zha99, dlP92, dlPMSS94, dlPMS95, KN97, LdlP97]. -Upper [HP93]. -Valued [CS93b, IMM^{+}90, Wen90, Alb96, Duh98]. -Variation [Dud92, MR92a, Qia98].

1 [GP92].

89m [BL97].

95m [Lig99]. 96b [Pin99]. 96f [Wu99].

A.S [HL96]. A.S. [Ein92]. Absence [WW98]. Absolute [Fit97, She91a, HS96]. Abstract [Enc93b, Iof91]. abundance [BCD98]. accessible [BW96b]. Acknowledgment [PR92, She91a, Yag91]. Acyclic [PWM90, PWM92]. Additive [Jai90, MR99, AP98, DK97a, DK97b, MR96a, MR96b, NdA98, TZ97]. Adjoint [DeB99]. advection [Kom96]. Age [Oel90]. Age-Structured [Oel90]. Ages [DJ92]. Aggregation [LBG92, Law95]. Aldous [AT93]. aléatoire [Pia98]. Aléatoires [Fer90a, Fer92, dF97]. Algebraic [Deu94]. Algorithms [Jer92]. Aligned [DKZ94b]. Almost [Ber91a, Ein91, HV92, Hou95a, KZ94, OS90, Sch91, Sep93, Tal90a, BG96c]. Along [ALP99]. Among [Szn93]. Analysis [CD90, DIS90, Jer92, Mik90, Yag90, Yag91, Kle97, Hsu98]. Analytic [Bas97, Bob99a, ORS95, dA96]. Ancestral [Don91]. Animals [Szn95]. Ann [BL97, Lig99, Pin99, Wu99]. Ann. [Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90g, Ano90h, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano92a, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h]. annealed [Pov97]. Annealing [Cat92]. Annihilating [Sud90]. Anticipating [MNS92, PP90]. Any [BLPS99, Sat93]. Applicable [Iof91]. Application [AL99, CPU94, Cat92, LZ94, PP94, AIM96]. Applications [Bal91, Bal95, BL90, DeB99, HP90, He93, IM94b, Kho93, LT90, Mel92, Slu94, CKS98a, DP97, DFL98, KS96, RS96, Zha97, Bra96]. Approach [BCGH93, Bus99, DIS90, FKMP95, GP92, ORS95, Pin93, Stu94, AGM96, FK97, RW97, Bas97]. Approximate [RR91, Tal94c]. Approximating [RHHQ90, Sch94]. Approximation [Ale97, BMSS95, BCL92, CC92a, Cos91, Erh99, HK97, HJ90, Hor90, LU95a, LL99, Maj90, Rio93a, Rio93b, Sch96, SK90, Kom96, Zha97, AT93]. Approximations [AT93, AST95, BDT95, Bro90, CM91, HK95, Pem90, Rob91, GM97, MS98b]. Arbitrary [BK95b, HK97, NdA98]. Area [RW91, BCD96, HL98a]. Arising [BL92a]. Armed [Dal90]. arrangements [BD98b]. Arrays [CS90, Hal91, Jan94, Ros95b]. aspects [Sch97]. Assemblies [AST95]. Associated [DJ94, MWZ93, Har96, Yu96]. Association [LRS90]. assumptions [BP98]. Asymmetric [FKS91, FF94, Sep99, Lan96, LOY96]. Asymptotic [AL99, BD94b, CC92a, CMS95, CH90, CHM91, CM94, DIS90, FJ92, GT99, HM95, HKW90, HW92, Hal90a, Hog90, Hsi95, Hsi99, IS99, JP98, Jak93, KK91, Lah93, LW94, Maj94, MS92, Mik90, Myk93, OR93, Pin90a, Por90, Rob91, SK90, TZ97, Wan95b, Wu94b, Lac97]. asymptotically [GGM97]. Asymptotics [AST95, Bal95, BL88, BH94, FS97, IM94a, IM94b, Ros92a, BL97]. attracted [AD98]. attracting [BS97]. Attraction [AT92b, Mee90, Mee91]. Attractive [BG94b, Bol94, Deu94, Lan91, Lan93, Lig91, Ste91a, Yag90, Yag91]. Attractive/Repulsive [Yag90, Yag91]. Automata [BN94, Sch92b]. Automaton [CL93, Fis92a, Ste94]. Automorphism [Pem92a, Häg97]. Autoregressive [BP92]. Average [Far90, GH94, Sto90b, Sto90a]. Averages [AT92b, DR91, Hsi99, HH97]. Averaging [FW93, Ver99]. avoided [ABP98]. Avoiding [BEI92, Tot95].

back [Ano90a, Ano90b, Ano90c, Ano90d, Ano91a, Ano91b, Ano91c, Ano91d, Ano92a, Ano92b, Ano92c, Ano92d, Ano93a, Ano93b, Ano93c, Ano93d, Ano94a, Ano94b, Ano94c, Ano94d, Ano95a, Ano95b, Ano95c, Ano95d]. Backward [CK96, CKS98b, EKP^{+}97, Pra98]. Backwards [Dar95, DP97]. Bahadur [DM90a, Ein96]. Balayage [BL92b]. Ball [LL99, Tal94d]. Ballot [Kal99a]. Balls [Hug92, Lin91a, Har96]. Banach [Hah94, Pin99, Che93a, Che94b, Ein93, Kla90, Pin94]. Band [Hou95b]. Bandit [Dal90]. Barbour [AT93]. Based [KW90, Tue93, DE96]. Bass [Hsu98]. Bayes [BS90]. be [GJW99]. Behavior [AL99, Ber91a, CH90, Ein91, Ein92, HW92, IS99, KK91, LPP95, MSS94, Maj99, Nob92, Pin90a, PP93, Sch92b, Tri92, Wan95b, Wu94b, CDM97, DP98, DE96, EM96, Wüt98]. Belong [Mee90]. Benedicks [CMS99]. Bernoulli [Ale90, AP96, KW92, KZ90, Zha99]. Bernoullicity [Ste91a]. Berry [BG96a, Blo99, HK95, JW99]. Bessel [Lee90, PY98]. Best [Hit90, Jac91, CK97]. Between [CD95, GMdB99, GH94, KK94, MSV90, PP94, GMdB03]. Bi [PW94]. Bi-Tree [PW94]. Bifurcation [BB99]. Big [FGL94]. bilinear [KN97]. binary [Lig96]. Binomial [Woo92]. Block [BGM93]. blockwise [Pel98]. Boltzmann [GM97]. Bond [DG99, Tot95, GS98]. Book [Bas97, Hsu98, Wih97]. Books [Oco90]. Bootstrap [And93, CC99, Hal90a, Sch92b, Sep93, AGM96, Pel98]. Bootstrapping [GZ90]. Bootstraps [PW93]. Borderline [Tal95a]. Borel [MPS91, MS93]. Bottlenecks [CM91]. Bound [Blo99, Din91, JW99, BG96a, GM96, HK97]. boundaries [CS96]. Boundary [CMS99, Cos91, Day92, Don91, FW92, Hus90, NP91, PW94, Pin93, Rob91, Roc92, Sow94, CDM97, Lot96]. Bounded [Ham95, Hou91, Pin87, Pin90b, SC92]. Boundedness [Ale96, GM95, Tal90b]. Bounding [Mar96]. Bounds [Ale90, BG99, Bro90, DEW91, Jai90, Kla90, Lig95, Pin94, Tal94a, dlPK94, KN97, NdA98, Pin99, Shi97, Wan97, dA97]. Branching [Ath92, BCGH93, Big92, BCG93, Cha91, DLM95, Del99, Dyn91, EP99, FG92b, GW91, LS92, Le 91, LL98b, LL98a, Lee90, Lee91, Lee92, LPP95, Sud90, Wu94b, BK97, BCG97, DFL98, DP98, DR98, Kle97, Kle98]. Bridge [Pit99b]. Brownian [Dur93, ABP98, AFL91, AR93, AS95, Ald94, Ald97, AL95, BP94a, BH91, Bis92, BH94, BD98a, Bra92, Bur90, BL90, BT95, BW96b, Car97, Cha91, CMS99, CM96, DM96, DFLM95, DeB96, DeB99, DD99, DR98, DL98, DMY97, DFST96, DW94, FT96, Gol98, Gri99a, GK98, GJW99, Ham97b, HL98a, HW90, Hug92, Kes93, Kho92, Kho94, Kho95, KS99, Lac97, Lal92, LW99, Le 91, LP95, Lee91, Lee93, LR95, Leu97, LZ95, MSV90, OS99, Per90, Pin90a, Pin95, PY98, Pit99b, Pov97, Rob91, RY91, Sch92a, Shi98a, Szn93, TZ97, Tal94d, Tal95b, Ver95, Ver96, Wüt98]. brownien [Leu97]. brownienne [Gol98]. Bubbles [Kho95]. Burgers’ [GN99]. Burnside [Sto95b].

Calculus [Oco90]. Cambridge [Sha98]. Can [Sat93]. Cantelli [GH92, Sch93a]. Capacitary [IM90]. Capacity [BPP95, FK92, KS99, Lyo92, Szn97]. Case [BDT95, Har99, Jac91, Lig91, Shi92c, BBE97, MS97]. Catalyst [DFLM95]. catalytic [DP98]. cell [Gol98]. Cellular [BN94, CL93, Fis92a, Sch92b]. cellule [Gol98]. Censored [GG99, HKW90, DE96]. Censorship [GL90]. Central [GMdB03, GT98, He93, KdHvdH97, PU97, AGM96, Gen96, Kön96, Pia98, Ren96, SX96, AMZ90, Ber92a, CPU94, Che91, Cog91, GMdB99, Gou93, Han89, Han91, Led95, PP93, Tan93]. Certain [AMZ90, Als92, HV92, Iof93, Raj93, Raj94, BD98a]. Chain [Che99, Don91, NdA98]. chaining [Tal96a]. Chains [BP94b, BPP95, Cog90, Cog91, Erh99, HSC93, Hog91, Ore91, Sep94, Wih97, dTB91, dA97, Num98]. Change [FHU91, Nai90, DFST96, FT96, Jan98, Jan02]. Changes [Kur91].

Changing [HV92]. Chaos [Zhe95, MR96a, RS96]. Chapman [CH94]. Characteristic [AD98, And93, Day92, PR91, PR92]. Characteristics [BP93b]. Characterization [DD99, GZ91, KM94b, Tal90a, dTB91, BB96]. Characterized [BK93]. Characters [Ros94]. Charge [Adl90]. Chemical [Blo91, AP96]. Chen [BB92a, HJ90]. Cheng [Sta90]. Choosing [Pem91]. Choquet [FK92]. Chung [EM94, Kes97, KLT94, Rem94]. Chung-Type [EM94, Kes97]. circle [Nor98]. Circular [Bai97]. Claims [MS95]. Class [BL92a, BB92b, BL88, CGST93, CdlPT93, Cos91, GW91, MNS92, Sim95, SvdB99, Wen90, Zhe95, dlP99, BL97, DFL98, KX96, PW98]. Classes [Alb94, AL95, Dud92, GZ91, SW92, Bun97, Lac97, vdV96]. Classical [SL95, Kal96, LS97]. close [Ber98, Fel96]. closed [Leu97]. CLT [Wu99, BG96b, Got91, Wu95]. Clumping [AT93]. Cluster [Ein95, EF96, Gri95, KZ90, Sep98a]. Clustering [Fis92a, Kle98, Kle96]. Clusters [BLPS99, LS99a, Rev93, Häg97, Kle97]. Coalescent [DJ92, Ald97, AP98]. Coalescents [Pit99a]. Codes [KW92]. coding [Shi97]. Codings [SvdB99]. coefficient [LOY96]. Coefficients [CEF91, PP90]. Coexistence [Lig94, DP98, Lig99]. Coexistent [Han99]. Coffman [Jer92]. Collectives [PF91]. Collisions [Pit99a]. Color [Fis92a]. Combinatorial [AST95, RW97]. Commutation [Che93b]. Compact [Pin93]. Comparison [Blo91, DSC93]. Compensated [Kho93]. Compensators [BB92a]. Complete [Han99, GK98, SS98, Zha96b]. Completely [MvW91]. Component [Bob99b]. Composition [Bun96, Gne97]. Compound [BCL92, Erh99, Kho93, Wan91b]. Compression [OS90]. Computer [KK94]. Concave [Bob99a, Van95, Bob96]. Concentration [Bar94, Dem97, Mar96, RS96]. Concept [SL95]. Conceptual [LPP95]. concernant [Gol98]. Concerning [HP93, Gol98]. Concomitants [GH94]. Condition [BA91, EK93, KL92, NY93, vdB97]. Conditional [AL99, Day92, Gog94, Hu91, Lee91, LZ99, MWZ93, Sch93b, SZ92, Stu91, Wan91a, Zhe95]. Conditionally [Hit94]. Conditioned [CMS95, GJW99, Kee92, Lee90, Pin93]. Conditioning [BD94a, Kal99b]. Conditions [Bra93, BS90, Cos91, DL94, FM93, FW92, KK95, Kur91, MS92, Tal90b, Tal92, GT98, Mue96]. Cone [DeB99]. Configurations [COP99, Lu95b]. Congruence [Elo90]. Conjecture [Kri94, PR91, PR92, Ben97, Gol98]. Connectivity [Ale90]. Consequences [DD99]. Conservation [Lan93, Qia96]. Conservative [JLQY99]. Consistent [MS94, MS96]. Constant [Mas90]. Constants [Hit90, Jac91, Wan99, BH97]. constraints [CKS98b]. Constructing [Dar95]. Construction [AR93, EP99, Pol92]. Constructions [Rio93a, Rio93b]. Constructive [Num98, vdB97]. Contact [BG90, BG91, BDS91, DS95, Kle94, Lal99, Lig92, Lig95, MSS94, Pem92b, SS99, LS98, Lal02, Lig96, NV96, NW97, SS97, SS98, Sta96, Zha96b]. Contents [Ano96]. Continues [Str90]. Continuity [CC92b, Imk93, IMM^{+}90, Tal92, Zab92, Fit97]. Continuous [AL93, DFL98, GW91, Str90, Tal90a, Wan95a, EF96, MR96b, Dur93].

Continuous-Time [Wan95a]. Continuum [Ald91, Ald93, Fon93, Roy90]. Contrôle [Leu95]. Contraction [dlPMSS94]. contractions [Pit98]. Contractive [Ste99]. contrepartie [Gol98]. Control [Hij90, Sto90b, Sto90a, Leu95]. controlled [BB96]. Convergence [ALP99, AT92b, Bai93a, Bai93b, BGZ94, Big92, Boi93, Cog90, DH95, Don91, DJ92, GZ94, Gog94, Got91, Hal90b, Han99, HP90, HV92, Hil92, Hou95a, Jak96, Kho94, KR91, Kur91, LO98, Lig91, Sch93a, Sil90, Tue93, Wan91b, Xia94a, Yu94, Ale97, BG96b, GM97, HL96, Pri96, Ren96, RdH96, SS98, SY96, Voi97, Zha96b]. Convex [Bár99, Kho92, LO99, Raj93, Raj94, CK97, Ver96]. Convolution [Sat93]. Convolutions [Bro90]. coordinates [Ker96]. Correction [Alg95, BL97, FK02, GMdB03, Han91, Lal02, Lig99, LP03, MS96, NP95a, Nol92, Pin99, Pin90b, PWM92, Raj94, Wu99, Jan98, Jan02]. Correlation [Har99]. Correlations [Fon93]. Corresponding [EF93]. Cost [SC92, Tal94b]. Countable [Num98, DK96]. counterpart [Gol98]. Coupled [Fre91]. Coupling [EK98, HJ90, Mue93, O’C95]. courbure [dF97]. Covariance [Bai93b, BY93, BS99, HPA95, Sil90, BS98]. Covered [Rev90]. covering [PT96]. Cramér [MW90]. Creating [Sha98]. Criteria [GW91, LPP95, HL96]. Criterion [Kal91]. Critical [Ath92, BA91, BLPS99, BG90, BG94b, BD93b, BCG93, DKZ94a, Deu94, GW91, GS98, Lee90, Lee91, Lig91, Lig95, MSS94, Mou92, PP95a, Pov97, Roy90, Sch92a, Ste94, Ald97, BBE97, BCG97, DR98, MS98b]. Crofton [Gol98]. Crossing [ALP99, Szn95, Lot96]. Crossings [Far90, Hus90, PW98, Slu94]. cube [Bob97]. Current [FF94]. Curvature [Led95, dF97]. Curve [Slu94]. Curve-Crossings [Slu94]. Curves [DZ95, BCD96, DM96]. cut [Por96]. cut-off [Por96]. Cycle [AT92a, Lal96]. Cyclic [Fis92a].

D [AT93, dA96, vdB97]. D. [AT93, Dur93, Gol98]. Dans [Fer92]. Darling [Ber98]. Data [GG99, OS90, DE96]. Dawson [Sch97]. Decay [BG91, Deu94, Fon93]. Decaying [BH94]. Decision [Sha98]. decomposable [Bun97]. Decomposition [BT95, LZ94, Mee91, MS95]. Decoupling [dlP92, dlPMSS94, dlPMS95]. Decreasing [DFST96, FT96]. Defined [KH94, Sil90]. Degenerate [BM95]. Degree [OR93]. Delay [BM95]. Denjoy [Bis92]. Denormalized [CdlPT93]. Densities [BM95, Imk93, IN94, KK94, Roy90, Wen90]. Density [AFL91, AR93, BD93a, Blo94, Mik90, She91b, Xia94b, AN98, Blo96, DP96]. Density-Dependent [Blo94]. Denumerable [Pol92]. Dependence [BF90, HP90, Ver99, DFL98]. Dependent [Blo94, Bry92, BGM93, Dal99, DH95, De 93, EP99, Lah93, Rio95b, Wu95, Blo96, Häg97, SY96, Wu99]. Derived [Pit99b, PY97]. determined [BMR97]. Deterministic [Blo91, Lu95b, MSS94]. Deviation [BET99, Bry92, Cat92, CS93a, CC92a, Din93, IKL92, Jai90, Lee92, PP99, BET02, BG96c, NdA98]. Deviations [Bal91, BD93b, BDT95, Din91, DZ92, DEW91, EF93, Fen94, Geo93, GdH94, IL93, Iof91, Lac90, Lan92, Lee95, LR95, Mik91, MNS92, Mog93, Qua95, RHHQ90, Sep94, Sow92, Ver99, Wu94a, Wu95, Col96, DR98, Gra97, HS96, KX96, Pov97, Rez98, Sha97, Wu99, Zer98, dA97]. Dichotomy [FK92]. Dies [BG90]. Difference [GH94, SC92, BBE97, CK97]. Different [Kle96]. Differentiability [Dud92, dS92]. Differentiable [Slu94]. Differential [AFN95, AI99, BMSS95, Bur94, DD99, Dyn92, Dyn93, Har93, Kat91, KH94, KP91, Lee93, LS99b, MNS92, MS99, NP91, San93, Wan95a, CK96, CKS98b, JP98, KX96, MS97, PT97]. Diffusion [ALP99, Bal91, Bas97, Blo91, Blo94, Blo96, CGST93, CM91, Cos91, DN94, FK99, FK02, Fre91, FW92, FW93, Gri99a, Kom96, LS92, LBG92, Law95, Nob92, She91b, Sow92, Sow94, Stu98, Wic90, dL95, Ata98, DK96, LOY96, Pra98]. Diffusions [Adl90, ARZ93, BB92b, Bor93, CEF91, CMS95, CG94, Day92, EP99, Iof93, Lee92, PZ95, Pin87, Pin90b, Pin93, PP93, CDM97, Fit97, Ker96, Kle96, Pin96]. Dilatations [LO99]. dilation [Shi98b]. dilation-stable [Shi98b]. Dimension [Kal95, MSS99, Tal94b, Xia95, Jan98, Jan02]. Dimensional [ARZ93, And93, BY93, BS99, Bol94, BN94, Bus99, COP99, CC99, CMS95, DeB90, FGL94, Fis92a, FHU91, LS92, LP95, LR95, Lig92, Lu94, Neu90, San93, Yag90, Yag91, Ale96, BS98, BS97, CS96, FU96, Ham97a, Kön96, Lan96, LN96, NV96, NW97, Pov97, Wüt98]. Dimensions [Ale90, BCG93, BEI92, NP95a, NP95b, NY93, BCG97, DF98, FG97, HM98, Pen96]. Direct [Cog90]. Directions [Ale90]. Dirichlet [ALP99, CL94, IM90, LZ94, Pin90a, PY97]. Disagreement [VM94]. Disc [Rev90]. Discontinuous [CEF91, DEW91]. Discrete [Bar94, CM91, Kif90, Sch93b, Bob97]. Discrete-Time [Kif90]. Distance [GMdB99, Szn96, AP96, GMdB03, Mar96]. Distances [KK94]. Distinguishing [OS94]. distributed [BP96]. Distribution [BB92b, CHM91, CM94, DKZ94b, DMS93, EH92, FG92a, GZ94, Gog94, GT99, GL90, HKW90, Ham95, Kho95, Kur91, PR91, PR92, PWM90, PWM92, dMK99, dlC90, BS98, PY97]. Distributional [Szc92]. Distributions [Bai93a, Bai93b, Bob94, CL94, CD90, DK91a, FR93, FKMP95, GMdB99, GP92, Hal90a, HK91, HS90, Hsi99, Hug92, LW94, Lee91, Maj94, MS94, OR93, Pin94, RR91, RS93, Sat93, Bob96, GMdB03, Har96, JP98, Pin99, MS96]. Divergence [NP95a, NP95b, Mar96]. Divisible [Raj93, Raj94, Ros90b, RS93, ST93, ST94, Szu91, Tal92, Tal93]. Doeblin [Lin91b]. Does [Che99, ZZ92]. Domain [AT92b, Bal95, CMS99, Mee90, Ata98, GT98]. Domains [BH91, Bis92, DI93, Mee91]. Domination [Hit94, LSS97, PP94]. Domino [BP93b]. Donné [Str90]. Donsker [Dud92, GZ91, SW92, vdV96]. Doob [Jac91]. Doob-like [Jac91]. D’Ornstein [Fer92]. Double [BET02]. Doubly [BET99]. Drift [BM95, Kat91, Pin90a]. driven [Jak96, PT97]. Dual [Roy90, KS96]. Duality [DF90, Kal99b, SL95, LS97]. D’un [BL99, Leu97, dF97]. d’Une [Leu95]. Dvoretzky [Mas90]. Dykstra [Win90]. Dynamical [FKMP95, Mik90]. Dynamics [Cha94, JLQY99, AG97, GG96, LSV96]. Dynkin [CK96].

E. [Jer92]. Eating [Liu95]. Eaton [PR91, PR92]. Edge [Pen99, Sch92a]. Edgeworth [KW90]. Edwards [KdHvdH97]. eigenfunction [FS97]. Eigenvalue [BY93, dMK99, FS97]. Eigenvalues [BS99, Sto95b, Sto95a, BS98, Szn97]. Eigenvectors [Sil90]. Elementary [Maj99, Bob97, Ein96]. elements [Tho96]. Ellipsoids [MWZ93]. elliptic [CZ96, DP97]. Embedded [Ber99]. Embedding [Kho93, Sho91]. Empirical [BK95a, CH90, DM90b, DM92, DM94, DM95, DK91b, DK91a, Ein92, Fen94, GZ90, GMdB99, Hor90, IL93, Maj90, Mik91, PW93, Rei95, Sch93a, SY91, Tal94a, Tal94b, Wu94a, Yu94, DE96, GMdB03, Pel98, Qia98, SY96, dA97]. Endpoint [DR91]. Enlargement [Ant95, Szn97]. Entrance [Don91]. Entropy [BP93b, FG97, Geo93, KW92, LL99, MS94, Sep98a, Shi92b, Wen90, Yag90, Yag91, MS96]. Enumeration [DH96]. Envelopes [PF91]. Environment [BCGH93, Boi93, BDS91, DLM95, FG92b, GdH94, Kle94, LP92, PP95a, PP99, SZ99, Wic90, CMP98, HS98, Kom96, LP03, Pia98, Zer98]. Environments [CL92, Cog90, Cog91, Lig92, MSS94, PP93, Myt96, NV96]. environnement [Pia98]. Epidemic [LU95a]. Epidemics [Zha93]. Equality [Roy90]. Equation [Cha94, CdlPT93, CH94, GN99, Kat91, Lan91, MSS99, Mue93, Sow92, BBE97, DF98, Dav96, GM96, LT98, Mue96, Mue97, Myt98]. Equations [AFN95, AI99, BMSS95, BM95, BK93, BKK95, Bus99, DD99, DN94, Dyn92, Dyn93, Fre91, FW92, KH94, KP91, Lee93, MNS92, MS99, NP91, PP90, San93, Sow94, CK96, CKS98b, JP98, KX96, LN98, MS97, PT97, Har93]. Equilibrium [JLQY99, Lal92, Lan93, Lu94, Nob92]. Equivalent [Bra93]. Erdos [Ber98, Gan98, KLS96]. Ergodic [Bac92, CG94, Hou92, Sch91, Shi92c, Wei91, AN98, EK98, dF97]. Ergodicity [BCG93, DMT95, Fer90b]. ergodique [dF97]. Erratum [BET02]. Error [Bro90, JP98, Tak97]. Escape [Pru90]. espace [dF97]. Esseen [BG96a, Blo99, HK95, JW99]. estimate [Tak97]. Estimates [ALP99, BGZ94, BD94b, BL90, CK91, Cat92, HSC93, LL99, Rev90, She91b, Sow95, Wan99, GM97, MS98a]. Estimating [KK94]. Estimation [AP91, Lat97, MS94, AN98, MS96]. Estimator [GL90]. Estimators [Hal91, Xia94b]. Euclidean [Ein95, IS99, MS98a]. Euler [JP98, PT97]. Evolution [BK93, LN98]. Evolutions [Wat96]. Exact [BS99, Bal95, Kho94, Sep98b, Ren96]. Examples [CPU94, MS97]. Exceedances [DK91b, DK91a, HJ90]. Exceptional [AL95]. Exchangeability [Woo92]. Exchangeable [DZ92]. Exchangeably [PW93]. Exclusion [FKS91, FF94, FHU91, Lan92, Liu95, Sep99, Yag90, Yag91, Jan98, Jan02, LOY96, SX96]. Exclusion-Eating [Liu95]. Excursion [Ald94, Pit99b, VA94, HL98b]. Excursions [Le 91, Ald97]. Existence [FKMP95, GKRS91, Imk93, IN94, Sep99, Ska93, SvdB99, Sto90b, DeB96, Kal96, Mue97, Sta96]. Exit [Bal95, DeB90, FJ92, GM92, GM94, Hug92, IM94a, IM94b, Mik91, Har96, She96]. Exits [Day92]. Expansion [KW90]. Expansions [IL93, Lah93, LW94, Myk93, Pit98, Por90]. Expectations [Gog94, dlPK94, KN97]. Expected [Bai93a, Bai93b, Por90]. Explicit [EP95, KW92, Wan97]. exploration [LL98b]. Explosion [PP94]. Explosions [KK95]. Exponential [Ata98, Ber91a, BG91, Bob94, Bro90, Cat92, DMT95, FGL94, JW99, LZ99, Ros95b, SZ92, dlP99, HK97]. Exponents [And93, BA91, BL90, Jur92, LW99, Lia97, MS97, Szn96, Zer98]. Expression [Hog90]. Extended [Xia94a]. Extension [Ber92b, Mat91]. Extensions [GK91, Sch91]. Exterior [BG94a, Pin90a]. Extinction [Kle94, Tri92, Pin96]. Extremal [Alb90, Bob96, SS99, Alb98, SS97]. Extreme [CHM91, Hus90, Jan94, RdH96]. extreme-value [RdH96]. Extremes [DR91].

F [Bra96, Hsu98]. Factors [BGM93]. Fair [KL92]. Familiar [Bor93]. Family [CD90]. Fast [Mar99]. Fayolle [Num98]. Feller [PZ95, Qia96]. fermé [Leu97]. Field [AFN95, BH94, COP99, Che93b, Fen94, FG92b]. Fieldo [GH90]. Fields [AP91, BD93b, Bra93, Bry92, Che91, Geo93, Nol88, Nol92, SvdB99, Sun93, VM94, Van93, Xia95, Ale96, DFST96, FT96, Tho96]. Filtering [BKK95, Enc93b, Ata98]. Filtration [BP94a]. filtrations [Fel96]. Final [LU95a]. finance [CKS98a, KS96]. Finetti [AL93]. Finitary [SvdB99]. Finite [Blo99, CC99, CL92, DR91, DSC93, GH90, Hog90, Hog91, KZ90, KW90, Lal93, MS98b, PV99, Sho91, Tow95, Woo92, Zha93, DH96, Pin96, Uch98]. Fires [Zha93]. First [Als92, DeB90, GG96, GM92, GM94, HM95, Jan98, LW94, LP92, FS97, LN96, LP03, Sep98b, WW98, Jan02]. first-order [FS97]. First-Passage [LP92, LN96, LP03, Sep98b, WW98]. Fisher [dS92]. Fitting [Win90]. Fixation [DS93]. Fleming [DK96, EG93, EK98, ORS95, Sch97]. Flow [CM91]. Flows [Har93, IS99, FK97, Fla96, Lia97, MS98a, Pit98]. fluctuation [GK98, Ham97a]. Fluctuations [CFR91, FF94, Law95, Lu94, NP95a, NP95b, Sch92a, Szn96]. Follmer [MS95]. fonctionnel [Pia98]. Fonctionnelle [Tou90]. Fonctions [Fer90a, Fer92]. Forced [Kat91]. Forest [Pit99b, Zha93]. Forests [Ale95]. Form [DF90, Hil93, Szc92, Kes97]. formation [EF96]. Forms [CGZ95, GT99, IM90, LS99b, Slu93, dlPMSS94, dlPK94, BG96b, Gad98, GT98, KN97, RS96]. Formula [Fan99, Van95, ZZ92, Tak97]. Formulae [RY91]. Formulation [Sto90a]. fortes [dF97]. Four [BEI92]. Fourier [MR96b, Tal90a, Tal92, Tal95a]. Fractal [BL90, CD90, DM95]. Fractional [AS95, Tal95b]. Frank [Ano93i, Ano93j, Ano93k, Gri93, Kes93]. Frechet [Dud92]. Free [Lal93, BP96]. Freidlin [Mik91]. French [BL99, Cép98, Fer90a, Fer92, Gol98, Leu95, Leu97, Pia98, Str90, Tou90, dF97]. Friedlin [Kif90]. front [Ano90e, Ano90f, Ano90g, Ano90h, Ano91e, Ano91f, Ano91g, Ano91h, Ano92e, Ano92f, Ano92g, Ano92h, Ano93e, Ano93f, Ano93g, Ano93h, Ano94e, Ano94f, Ano94g, Ano94h, Ano95e, Ano95f, Ano95g, Ano95h]. Frontier [LS92]. Frustrated [DG99]. Full [PW94, Ver99]. Function [Ale90, Din93, DK95, EG93, GL90, PR91, PR92, Ste99, Van95, Wan91a, CP97].

Functional [DM90b, DM92, DM94, Gan98, Gou93, GP92, GL90, Han89, Han91, KLT94, MW90, Rio95a, Kal96, MS97, Vit96, Pia98, Tou90]. functional-differential [MS97]. Functionals [Arc94, DK91b, ES93, EP95, HPA95, Hsi99, IL93, Jai90, Kho92, MR99, RS93, SZ92, BD98a, DE96, DMY97, DK97a, DK97b, Har96, HH97, LL98a, MR96a, MR96b, NdA98, TZ97]. Functions [ALP99, BLBE91, Bar94, Bas97, DW94, GZ91, GM90, HS90, Kal94, Pin87, Pin90b, SW92, Sil90, Wan99, Zha99, dlC90, AD98, Ale97, BS96, CDM97, CK97, CMP98, Fer90a, Fer92, KN97]. Fusions [EH92].

G [Bas97, Gol98, GP92, Jer92, Num98]. gains [CKS98b]. gains-process [CKS98b]. Galton [PP95b]. Gambler [GM94]. Gambling [Alg92]. Game [Gne94]. Games [KL92, MS93, CK96]. Gamma [Dar95]. Gamma-Martingales [Dar95]. Gap [FGL94, LSV96]. gaps [MP96]. gasket [Ham97b, HL98a]. Gauss [Bob97]. Gaussian [AMZ90, Alb94, Alb96, AP91, Arc94, BG99, Ber91b, Ber92a, BD93b, CP97, CS93b, DM99, DP96, Fer90a, GZ91, HP93, Har99, HSC93, HS90, HJ90, Hus90, KLT94, LO99, LL99, Lin91a, Lot96, MR92b, MR92a, MR96a, MWZ93, SSZ98, Slu94, Sow92, Sun93, Tal94a, Vit96, Voi97, Xia95]. Gaussiennes [Fer90a]. Gel [PWM90, PWM92]. Genealogical [DJ92]. General [BL88, Del99, DEW91, Ein93, GZ90, MS95, PS99, PW93, dlP99, BL97, PT96, Glo90]. Generalization [Fin92]. Generalized [Als92, BP92, Ein95, Gou93, LU95a, Mee91, Tót96, Tow95, GM97, KN97]. Generating [Mat91]. Generators [Xia94a, LN98]. generic [Tal96a]. Genetics [Don91]. geodesic [Har96]. Geodesics [LN96, WW98]. Geoffrey [Bez92]. Geometric [Bor93, BP93a, Bro90, Sch97, dL95]. geometrically [Zha97]. George [Jer92]. GI [GP92]. GI/G/1 [GP92]. Gibbs [FP97, OS99, vdB97]. Gibbsian [GH90, Szn93]. Ginzburg [Lu94]. Girsanov [ARZ93, ZZ92]. Given [GKRS91, Ska93, Str90]. glass [AG97]. Glivenko [GH92, Sch93a]. Global [FP97]. Googol [Gne94]. grands [dF97]. Graphs [Ale95, FW93, Tho92, Ald97, BS96]. Gregory [Bra96]. Grid [SY91]. Grimmett [Bez92]. Group [BLPS99, Sto95b, dS92, LPP96]. Groups [Bar94, DSC93, Fan99, Glo91, HSC93, Lal93, Lia92, Sto95a, Sto95c, DH96]. Growth [Che92, Che94a, KZ94, Lal99, Lal02, Pin95, GG96]. GZH [He93]. GZH-Semigroups [He93].

H. [Har93]. half [Bob96, WW98]. half-plane [WW98]. half-spaces [Bob96]. Halmos [KS96]. Harmonic [Bas97, Ker96, Pin87, Pin90b, Tal92, BS96, CDM97, Duh98]. Harnack [Wan99]. Harris [Che99]. hasard [Pia98]. Hausdorff [LP95, Tal95b, Van93]. Hawkes [BM96]. Heat [Mue93, Mue96, Myt98, Stu98, Wan97]. Heavy [Hal90a, SK90, Szc92]. Heavy-Tailed [Hal90a]. Hedging [MS95]. Height [Pit99b]. heights [CP97]. Heisenberg [PS99]. Hereditary [BM95]. Heuristic [AT93]. Heyde [BK97]. Hierarchical [BEI92]. hierarchically [EF96, Kle96]. High [Hus90, NY93, BCG97, Pen96]. Hilbert [De 93, Ein91, FG92a, PZ95]. Historical [BP94a, EP95, Sha98, Ver95]. Hit [Pin93]. Hitting [BG94a, Col96, Rob91]. Holder [BMSS95, BH91, Fin92]. Holomorphic [CDM97]. Holst [AT93]. Holtsmark [PWM90, PWM92]. homeomorphisms [Pit98]. Homogeneous [Arn94, Lal93, Lal99, LS98, Lal02, SS98]. Homogenization [Bal91, FK97]. Hopfield [BG96c, Gen96]. hull [Ver96, Kho92, Van95]. Hurst [Dal99]. Hydrodynamic [CS96, FHU91, Lu95b, Qua95, Jan98, Jan02]. Hydrodynamical [Lan91, Lan96]. Hydrodynamics [Sep99]. Hypercontraction [KS91]. Hypercontractivity [Szu90]. hypergroups [Voi97]. hyperplane [BD98b].

I.I.D [CGZ95, DK91b, DZ95, Din91, KK91, Muk94, SC92, Ber98, HK97, Maj99]. identically [BP96]. Identifying [Din93]. Identities [HPA95, PY98]. Idle [GP92]. II [Bai93b, FT96, IM94b, Muk94, Rio93b, SS99]. II. [BL90, MS97, SL95]. IID [Ein91]. III [Ald93, Hij90]. Impedances [BP93b]. implies [SS98]. Implosion [Ken95]. Improved [Lig95]. Improvements [HK95]. Increase [Don96]. Increment [Tal90b]. Increments [CC92b, CS93b, DM92, DM95, MR92a, Mog93, Ste91b, EM96]. Independence [Hsi95, Kal91, RR91, Tal96b]. Independent [Che93a, Hil92, Hit94, Jak93, Jan94, KS91, LT90, Lee95, LRW92, Mog93, CK97, Lat97]. Index [Dal99, Nol88, Nol92, Xia95]. Index- [Nol88, Nol92, Xia95]. Indexed [Arn94, BP94b, BET99, CS90, DM94, Eva92, Rio93a, Rio93b, BET02, Enc98, Nor98]. Indices [LRW92, PT96]. Indistinguishability [LS99a]. Induced [IM90]. Induction [Sha98]. Inequalities [Ber91a, Bob94, Bob99a, Bob99b, CPU94, CS90, CC92b, Fin92, Ham95, HPA95, IM90, Jac91, Kho92, KS91, ST93, ST94, Sha95, Tho92, Wan91a, Wan95a, Wan99, Zha99, dlP92, dlPMSS94, dlPMS95, dlP99, CK97, Dem97, RS96]. Inequality [Fan99, Fin92, Har99, Hit90, Mas90, Rio95b, SC92, Bob97, LY98]. Inf [KLT94]. Infinite [AFL91, ARZ93, Ale95, BBK94, BLPS99, CG94, DH95, Häg97, Hsi99, KW92, KM94a, Roc93, BS96, FG97, Kle98]. Infinite-Variance [Hsi99]. Infinitely [Raj93, Raj94, Ros90b, RS93, ST93, ST94, Szu91, Tal92, Tal93, BMR97]. Infinity [Maj99]. Information [Ano90i, Ano91i, Ano92i, Ano93l, Ano94i, Ano95i, Dem97, Van93, dS92]. informational [Mar96]. inhomogeneity [NW97]. Inhomogeneous [CL92, MSS94, Zha94, Lan96]. Initial [FR93, Lu95b, Mue96]. Instant [Str90]. Integer [Pem91]. Integrability [Ros95b]. Integrable [Wan95b]. Integral [EP95, Ham95, Imk93, She91a, Ahn97, Tak97]. Integrals [CK91, Kal91, KP91, LZ99, LS99b, Rem94, RW91, Sch94, Szu91, Jak96, RS97, RW97]. Integrated [Bas97, GJW99]. Integration [Bur94, Fan99, IN94, Tow95]. Integrator [Ham95]. Intensities [AL99]. Interacting [Adl90, CG94, DPPV94, Fen94, Geo93, Gri93, Sim95, Sud90, SL95, Kle96, Lig97, LS97, Tót96]. Interaction [Bol94, Gri99a]. Interactions [Gri99b, Zha93, vdB97]. Interactive [ORS95]. interface [BdH97]. Intermediate [CM94, DS95, Sta96]. Intermittency [Sow95]. Intermittency-Type [Sow95]. Internal [LBG92, Law95]. Interpretation [Der99]. Intersection [AFL91, AR93, BBK94, Ber92a, LW99, MR99, Ros90a, SSZ98]. Intersections [Ber91b, RY91, Ros92b, Bra96]. Interval [dlC90]. Invariance [FK99, FK02, HP93, dS92, Yu96]. Invariant [BK93, BCG97, CH94, Glo91, Lal99, Mik90, Mue93, Pem92a, SS99, Häg97, Kle98, Lal02, SS97, Sep98a]. Inversion [Gan93]. Involving [dlPK94]. Irreducibility [PZ95]. Irreducible [BL88, BL97]. Irregular [Hou95b]. Ising [Fon93, Neu90]. Isoperimetric [Bob94, BH97, Bob99a, LT90, Tho92, Bob97]. Isoperimetry [Hah94]. isotropic [Pit98]. Itô [VA94]. Iterated [BK95a, BP93a, BS95, CK91, Che93a, Che94b, DM90b, DM92, DM94, DL94, Ein93, EM94, Ein95, Fis92b, Gan93, GG99, GL90, Hae93, Hal91, KH94, LRW92, LZ99, MR94, Mas94, Rem94, Rio95a, Ste99, HS96, Kes97, Tou90]. Iterative [Win90]. Itéré [Tou90]. Itself [OS94]. IV. [LS97].

J [Glo90]. Jackson [IM94b]. Jacobi [Voi97]. Jan [Sha98]. Janson [AT93]. Joint [HKW90, MS94, Har96, MS96]. Julia [Lal92]. Jump [DK95, IM94a, IM94b, KK95]. Jumps [Fen94].

Kac [COP99]. Kendall [Gol98]. Kernel [Xia94b, FU96, Tak97]. Kernels [MvW91, Stu98, Wan97]. Khintchine [dlP92]. Kiefer [Mas90, DM90a, Ein96]. Kiefer-Type [DM90a]. Killed [CMS99, DeB99]. kinetic [Rez98]. Kingman [Sch91]. KMT [Rio93a, Rio93b]. Knight [Tót96]. Kolmogorov [CH94, DL98]. Konig [MPS91]. KPP [LT98]. Kunita [Har93]. Kuratowski [LO98].

L [AT93, PR91, PR92, Wih97]. Ladder [CP97]. lamplighter [LPP96]. Landau [Lu94]. Langevin [AG97]. Laplace [BDT95, LL98a]. Large [Bai93a, Bai93b, BY93, BS99, Bal91, BD93b, BDT95, BET99, Bry92, CL93, Cat92, CS93a, CC92a, CM96, CS93b, CGZ95, DK91b, DK91a, Din91, DZ92, Din93, DEW91, EF93, Fen94, Gad98, Geo93, GH92, GdH94, GM94, HS96, IL93, Iof91, Jai90, KX96, Kat91, KZ90, Lac90, Lan92, Lee92, Lee95, LR95, Lin91a, Mik91, MNS92, Mog93, NdA98, O’B99, Pin95, PP99, Qua95, Rei95, Rez98, RHHQ90, Sep94, Sow92, SZ99, Ver99, Wu94a, Zha96a, dF97, dMK99, BS98, BP96, BET02, BG96c, Col96, DR98, FS97, Pov97, Sha97, Zer98]. Large-Deviation [Lee92]. large-dimensional [BS98]. Largest [Rev90, HS96]. Lattice [BEI92, Deu94, Pem91, Szn95, LT98]. Lattices [Yag90, Yag91]. Law [BP93a, BS95, Che93a, Che94b, CM96, DL94, Ein93, EM94, Ein95, Fis92b, Gad98, Gan93, GH92, Hae93, KH94, LRW92, Mas94, MSS99, Rem94, Rio95a, Szc92, Tal94c, Xia94b, Zha96a, AGM96, Bai97, BP96, Ber98, HS96, Kes97, Tou90, Pen99, Rei95, SZ99]. Lawler [Bra96]. Laws [AR93, AL99, BK95a, CL93, CMS95, CGZ95, DM90b, DM92, DM94, GG99, GL90, Hal91, Kal99a, MR94, MW90, ST91, Ste91b, dF97, AD98, Bun97, DMY97, Gan98, RS96, Rio95b, Zha99, Deh97]. Lectures [O’C95, Wat96]. Ledoux [Hah94]. Lemma [BT95, MPS91]. Level [Far90, Wu94b, Ale96]. Lévy [Ber92b, Car97, DW92, Don96, FG92a, Lac90, LL98b, LL98a, MR94, MR96b, MR99, MP91, PT97, PT96, RS96, ST94, Shi98b]. Liapounov [GM96, Lia97]. Lie [Jur92, dS92]. Like [ZZ92, Jac91]. LIL [Ein92, GK91, KLT94, Wu94a]. Lim [KLT94, MS93]. Limit [AMZ90, AR93, AG93, Arc94, BY93, BG99, BD93a, BP98, Ber92a, BP93b, CPU94, CS93a, CKS93, Che91, Cog91, CMS99, CS93b, Dar95, DK91b, DK91a, DKZ94b, FR93, FHU91, GMdB99, GH90, Gou93, GL90, Han89, Han91, He93, Hen97, HH97, Jak93, Kee92, KM94a, KP91, LS92, LS98, LT90, Led95, Lee90, Lee91, Lee93, Maj99, Oel90, Pia98, PP93, Qua95, Ros92b, Sch93b, Sim95, Tan93, Tot95, Wen90, dlC90, AGM96, Gen96, GMdB03, GT98, Jan98, Jan02, Kön96, KdHvdH97, Lan96, PU97, Ren96, Rez98, SX96, Tót96]. limite [Pia98]. Limited [Hou95b, LBG92, Law95]. Limiting [DZ95, HS90, MW90, Ale97, BS98, DE96, Sep98b]. limiting-shape [Ale97]. Limits [Blo94, CGST93, DJ94, HW92, KM94a, Lu95b, LS99b, Blo96, CS96, Dav96, HS98]. Lindvall [O’C95]. Line [GN99, Lee95]. Linear [AI99, AL95, Blo91, Hal91, Hou91, Mat91, Sto90a, DK97a, MS97, PU97]. Lines [Cha91, Ale96]. Liouville [Sta90]. Lipschitz [BH91]. Little [Szc92]. Local [AMZ90, AFL91, AR93, BBK94, BK95a, Ber91b, Ber92a, BP93b, CS93a, DFLM95, DM94, EK93, FG92a, GH90, Gri99a, Gri99b, HK95, Jac93, Kho93, Kho94, Kro93, Lac90, Lac97, Lan93, Leu95, MR92b, MR92a, MR94, MR99, Mar99, Pit99b, RW91, Ros90a, Str90, DE96, Deh97, Pin96, SS98]. Locales [Str90]. Localization [Bol94, BdH97, FG92b]. Locally [CG94, Hus90, Raj93, Raj94, Ste99, Wan95b]. Locaux [Leu95]. Log [BD93a, Bob99a, LPP95, Wan99, Bob96]. Log-Concave [Bob99a, Bob96]. Log-Sobolev [Wan99]. Logarithm [BK95a, BP93a, BS95, Che93a, Che94b, DM90b, DM92, DM94, DL94, Ein93, EM94, Ein95, Fis92b, Gan93, GG99, GL90, Hae93, Hal91, KH94, LRW92, MR94, Mas94, Rem94, Rio95a, Xia94b, HS96, Kes97, Tou90]. Logarithme [Tou90]. Logarithmic [AST95, Fan99, LY98]. Loi [Tou90]. lois [dF97]. Long [Dal99, DP98, Mue97]. Long-Range [Dal99]. Long-time [DP98]. Longest [Pen99]. Look [ZZ92, Tal96b]. Loop [Fan99]. loops [Enc98]. Low [BCG93, PP99, HM98, NW97]. low-dimensional [NW97]. Lower [Alb94, Ale90, BGZ94, Jai90, Kla90, Ste91a, Lif97, NdA98, Wan97, dA97]. Lowest [SS99, SS97]. Lueker [Jer92]. Lyapounov [Szn96, Zer98]. Lyapunov [CMP98, DW94, MS97].

M [Dur93, Num98]. M. [Glo90, PR91, PR92]. Machlup [SZ92]. Magnitude [dlPK94, KN97]. main [Ein96]. Majorization [Ber91a]. Majorizing [Tal96a]. Making [KL92]. Malliavin [Oco90]. Malyshev [Num98]. Manifold [Kat91, LZ95, MS99, GK98]. many [BMR97]. Mapping [Lal92].

Mappings [Han89, Han91]. marche [Pia98]. Marcinkiewicz [Rio95b]. Marginals [GKRS91, KK94, ST91, Ska93]. Mark [Wat96]. Markov [BL97, Glo90, Num98, Pol92, AMZ90, AFN95, AP91, BP94b, BPP95, BK93, BB96, BL88, BDT95, Che99, Cog90, Cog91, DW92, DK91a, Don91, DMT95, DEW91, EK93, EF93, Elo90, Erh99, Fit90, FK92, GM90, Glo91, HP90, HSC93, Hij90, Hog91, IM94a, IM94b, Jai90, KK95, Lia92, MR92b, MR96a, Mel92, MN90, NdA98, Ore91, Sch93b, Sep94, She91b, Shi92a, Tou90, VM94, VA94, Wih97, Xia94a, dA97, dTB91]. Markovian [CH94, FR93]. Martin [BPP95, PW94, Pin93, Roc92]. Martingale [BK93, Gou93, Hit90, Ken95, Leu95, Sto90b, Sto90a, Stu94, Wan91a, CK97, FK97, Ren96, Sch96]. Martingales [Big92, Cha91, Dar95, Dur93, Fis92b, Myk93, Pin94, Str90, Wan95b, Wan95a, dlP99, CKS98a, Gra97, Pin99]. Matching [DKZ94a, Shi92c, SY91, Shi97]. mathematical [CKS98a, KS96]. Mathematics [Sha98]. Matrices [Bai93a, Bai93b, BS99, CL94, CN90, Kal95, Muk94, Sil90, Sto95c, dMK99, BS98, Hen97]. Matrix [BY93, Kal94]. Maxima [CGZ95, Jak93, KZ94, Leu95, Sun93]. Maximal [CS90, DKZ94b, Rio95b, Sha95]. maximality [Pes98]. Maximum [Ath92, DM99, Geo93, Hsi95, Lac90, DP96, Pes98]. Maximums [Leu95]. McKean [Cha94]. Mean [Fen94, FG92b, LPP95, MS95, PP95b, Sep93, AGM96, Car97, Uch98]. Mean-Field [Fen94]. Mean-Variance [MS95]. Means [BLBE91]. Measurable [MPS91]. Measure [BKK95, Che93b, CN93, DLM95, DK95, DK99, EP99, Har99, Kal99b, Lal92, Le 91, LP95, Lin91a, SS99, Tal94b, Tal95b, Wu94b, BMR97, Dem97, DK96, DFST96, FT96, HL98b, Kön96, Mar96, Pin96, SS97, SSZ98, Sch96]. Measure-Valued [BKK95, DLM95, DK95, DK99, EP99, Le 91, DK96, Pin96]. Measures [Als92, BL92b, BK93, Bob99a, BET99, CH94, GZ90, Gri95, GKRS91, Jur92, KK94, Lal99, LO99, LL99, Mik91, Mue93, OS99, PF91, Pem92a, Raj93, Raj94, Rei95, ST94, Shi90, SY91, Ska93, Szu91, Zab92, dS92, BB96, BH97, BET02, BCG97, Fel96, FP97, Kle98, Lal02, LSS97, PW98, Sep98a, She96, Tal96a, Voi97, dA97, vdB97]. Mechanism [Del99]. media [LT98]. Menshikov [Num98]. Meromorphic [BLBE91]. Method [BCL92, BB92a, Erh99, HJ90, Rei95, JP98, Mar96, Szn97, O’C95]. Methods [KS91, LT90]. Metric [LL99, Van93, Stu98, dF97]. Metrics [Tue93]. métrique [dF97]. Meyer [Kur91]. Meyn [Wih97]. Michel [Hah94, Hah94]. Microscopic [FKS91]. Migration [Neu90]. Min [Pen92, Wei91]. Min-Stable [Wei91]. Minima [Alb96]. Minimal [Ale95, KZ94, Pen99, Pen96]. Minimax [HK91, SY91]. Minimax-Optimal [HK91]. minimization [FG97]. Miscellaneous [Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90g, Ano90h, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano92a, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h]. Mises [FM93]. Mixing [Bra93, Che91, Hsi95, Rio95a, Sha95, Yu94, BP98, vdB97]. Mixtures [Woo92]. Model [Blo94, COP99, Che92, Che94a, Der99, Lu94, PWM90, PWM92, Blo96, BG96c, BCD98, DP98, EF96, Gen96, KdHvdH97, Sep98b]. Models [BB99, Blo91, DK99, Fon93, Han99, Lig94, Neu90, Pem90, PS99, BCD96, GM97, GS98, Kle97, Kle98, LY98, Lig97, Lig99]. Moderate [IKL92, Wu94a, Wu95, dA97, Gra97, Wu99]. Moderately [IL93]. Modern [Sha98]. Modes [Sat93]. Moduli [CC92b]. Moment [CKS93, Kho92, KS91, Slu93, Zha99]. Moments [AL99, LdlP97, LZ99, MWZ93, Mee90, SZ92, AIM96, Lat97]. Money [KL92]. Monotonicity [ST93]. morphisms [Duh98]. Mosco [LO98]. most [Leu97]. Motion [AS95, AL95, BP94a, BH91, Bis92, BH94, Bra92, Bur90, BT95, Cha91, CMS99, DeB99, DD99, GJW99, Kes93, Lal92, LW99, LP95, Lee93, LR95, LZ95, OS99, Per90, Pin90a, Pin95, Rob91, RY91, Tal95b, Ver95, ABP98, BD98a, BW96b, Car97, DeB96, DR98, DL98, DFST96, FT96, GK98, Ham97b, HL98a, Lac97, Leu97, Pov97, Shi98a, TZ97, Wüt98, Dur93]. Motions [AFL91, DW94, Gri99a, Lee91, MSV90]. mouvement [Leu97]. Moving [AT92b, DR91, Hsi99, CS96, HH97]. MR [BL97, Lig99, Pin99, Wu99]. Multi [HV92]. Multi-Sample [HV92]. Multiclass [HW90]. Multidimensional [GM95, LRW92, Sow94, Zer98]. Multidimensionally [CS90]. Multilinear [dlPMSS94, Gad98, RS96]. Multiparameter [Nai90, Tal95b]. Multiple [Kal91, Kle97, Lig96, Per90, Pit99a, Shi92a, Shi98b, Szu91, Ham97a]. Multiplicative [dlC90, Ald97]. Multiplicity [BBK94]. Multitype [Ath92]. Multivalent [Cép98]. Multivariate [CH90, Got91, Hou91, dlPMS95]. multivoque [Cép98]. mutually [DP98]. MWI [Slu94].

N [Ros94]. Natural [DK97a, Sto95b]. Nature [DM95]. Navier [Bus99]. Near [Tri92, BdH97]. Nearest [CL92, CD90, HM98, Lig91, Mou92, Sch92a, Yag90, Yag91, Hof98, MS98b]. Nearest-Neighbor [CD90, Yag90, Yag91, HM98]. Nearest-Particle [Sch92a]. nearly [FS97]. Necessary [DL94, EK93, KL92, MS92, Tal92]. Negative [Led95, dF97, dF97]. Neighbor [CD90, Yag90, Yag91, HM98, Hof98]. Nested [Bun97]. Net [Adl90]. Networks [Bac92, CM91, IM94b]. No [BS98, BW96b, BLPS99, BL97, Lig99, Pin99, Wu99]. Noise [Day92, Mue93, Sow94, Enc98, Mue96, Mue97, Myt98]. nombres [dF97]. Non [BG99, CH94, DM99, DKZ94b, FR93, Ken95, Sow92]. Non-Aligned [DKZ94b]. Non-Gaussian [BG99, DM99, Sow92]. Non-Implosion [Ken95]. Non-Markovian [CH94]. Non-Poisson [FR93]. Nonabsorption [CMS95]. Nonamenable [BLPS99]. noncompact [Ata98]. Nonconvergence [Pem90]. Nondegenerate [BDT95, She91b, Sow95, dL95]. Nondeterminism [Ber91b]. Nondifferentiability [DM96]. Nonexistence [SvdB99]. Nonexplosions [KK95]. Nongradient [Lu94, Qua95]. Nonhomogeneous [AP91]. Nonincrease [Bur90]. Nonintersection [BL90]. nonisomorphic [Fel96]. Nonlinear [AP91, Arc94, BKK95, Blo94, Dyn92, DK97b, Enc93a, Enc93b, HS90, Hu91, Mel92, Ove96, Woo90, Ata98, Blo96, BM96, Fla96, FW98]. Nonnegative [BCL92, BD94a, CN90, Zha99, AIM96, KN97]. Nonnormal [HW92]. Nonparametric [Hal91]. Nonpolar [Eva92]. nonquasilocality [FP97]. Nonreversible [BDT95]. Nonsmooth [DI93]. Nonstandard [DM90b, Hae93]. Nonstationary [Hou95a, Jak93]. Nonuniform [Che91, DL94, JW99]. Norm [BMSS95, Leu95, dMK99]. Normal [CKS93, Hug92, Mee90, Slu93, GGM97]. Normality [MS92]. Normalization [Sep93]. Normalizations [GK91]. Normalized [HKW90, HW92, JW99, Ber98, HS96, Sha97]. Norme [Leu95]. Normed [KS91]. norming [BK97]. Note [Bra92, Din91, Hil92, Hor90, Hsi95, Jan02, Maj90, SZ92, Szu90, Zab92, Zha94, Ren96]. Number [Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90g, Ano90h, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano92a, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, BS90, Far90, HJ90, Sat93, Slu94]. Numbering [BL99]. Numbers [AI99, CL93, CGZ95, BP96, CM96, Gad98, GH92, Rei95, SZ99, Zha96a, dF97]. Numérotation [BL99].

Oblique [DI93]. observation [HS96]. Observations [AP91, SC92]. Observed [Hij90]. obstacle [EKP^{+}97]. Obstacles [Ant95, Tan93, Szn97]. Occupation [BB96, DR98, Imk93, IN94, Lan92, PW98]. off [Por96]. Often [Che99]. One [BD94b, BN94, COP99, CMS95, FGL94, Fis92a, FHU91, GH92, HV92, LS92, Lig92, Lu94, Mas94, Neu90, San93, Tal94c, Yag90, Yag91, AGM96, CS96, Kön96, Lan96, NV96, Pov97, RS96]. One- [HV92]. One-Dimensional [BN94, COP99, CMS95, FGL94, Fis92a, FHU91, LS92, Lig92, Lu94, San93, Yag90, Yag91, CS96, Kön96, Lan96, NV96, Pov97]. One-Sided [BD94b, GH92, Mas94]. Ones [Hit94]. Onsager [SZ92]. Onto [Kat91]. Operator [Jur92, Roc92, FS97]. Operators [SL95, LS97, Tak97]. Optimal [BG96b, BG99, CD95, Hal90b, HK91, Jac91, Jac93, KK91, MSV90, Nua92, Pes98, Sch96]. optimality [BB96]. Optimum [Pin94, Pin99]. Options [BS90]. Order [GH94, HP93, Jan94, KM92, KN97, Maj99, Mat91, dlPK94, Fla96, FS97, Jan98, Jan02, RdH96]. Order-of-Magnitude [dlPK94]. Ordered [Fit90, Fon93, Shi90]. Orders [AR93]. Organizing [CGST93]. Oriented [NY93]. Orlicz [NS94]. Ornstein [Fer92, IMM^{+}90, Nor98, Roc92]. Orthogonal [MvW91]. Orthogonality [MvW91]. oscillators [FW98]. other [PS99]. overlap [Gen96]. Overshoot [GM95].

P [Wih97]. P.D.E. [DK97b]. Packing [Jer92, PT96]. pages [Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90g, Ano90h, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano92a, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h]. Pair [MSV90]. Palm [Kal99b]. Parabolic [BMSS95, DD99, Dyn92, Roc92, DK97b, Fla96]. Paradox [KK94]. Parameter [KH94, Pin90a, Sch93b, PY97]. parameters [DFL98]. Part [Ken95, LZ95, Bai93a, Bai93b]. Partial [BMSS95, DH95, DD99, DK91b, DK91a, Dyn93, Hsi99, KZ94, Mat91, Rio93a, Rio93b, Sha95, EM96, HK97, HS96, KX96, Qia98]. Partial-Sum [Rio93b]. Partially [Hij90]. Particle [CGST93, CL92, DPPV94, Der99, DK99, DN94, Dyn91, Fen94, GW91, Gri93, Gri99b, Lan91, Lan93, Lig91, Liu95, Mou92, Sch92a, Sim95, Sud90, SL95, CS96, GM97, LS97, MS98b, SX96]. Particles [FR93, Kom96]. Particular [Har99]. Partitioning [Jer92]. Partitions [PWM90, PWM92]. Parts [Fan99, IN94]. Passage [Als92, AIM96, HM95, LW94, LP92, GG96, LN96, LP03, Sep98b, WW98]. Passage-time [AIM96]. Path [Bol94, Del99, Fan99, MR92b, Nol88, Nol92, MR96a]. Paths [ALP99, BL90, MPS91, NS94, RS93, Shi92a, BPP98]. Pathwise [Enc93b]. Patterns [DPPV94]. Payoff [MS93]. PDE [DP97, EF93, EKP^{+}97, Zhe95]. Pemantle [Ben97]. Percolation [Ale90, Ale95, And93, BA91, BK95b, BLPS99, BG91, CC99, DG99, HM95, KZ90, Kle94, Lyo90, Lyo92, LP92, LS99a, NW97, NY93, Roy90, Sch92b, VM94, Zha94, AP96, BPP98, GS98, Häg97, HM98, LN96, LP03, Sep98b, WW98, Bez92]. Period [GP92]. Periodicity [Cog90]. Periodogram [DM99, TW90]. Permutations [AT92a]. Persistence [GW91]. Persistent [NV96]. Perspective [Sha98]. Perturbations [BN94, Iof93, Sow92, Sow94, dS92, FW98]. Perturbed [FW92, Mik90, Pin90a, Dav96]. Petri [Bac92]. Phase [Asm92, BD93b, DS95, Fon93, PS99, Sta96]. Phase-Type [Asm92]. Phases [DPPV94]. Phenomena [DKZ94a]. phenomenon [Por96]. Philosophy [Sha98]. Photograph [Ano93i]. Physical [Cos91]. Physics [Sha98]. Pinned [She91a]. Pinsky [Bas97, Wat96]. Pioneering [Gri93]. Pitman [Ber92b]. plan [Gol98]. Planar [Bra92, LW99, BS96, BW96b, Ver96]. Plane [MN90, Nua92, Pin87, Pin90b, Rev93, RY91, Ros92a, Ker96, Sep98b, Shi98a, WW98]. Plato [Sha98]. plus [Leu97]. Point [BB92a, BL92b, BL99, BF90, DH95, DFLM95, MN90, Nai90, Ste94, BW96b, Ham97a, Leu97]. Points [Bár99, BBK94, KM94a, Mar99, Nua92, Pem90, Per90, Shi92a, Ver95, Lig96, Shi98b, dF97]. Pointwise [Hou92, dF97]. Poisson [AT93, AT93, AST95, BCL92, BH94, BL99, CC92a, Erh99, FR93, GM96, HJ90, Hor90, HPA95, Kho93, LU95a, PY97, Wan91b]. Poissonian [Pov97, Wüt98]. Polar [Eva92, PP95b, Per90]. Pólya [Gou93]. polymer [BdH97, CM96, Kön96]. Polymerization [PWM90, PWM92]. Polynomial [Far90, Slu93]. Polynomials [Maj99]. Ponctuel [BL99, dF97]. Population [Blo99, Che92, Che94a, Don91, DJ92, DK99, Blo96]. Populations [KW90, Oel90]. Portfolio [Alg92]. Position [Bár99, GM92, GM94, Har96]. Positive [Ber92b, BF90, GJW99, Kee92, O’B99, Str90, Hen97, Bas97]. Positives [Str90]. possible [CK97]. Potential [BH91, CZ96, FU96, Car97, FS97, Pov97, Wüt98]. Potentials [GH90]. Pour [Tou90, Pia98, dF97]. precision [HK97]. predictability [HM98]. Prediction [Alg92, Alg95, Hou91]. Preferences [Pit93]. Prefixes [Shi92b]. Prescribed [Dar95]. Preserving [MvW91]. pricing [Sch96]. primitives [Lac97]. principal [FS97, Szn97]. Principle [BET99, Bry92, FK99, FK02, FW93, Geo93, HP93, PP99, Ver99, BET02, BG96c, Pes98, Yu96]. Priority [PR92, She91a, Yag91]. Prises [Str90]. Probab [Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90g, Ano90h, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano92a, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, BL97, Lig99, Pin99, Wu99]. Probabilistic [Bas97, Bus99, Der99, Sta90, Van95, Hsu98, Jer92]. Probabilities [Cog90, FJ92, Ore91, RHHQ90, Sun93, dlPMS95, Col96, GS98, Lif97]. Probability [Bal95, Bár99, Bob99a, Bra96, EH92, Fin92, GKRS91, HK95, He93, Hog90, Jur92, KZ90, MWZ93, PF91, Raj93, Raj94, ST91, Sha98, Ska93, SL95, Zab92, BMR97, BH97, LS97, Qia96, AT93, Hah94, dA96]. Problem [Dal90, Ein95, Hog91, PP94, Pin90a, Pit93, Roc93, Slu93, Tal94d, Cép98, KLS96]. Problème [Cép98]. Problems [Asm92, BK93, BS90, HK91, Mik91, NP91, Sto90b, Sto90a, EKP^{+}97, Lot96]. Procedure [Win90]. Process [Adl90, AFL91, Alb94, AP91, AST95, Ath92, BK95a, Ber92a, BG90, BL99, DH95, DeB90, DeB99, DS95, Ein92, EK93, EG93, Eva92, Fen94, FKS91, FF94, FGL94, Fit90, FG92a, Gri99b, Gri95, Lal99, Lan92, LU95a, Lig95, MSS94, Maj90, MP91, Mou92, Nob92, OW90, Pem92b, PW93, SS99, Sep99, She91a, She91b, Sho91, Slu94, Sud90, Wu94b, Yag90, Yag91, dL95, CKS98b, LS98, Lal02, LL98b, Lig96, Pes98, PT96, Qia98, SS97, SS98, Sta96, Zha96b]. Processes [AMZ90, AR93, AS95, Alb90, AK92, AG93, Arn94, Bal91, BB92a, Ber91b, Ber92b, BG91, BK93, BL88, BDT95, BET99, BP92, BDS91, BCG93, BF90, BGM93, Cos91, CC92b, CH90, CS93b, Dal99, DLM95, De 93, DM90a, DM90b, DM92, DM94, Deu94, DK95, Don91, DJ92, DMT95, DEW91, EF93, Elo90, FK92, FW93, FHU91, GM90, Glo91, HP90, Hij90, Hil93, HS90, Hor90, Hou91, Hou95b, Hus90, Imk93, IL93, IMM^{+}90, IM94a, IM94b, Jai90, JLQY99, Kal99a, KK95, Kho93, Kle94, KH94, Lac90, Le 91, Lee90, Lia92, Lig92, Liu95, LZ94, LPP95, MR92b, MR92a, MR94, MR99, MN90, Mog93, Nai90, NS94, O’B99, ORS95, Pen92, Rio93a, Rio93b, RS93, Ros95a, Sch93a, Sch93b, Shi92a, Ste91a, Ste91b, Stu94, Tal90a, Tal90b, Tal92]. Processes [Tal93, Tal94a, Tow95, VA94, Wei91, Wu94a, Xia94a, Yu94, AN98, Alb96, Alb98, AIM96, Ata98, BB96, BL97, BET02, BCG97, BM96, DE96, Deh97, DR98, DP96, Don96, EK98, FG97, Hof98, HL98b, Jan98, Jan02, Lan96, LOY96, LL98b, LL98a, MR96a, MR96b, MP96, NV96, NW97, Nor98, PU97, Pel98, PY98, Qia98, Sch97, SY96, Shi98b, Stu98, Tou90, Vit96, Glo90, Hah94, Pol92]. Processus [BL99, Tou90]. Produce [Sat93]. producing [Fel96]. Product [Cha91, GL90, dS92, BH97, LSS97, Tak97]. Product-Limit [GL90]. Products [CN90, CGZ95, Muk94, Hen97, Zha96a]. Profile [Lal99, Pit99b, HM98, Lal02]. Programming [Sto90a]. Projection [Sch93b]. projections [BMR97, FP97, RS97]. Proof [Kri94, PR91, PR92, Sta90, Bob97, Ein96, SS98]. Proofs [LPP95]. Propagation [Zhe95, LT98, Pra98]. Properties [Bor93, BD94b, BF90, Del99, EP99, Hal90a, He93, MR92b, MN90, Nol88, Nol92, RR91, Wei91, dL95, Bob96, LOY96, MR96a, Pin96, TZ97]. Property [AFN95, Che93b, DW92, EK93, PZ95, Qia96]. Prophet [SC92, CK97]. prove [Mar96]. Publications [Ano93j].

quadrant [AIM96]. Quadratic [CGZ95, GT99, dlPK94, BG96b, GT98]. Quantile [DM90b, DM92, Xia94b, Deh97]. quantitative [KS96]. Quantum [LS97, SL95]. Quasi [BS90, FKMP95, LS97, dS92]. Quasi-duality [LS97]. Quasi-Invariance [dS92]. Quasi-Stationarity [BS90]. Quasi-Stationary [FKMP95]. Quasilinear [Zhe95]. Quasireversibility [HW90]. Quenched [PP99, Zer98]. Question [Bár99]. Queue [GP92]. Queueing [Asm92, BL92a, DIS90]. Queues [SK90]. Qun [Pol92].

R [Wih97]. Rademacher [DMS93, ES93]. Radial [Ken95, LZ95]. Radially [Iof93]. Ramer [ZZ92]. Random [AL99, Ald91, Ald93, Ald94, AK92, AP91, ALP99, Ant95, AT92a, Asm92, Bai93a, Bai93b, BCGH93, BCL92, BS96, Ber91a, BD94a, Big92, Boi93, Bol94, BET99, BP93a, Bra93, Bra96, BDS91, BN94, BD98b, Bry92, BL90, CD95, COP99, CL94, Che91, CL92, Che93a, Che93b, Cog90, Cog91, CGZ95, DJ94, DR91, DH95, DLM95, DSC93, DZ92, Ein91, Far90, FR93, FG92b, FW98, Geo93, GH90, GdH94, GM92, GM94, Gri95, GL90, Han89, Han91, HSC93, Hil92, Hil93, Hit94, HJ90, Hsi95, Hu91, Jak93, Kee92, KK91, Kes93, KM94a, KR91, Kla90, Kle94, Kur91, KS91, Lah93, LW94, Lal93, LR91, LT90, LRS90, Lee95, LRW92, Lig92, Lyo90, Lyo92, LP92, LPP96, LP03, MPS91, Maj99, MR94, MR96b, Mat91, Mee90, Muk94, Nai90, OS94]. Random [PP95a, PV99, PP93, PP99, PY98, Pit99b, PWM90, PWM92, Pit93, Pru90, Rev90, Rev93, Ros90a, Ros94, Ros90b, Ros95b, ST93, ST94, Sch94, Sep93, Sep94, Sil90, Slu93, SvdB99, Sto95b, Sto95a, Sto95c, Sun93, Szn95, SZ99, Szu90, Szu91, Tal90a, Tal92, Tal95a, Tan93, Tho92, Wat96, Wen90, Wic90, Wu95, dMK99, AD98, Ald97, Ale96, AIM96, BBE97, Ben97, BP96, BK97, BS97, BET02, Bun96, Car97, CP97, CMP98, DP97, Dav96, DH96, FK97, Fer90a, Fer92, FU96, Gan98, HK97, Ham97a, Ham97b, Hen97, HL98b, HS98, KN97, Kom96, Lat97, LY98, LT98, LN98, Lif97, Lot96, Myt96, NV96, Pen96, Pia98, Por96, Pri96, Roi96, Sep98a, Shi98a, Tho96, Tót96, Uch98, Wu99, Zer98, Zha97, dF97, vdB97]. Random-Cluster [Gri95, Sep98a]. Random-Size [FR93]. Randomization [Dal90]. Randomized [Nua92]. Randomly [FW92, Kla90, Mik90, LdlP97]. Range [Dal99, FGL94, GH90, GK98, JLQY99, Lal93, LR91, MP91, PV99, Zha93, Duh98, Ham97a, Lan96, LSV96, MP96, SX96]. Rank [HP90]. Rankable [BS90]. Rao [MW90]. Rapid [Nob92]. Rate [Bai93a, Bai93b, BGZ94, Din93, DK95, Got91, Hor90, KZ94, Pin95, Pru90, Sho91, Voi97, BP98]. Rates [Hal90b, Kho94, Lig91, Yu94, Ale97, BG96b, Ren96, RdH96]. Ratio [CMS99]. Rational [Lal92]. Ratios [KM92, dlP99]. Ray [Tót96]. Reaction [Blo91, Blo94, DN94, Fre91, FW92, Sow92, Sow94, Pra98]. Reaction-Diffusion [Blo94, DN94, Fre91, FW92, Sow92, Sow94, Pra98]. Real [GN99, Lat97]. Recall [Roc93]. Reconstruction [Hou95b]. Records [BL99, DZ95]. Recovery [Hal90b]. Recur [Che99]. Recurrence [Pin87, Pin90b, PP93, Pin96]. Recurrent [BL92b, Che99, Iof93, MR94, Ham97a, Tou90]. Recurrents [Tou90]. Recursive [Ald94, BL92a, Ham97b]. Refinements [GH92, Lah93]. Reflected [BT95, DeB99, EKP^{+}97, AIM96]. Reflecting [BH91, DW94]. Reflection [Cos91, DI93, CK96, DeB96]. reflections [Por96]. Regenerative [Ber99, Kal99b]. Regime [BD93b, PP99]. regimes [Kle96]. Regression [Hal91]. Regular [BL92b, Che93b, RdH96]. Regularité [Fer92, Fer90a]. Regularity [DD99, Tal93, Fer92, Fer90a]. Reinforced [PV99, Ben97]. Related [AS95, Hou95a, O’B99, Sch92b, Wan95a, Wu95, DR98, EKP^{+}97, Wu99]. Relative [OS99, Wen90]. Relaxation [JLQY99]. Remark [GZ94]. Remarks [Ros95b]. Renewal [Als92, Ber99, Dal99, FKMP95, GM95, Hu91, Mel92, Ste91b, Woo90]. Renormalisation [AR93]. Renormalization [RY91, Ros92b]. Renormalized [Ber92a]. Rényi [Gan98]. Representation [BL92b, Jak93, Slu94, dTB91, Ahn97, BD98a, DK96, Gne97]. Representations [Asm92, De 93, DK99, EP95, Glo91, Kal94, Ros90b, Tue93]. Reproduction [Che92, Che94a, Nob92]. Repulsion [Tot95]. Repulsive [Yag90, Yag91]. Respect [Szu91]. Result [TW90, Ham97a]. Results [CC92a, DS93, KLT94, Sch93a, Xia95, Ale97, EM96]. Reveals [OW90]. Reversible [Boi93, CGST93, CL92, Lig91, Sch92a, dTB91, MS98b, SX96]. Review [Bas97, Bez92, Bra96, Dur93, Glo90, Hah94, Har93, Hsu98, Jer92, Num98, O’C95, Pol92, Sha98, Wat96, Wih97, dA96]. Reviews [AT93]. Revisited [FM93, Szn97]. Revuz [Dur93]. Reward [KK91]. Richard [Hsu98]. Riemann [CP97]. Riemannian [GK98, LZ95]. riffle [Lal96]. Robust [PS99]. Robustness [BKK95]. Roommates [Pit93]. Rosenthal [Hit90]. Ross [Bas97]. Rotation [AI99]. Rotational [Kal94, Kal95, dTB91]. Rotations [Ros94]. Rough [Cat92]. Ruin [GM94, Hog90, Hog91]. Rule [BS90]. Rules [HK91]. Russo [Roy90, Tal94c].

S [AT93, Jer92, Wih97]. S. [Sta90]. S.-Y. [Sta90]. sa [Gol98]. Same [KK94, PP95b]. Sample [BY93, BS99, HV92, MR92b, NS94, RS93, Shi92a, Sil90, Tal90b, Tal92, Wei91, BS98, MR96a, Bai93b]. Samples [FR93, Hou95b, KR91, KW90, KLT94, Pri96]. Sampling [OW90, Sho91]. Sausage [Bol90, CFR91, Por90]. Sausages [Ros92a]. Savage [KS96]. scale [Kle97]. Scaled [KR91, Pri96]. Scaling [CC99, Lu95b, Qua95, PY98]. Schedules [Cat92]. scheme [PT97]. Schemes [Alg92, Alg95]. Schrödinger [FG97, Tak97]. Schweizer [MS95]. Score [DKZ94b]. Scoring [DKZ94a]. SDE [Dar95, DP97, EKP^{+}97, Kal96, Pit99b]. SDEs [DI93, Pra98, Ver99]. Second [RdH96, SS99, Fla96, SS97]. Second-order [RdH96, Fla96]. Secretary [HK91, Roc93]. Segmental [DK91a, DKZ94b]. Selection [Alg92, BS90, FG92b, ORS95]. Self [AS95, Ald94, Ber91b, Ber92a, BEI92, CGST93, GK91, Gri99a, HKW90, HW92, JW99, Ros92b, Sha97, Tot95, VA94, Alb98, BS97, Tót96]. self-attracting [BS97]. Self-Avoiding [BEI92, Tot95]. Self-Diffusion [Gri99a]. self-interacting [Tót96]. Self-Intersection [Ber92a]. Self-Intersections [Ber91b, Ros92b]. Self-Normalizations [GK91]. Self-Normalized [HKW90, HW92, JW99, Sha97]. Self-Organizing [CGST93]. Self-Similar [AS95, VA94, Alb98]. Self-Similarity [Ald94]. Semi [Pen92]. Semi-Min-Stable [Pen92]. semiexponential [Gan98]. Semigroups [He93, Jur92, Bun96]. Semilinear [Lee93, DP97]. Semimartingale [Ahn97, DW94]. Semimartingales [Jac93, Jak96, PW98]. Seneta [BK97]. sense [LO98]. Separate [DPPV94]. Separation [BS99]. Sequence [Arc94, Ath92, DM99, DKZ94a, DKZ94b, KK91, OS94, Shi90, Wen90]. Sequences [BL92a, CKS93, HP93, Jak93, Rio95a, Sha95, Yu94, BP98, CK97, DFST96, FT96, Pel98, SY96, Yu96]. Serial [HP90]. Series [BP93a, DMS93, FJ92, HP90, Hou95a, KS91, Ros90b, Ros95b, SK90, Tal90a, Tal92, Tal95a, HL96, Lif97, MR96b, Zha97]. Service [HW90]. Set [Ein95, Lal92, Pin93, Rio93a, Rio93b, LS98, LO98]. Set-Indexed [Rio93a, Rio93b]. Sets [ABP98, Ber99, DM94, Eva92, Kal99b, LO99, PP95b, Per90]. Several [MR99]. Sexual [Che92, Che94a, Nob92]. Seymour [Roy90]. Shape [NP95a, NP95b, Zha93, Ale97, Sep98b]. Shapes [HM95]. Sharp [DW92, Ham95, KK95, Wan91a, Wan95a, Wan97]. Sharpe [Glo90]. Sharper [Tal94a]. Sheet [Kho95, KS99, Tal94d, DM96]. Sheets [DW92]. Shift [Bob99b, CH94]. Shift-Invariant [CH94]. shifting [Tho96]. Shifts [KW92]. short [Ein96]. shuffles [Lal96]. Sided [BD94b, GH92, Mas94]. Sierpinski [Ham97b, HL98a]. Signal [Hal90b]. Signed [Adl90]. Similar [AS95, VA94, Alb98]. Similarity [Ald94]. Simple [Ant95, FKS91, FF94, Lan92, Sep99, HS98, Jan98, Jan02, LOY96]. simplified [Sep98b]. Simulated [Cat92]. Simulation [Bal95]. Sinai [HS98].

Singular [BB92b, Bob99b, CGST93, Mue96, Wic90]. Singularity [DFLM95]. site [GS98]. sites [EF96]. Size [Bob99b, CC99, FR93]. Skorohod [BT95, Cép98]. Skorohod-Type [BT95]. Skorokhod [Shi90]. Slepian [ST93, ST94]. Slepian-Type [ST93]. SLLN [Hou92, HL96]. Slow [Fon93, Mar99, Ver95]. Slowly [HV92]. Small [CS93b, Day92, Gan93, KH94, LL99, MWZ93, MP96, Neu90, Pin90a, RHHQ90, Ste91b, Tal94d, Har96]. Smallest [BY93]. Smooth [BM95, DP96]. Smoothness [MSS99]. snake [Ver96]. snakes [LL98a]. Sobolev [Fan99, LY98, Wan99]. Sojourn [Kal99a]. Solution [BKK95, DIS90, Gne94, KK94, Pin90a, Sto90b]. Solutions [Kat91, EKP^{+}97, GM96, Kal96]. Solved [Pit99b]. Some [AL95, BH91, BD93a, Bor93, BD94b, CEF91, CC92a, CK97, DK95, DMY97, EM96, Fin92, GK91, IL93, Iof91, KW92, LOY96, LT90, Lee92, Lee93, MvW91, Mee90, Rem94, Sch92b, She91b, Sow95, Wei91, LY98, Lif97, Lot96, Ren96]. Space [ARZ93, Che93a, Che94b, De 93, Ein91, Ein93, Ein95, Enc93a, Fan99, IN94, IS99, Kla90, MSS99, Nai90, Roc92, Ste91a, Blo96, Bob97, Lan96, MS98a, RS97, dF97]. Space-Time [Ste91a, Blo96]. Spaces [Enc93b, Hah94, HPA95, KS91, NS94, PZ95, Pin94, Raj93, Raj94, Bob96, Stu98, Pin99]. Spanning [Ale95, BP93b, Pem91, Pen99, Pen96]. Spatial [BCG93, BCD96, DPPV94, IS99, MS98a, BCG97, BCD98, DF98, Kle97, Kle98]. Spatializing [BET99, BET02]. Spatially [EP99]. SPDE [Fla96, Kry97, Sow95]. species [BCD96, BCD98]. specifications [FP97]. Spectral [Bai93a, Bai93b, Che93b, Hou92, HL96, LSV96, Mee91, BS98]. Spectrally [Ber92b, O’B99]. Speed [FHU91, PP99, Jan98, Jan02]. Sphere [CN93, GM92]. Spheres [BG94a, GM94, RR91, Voi97]. Spin [BG94b, Fer90b, Ste91a, AG97]. Spitzer [Ano93i, Ano93j, Ano93k, Gri93, Kes93]. splicing [PY98]. Square [Wan91a, Wan95b, BS96]. Squares [AMZ90, HKW90]. Stability [BB92b, BM96, Che92, Che94a, RR91, TW90, Ata98, Bun96, Wih97]. Stable [AR93, AL99, AT92b, BW96a, DJ94, DeB90, Lac90, LR91, LRS90, MR92a, MS99, Nol88, Nol92, NS94, O’B99, Pen92, PWM90, PWM92, Pit93, Por90, Ros92a, Ros95a, ST91, ST93, Szu90, Tal90a, Wei91, Xia95, AD98, Ber98, MP96, PY97, Shi98b]. standard [AP98, GGM97]. State [ARZ93, LU95a]. states [She96]. Station [HW90]. Stationarity [BP92, BS90, SK90]. Stationary [Alb90, Alb94, AK92, Arc94, Arn94, Ber92a, Bry92, Che93b, DF90, FKMP95, GP92, HM95, HP93, HS90, Hou95a, Hsi95, Hus90, Kal99a, MR92a, Ore91, Rio95a, Ros95a, Ste91a, Sto90b, Tal90a, Yu94, HL96, HL98b, Pel98]. Statistic [Blo99, Stu94, BG96a, GGM97]. Statistical [Mel92]. Statistics [BGZ94, BG99, BD94b, CdlPT93, DEW91, FR93, GZ94, GH94, HP93, HP90, HV92, Jan94, KM92, KW90, Maj94, MS92, OR93, Stu91, Zha99, dlP92, dlPMSS94, dlPMS95, KN97, LdlP97]. Stay [BD94a, Kee92]. Stein [BCL92, BB92a, Erh99, HJ90, Rei95]. stepping [EF96]. stepping-stone [EF96]. Steutel [Kri94]. Sticky [Der99]. Stirrings [Fer90b]. Stochastic [AFN95, Alb94, AI99, BL92a, Bac92, BMSS95, BB99, BM95, Blo91, Bur94, CK91, CL94, CM91, CC92b, EP95, Fla96, GM97, Gri95, GN99, HL98a, Ham95, Har93, Imk93, IS99, Kal95, Kat91, KH94, KP91, LN98, Lig97, LS99b, MS93, MNS92, MSS99, MS99, Neu90, NP91, PP90, Pem90, Rem94, RW91, RW97, ST93, Sch94, Szu91, Tal90b, Tow95, Tue93, Wih97, AIM96, CK96, CKS98b, DF98, JP98, Jak96, KX96, Lia97, MS97, MS98a, Pia98, Pit98, PT97, RS97, Har93]. Stochastically [Ore91]. Stokes [Bus99]. stone [EF96]. Stop [HK91]. Stopped [Kla90, LW94, Lee90, LdlP97]. Stopping [Cha91, Fit90, Jac91, Jac93, KK91, KM94b, Nua92, Shi90, Pes98]. Strassen [Che94b, DL94, Gan93]. Stratonovich [San93]. Strict [BP92]. String [Shi92c, Shi97]. strings [CMP98]. Strong [Bur94, CS93a, Che93b, CS93b, DK91b, DK91a, DF90, HP93, IKL92, LT90, PZ95, Rio93a, Rio93b, Ros95b, Ste91b, Wan95a, Zha97, DE96, Yu96, CM96, Deh97, Gad98, GH92, Pen99, Rio95b, Zha96a, Zha99, dF97]. Strongly [Hsi95, MR92b, Rio95a, Tal90a, Kom96]. Stroock [dA96]. Structure [AT92a, FKS91, Ros95a, Lal96]. Structured [Oel90, EF96]. structures [Gne97]. Student [BG96a, Blo99, GGM97]. Students [Ano93k]. Studies [Sha98, MS97]. Study [VM94]. Sub [Zha99, LO98]. sub- [LO98]. Sub-Bernoulli [Zha99]. Subadditive [RS93, Sch91, Ale97]. Subcritical [BG91]. Subdiffusive [Law95]. Submartingale [CS90, Ham95]. Subordination [Bur94, Wan95a]. subordinator [PY97]. successive [Lac97]. Sufficient [DL94, EK93, MS92, Tal92]. Sum [DH95, Hsi95, KZ94, Rio93a, Rio93b, Qia98]. Sums [AT92b, CHM91, CM94, CGZ95, DK91b, DK91a, Ein91, GH90, Hae93, HKW90, HW92, Hil92, Hit94, Hsi99, Jan94, JW99, KM92, KM94a, Kla90, Lah93, LT90, Sha95, Ber98, CK97, EM96, HK97, HS96, Lat97, Zha96a]. Sup [MS93]. Super [AS95, DFLM95, DD99, LP95, Lee93, LR95, Per90, Pin95, DR98, DL98]. Super- [Pin95]. Super-Brownian [DFLM95, DD99, LP95, Lee93, LR95, Per90, DR98, DL98]. Supercritical [KZ90, Lal99, Pin95, PWM90, PWM92, AP96, LS98, Lal02, Pin96]. Superdiffusions [Dyn92, Kro93, DK97b, Kry97, She96]. Superdiffusive [Wüt98]. supermartingales [LO98]. Superprocesses [Del99, Dyn91, Dyn93, Myt96, Ros92b, Tri92, DFL98, DK97a, LL98a, Ove96]. Support [BMSS95, DLM95, EP99, LP95, Pin95, Ver95, BS98, Pin96]. Supports [Raj93, Raj94]. Sure [Ber91a, Ein91, HV92, Hou95a, KZ94, OS90, Sep93, BG96c]. Surely [Tal90a]. Surfaces [CFR91]. Survival [BH94, BN94, Lig92, Szn93, NV96, SS98]. Switching [CD95, MSV90]. Sylvester [Bár99]. Symmetric [ARZ93, AG97, DeB90, FG92a, HV92, Iof93, Lan92, LO99, Liu95, Maj99, MR92b, MR92a, MR94, Szu91, Fit97, Jan98, Jan02, MR96a]. Symmetries [GM90]. Symmetrization [RS96]. symmetrized [Zha99]. Symmetry [Glo91, Lia92, dL95]. System [DPPV94, Fen94, Gri99b, Qua95, Sud90, MS98b]. Systems [AFL91, BG94b, CGST93, CL92, CG94, DG99, DIS90, DS93, DN94, Dyn91, EF93, FR93, Fer90b, GW91, Gri93, Lan91, Lan93, Lig91, Mik90, Sch92a, Sim95, Ste91a, Ste99, SL95, CS96, CZ96, Kle96, Lig97, LS97, Pra98, SX96].

Tagged [Gri99b]. Tail [CH90, DM90b, Ein92, Sun93, dlPMS95, Lif97]. Tailed [Hal90a]. Taken [Str90]. Talagrand [Hah94]. Tanaka [RY91]. Taylor [KLS96]. Techniques [DSC93, Hsu98]. Temporal [PP93]. Temps [Leu95]. Tends [Maj99]. term [Mue96, Mue97]. terminal [DP97]. test [DL98, Uch98]. Their [He93, Maj99, RR91, CKS98a, DMY97, Zha97]. Théoràme [Pia98]. Theorem [AL93, BMSS95, Ber92b, CKS93, ES93, GH92, Han99, Hou92, Ken95, MS99, Pia98, Roy90, Sch91, Sim95, Wan91b, Woo90, Zha93, AGM96, Ein96, Gen96, KS96, Kön96, KdHvdH97, PU97, SX96, Zha96b, dF97, Ber92a, CPU94, Che91, Cog91, Han89, Han91, Led95, PP93, Sta90, Tan93]. Théorème [dF97]. Theorems [AG93, Arc94, BG99, BD93a, BP93b, CS93a, CMS99, CG94, CS93b, DK91b, DK91a, FR93, GH90, IKL92, Iof91, Jak93, Kal99a, Kee92, KP91, LS92, LT90, Lee90, Lee92, Lee93, Mik91, Oel90, Ros92b, Sch93b, Tot95, Wen90, BP98, Ber98, EK98, GMdB03, GT98, Hen97, HH97, Ren96, Tót96, Zha97, AMZ90, GMdB99, Gou93]. Theory [Alb90, BL92a, Bac92, BH91, Ber99, Glo90, GM95, He93, Hu91, Kif90, Mel92, Num98, Pol92, Sha98, SL95, VA94, dA96, Alb98, CZ96, LS97, RdH96, Tót96]. thinnings [LS97]. Three [Ale90, CC99, Fis92a, LR95]. Three-Color [Fis92a]. Three-Dimensional [CC99, LR95]. Threshold [DS93, Han99, Lig94, Ste94, GG96, Lig99]. Tight [Mas90]. Tightness [Muk94]. Tilings [BP93b, BS96]. Time [BBK94, Ber92a, DFLM95, DeB90, EK93, FGL94, FG92a, FJ92, Gan93, GW91, GM92, GP92, HP90, Hog90, Kho94, Kif90, Kur91, Lac90, Lan92, Pin95, RW91, Ros90a, Ste91a, Sto90b, Sto90a, Wan95a, AIM96, Blo96, DP97, DP98, DR98, GT98, Har96, Mue97]. Time-Average [Sto90b, Sto90a]. time-domain [GT98]. Times [AMZ90, AFL91, AR93, Als92, AL95, BK95a, DF90, Fit90, IM94a, IM94b, Jac93, Kho93, KM94b, Kro93, LW94, Leu95, MR92b, MR92a, MR94, MR99, Mar99, Pit99b, Rob91, GG96]. Tomography [KK94]. Topics [Num98]. Topologies [DL94, Jak96]. Torgny [O’C95]. torus [MR96b]. Total [AST95]. Totally [Sep99, Lan96]. Traffic [SK90, Szc92]. Trajectories [Tal95b]. Transfer [BP93b]. Transfer-Impedances [BP93b]. Transform [ARZ93, FG92a]. Transformations [Enc93a]. Transforming [Tho96]. Transience [Pin87, Pin90b, PP93, Pin96]. Transience/Recurrence [PP93]. Transient [CL94, Iof93, Tho92]. Transition [BD93b, Cog90, EG93, Kal94, Ore91, She91b, BdH97, Lig96]. Transitions [PS99, Sep94]. Translate [OS94]. Translation [Glo91, Zab92, Sep98a]. translation-invariant [Sep98a]. Transport [Cos91]. Transportation [Tal94b]. Traps [BH94, Szn93]. Travelling [FKS91]. Tree [Ald91, Ald93, Arn94, DS95, Eva92, Lal99, Pem91, Pen99, PW94, Pit99b, PWM90, PWM92, LS98, Lal02, Lig96, Pen96]. Tree-Type [PWM90, PWM92]. Trees [Ald94, BP94b, BP93b, Lal93, Le 91, Lyo90, Lyo92, LP92, Pem92a, Pem92b, PP94, PP95a, PP95b, PS99, Häg97, LP03, SS98, Sta96, Zha96b]. Triangle [BA91, NY93]. Triangular [Jan94, Ros95b, Sto95c]. Triangulations [Ald94]. Trigonometric [Far90]. Trimmed [Hae93, HW92, KM92, KM94a]. Triple [RY91, BW96b]. Trotter [Tak97]. True [GMdB99, Tot95, GMdB03]. Truncated [CGZ95]. Truncation [GL90]. Turbulence [FK99, FK02]. turbulent [Kom96]. Tweedie [Wih97]. Two [Ale90, And93, Bol94, Bus99, CD95, Dal90, DeB90, DKZ94b, LP95, Liu95, MSS99, NP95a, NP95b, Oco90, Wu94b, Ale96, DF98, FU96, Ham97a, LN96, PY97, Wüt98]. Two-Armed [Dal90]. Two-Dimensional [And93, Bol94, Bus99, DeB90, LP95, Ale96, FU96, Ham97a, LN96, Wüt98]. two-parameter [PY97]. Two-Particle [Liu95]. Two-Sequence [DKZ94b]. Type [Asm92, Bob94, BT95, DM90a, EM94, PWM90, PWM92, ST93, Sow95, Kes97]. Typical [COP99].

Uhlenbeck [Fer92, IMM^{+}90, Nor98, Roc92]. Unbounded [Fen94, FK97]. Uniform [Big92, Che91, CH90, DMT95, Dud92, GZ91, HK95, Kla90, Maj90, Mou92, SW92, Tal94b]. Uniformly [Pem91]. Unimodal [Sat93]. Uniqueness [BKK95, CEF91, Gri99b, Gri95, DeB96, Myt98]. Universal [Alg92, Alg95, EM94, Mas94, OS90, EM96, Kal96, Kes97]. Unknown [BS90]. unordered [CK97]. Unpredictable [BPP98, Hof98]. Unstable [PF91, Pem90]. Unusually [O’B99]. Upcrossing [AL99]. upon [DE96]. Upper [Alb94, Din91, DEW91, HP93, Lig95, Ste91a, Sto95c, HK97]. Urn [Gou93, Pem90]. Using [Dar95, HJ90, BS96, Erh99]. usual [HK97].

V [Num98]. Valeurs [Fer90a, Fer92, Str90]. Value [Lig95, Mou92, NP91, She91a, Van93, RdH96]. Valued [Ber91a, BKK95, Che93a, CS93b, DLM95, DMS93, DK95, DK99, EP99, IMM^{+}90, Kla90, Le 91, Ros95b, Wen90, Alb96, DK96, Duh98, LO98, Pin96]. Values [Hus90, O’B99, dlC90, Fer90a, Fer92, Str90]. variable [DeB96]. Variables [BCL92, Ber91a, Che93a, CGZ95, DJ94, DR91, DH95, DK91b, DK91a, Ein91, Hil92, Hit94, HJ90, Hou95a, Hsi95, KK91, KS91, LT90, LRS90, LRW92, Maj99, OS94, SC92, Sch94, Slu93, Szu90, Wen90, Wu95, AD98, BP96, Ber98, Gan98, HK97, HL96, Lat97, Wu99]. Variance [BLBE91, DH95, Hsi99, MS95, Kle98, Sch96, Uch98]. variance-optimal [Sch96]. variates [CK97]. Variation [AST95, Dud92, MR92a, Qia98, RdH96]. Variational [BCGH93, CPU94, BD98a]. Variations [Tow95]. various [Jak96]. Vector [Ber91a, Ber92a, DMS93, HS90, Ros95b, Xia95, Fer90a]. Vector-Valued [Ber91a, DMS93, Ros95b]. Vectorielles [Fer90a]. Vectors [AL99, Arc94, DZ92, Lah93, Mee90, Ros90b, ST93, ST94]. Velocities [Szn95]. velocity [Car97]. Version [AL93, Hit90, Kif90, MPS91, KS96]. Versions [Iof91]. Vertex [Ben97, PV99]. Vertex-Reinforced [PV99, Ben97]. Via [BCL92, BK93, BP93b, DF90, GK91, MR92b, Pin93, Rio93a, Rio93b, AT93, Rei95, Shi97]. View [dA96]. Viot [DK96, EG93, EK98, ORS95, Sch97]. Viscosity [DIS90]. visité [Leu97]. visited [Leu97]. Vlasov [Cha94]. Volterra [PP90]. Volume [Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90g, Ano90h, Ano90i, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano91i, Ano92a, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano92i, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano93l, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano94i, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95i, Ano96, Bol90, Por90]. Voter [DS93, Han99, Lig94, Ste94, Lig99].

W [dA96]. W. [Lin91b]. Waiting [FGL94, GP92]. Wald [CdlPT93]. Walk [ALP99, Ant95, Asm92, BCGH93, BD94a, Big92, Bol94, BEI92, CL94, DSC93, FG92b, GdH94, GM92, GM94, Kes93, Lal93, LP92, PP95a, PV99, PP99, Pru90, Rev90, Rev93, Sto95b, Tan93, Tot95, BK97, FU96, HS98, LP03, Pia98, Zer98]. Walks [Boi93, Bra96, CD95, HSC93, Hu91, Kee92, KM94a, LW94, LR91, Lee95, Lyo90, Lyo92, MR94, Ros90a, Ros94, Sto95a, Sto95c, SZ99, Tho92, AIM96, Ben97, BS96, BS97, BD98b, CP97, CMP98, Dav96, DH96, HM98, Ham97a, LY98, Lot96, LPP96, Roi96, Shi98a, Tót96, Uch98]. Wasserstein [GMdB99, GMdB03, Tue93]. Watanabe [Sch97]. Watson [PP95b]. Wave [LT98, MSS99, DF98, Mue97]. Wavefront [Pra98]. Waves [FKS91]. Weak [AT92b, Boi93, Che93b, Dav96, Don91, DJ92, HP90, KP91, Myt98, Rei95, Sch93a, SY96, Sil90, Xia94a]. Weakly [Bry92, DH95, Lah93, Lal99, LS98, Lal02]. Wedge [DeB90, DeB96]. Weighted [Ein92, IL93, Maj94, OR93, PW93, Sch93a, AGM96, SY96, Zha97]. Welsh [Roy90]. Wentzell [Kif90, Mik91]. Which [BL88, Ham95, Mee90, BL97, MvW91]. White [Enc98, Sow94]. Wiener [Bol90, CFR91, CN93, Enc93a, Enc93b, ES93, Har96, HPA95, IN94, Kal91, LZ99, Roc92, She91a, Uch98]. Wigner [Bai93a]. Wills [Vit96]. Windings [Shi98a, BW96a]. Wishart [PR91, PR92]. within [Hog90]. without [BP98]. Wolfowitz [Mas90]. Words [BK95b]. Work [Gri93, Kes93].

Xiang [Pol92]. Xiang-Qun [Pol92].

Y. [Sta90]. Yang [Pol92]. Yor [Dur93].

Zero [FGL94, Tal94c, AGM96, Lan96, LSV96, RS96, SX96, Uch98, JLQY99]. Zero-One [Tal94c, AGM96, RS96]. Zero-Range [FGL94, Lan96, LSV96, SX96, JLQY99]. zeta [CP97]. Zheng [Kur91]. Zygmund [Rio95b].

References

Adelman:1998:SAB

[ABP98] Omer Adelman, Krzysztof Burdzy, and Robin Pemantle. Sets avoided by Brownian motion. Annals of Probability, 26(2):429– 464, April 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022855639.

Aaronson:1998:CFR

[AD98] Jon Aaronson and Manfred Denker. Characteristic functions of random variables attracted to 1-stable laws. Annals of Probability, 26(1):399–415, January 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022855426.

Adler:1990:NCP

[Adl90] Robert J. Adler. The net charge process for interacting, signed diffusions. Annals of Probability, 18(2):602–625, April 1990. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990847.

Adler:1991:ILT

[AFL91] Robert J. Adler, Raisa Epstein Feldman, and Marica Lewin. Intersection local times for infinite systems of Brownian motions and for the Brownian density process. Annals of Probability, 19 (1):192–220, January 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990540.

Alabert:1995:MFP

[AFN95] Aureli Alabert, Marco Ferrante, and David Nualart. Markov field property of stochastic differential equations. Annals of Probability, 23(3):1262–1288, July 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988183.

Arcones:1993:LTP

[AG93] Miguel A. Arcones and Evarist Gine. Limit theorems for Uprocesses. Annals of Probability, 21(3):1494–1542, July 1993. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989128.

Arous:1997:SLS

[AG97] G. Ben Arous and A. Guionnet. Symmetric Langevin spin glass dynamics. Annals of Probability, 25(3):1367–1422, July 1997. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1024404517.

Arenal-Gutierrez:1996:ZOL

[AGM96] Eusebio Arenal-Gutiérrez and Carlos Matrán. A zero-one law approach to the central limit theorem for the weighted bootstrap mean. Annals of Probability, 24(1):532–540, January 1996. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1042644731.

Ahn:1997:SIR

[Ahn97] Hyungsok Ahn. Semimartingale integral representation. Annals of Probability, 25(2):997–1010, April 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1024404427.

Arnold:1999:RNL

[AI99] Ludwig Arnold and Peter Imkeller. Rotation numbers for linear stochastic differential equations. Annals of Probability, 27 (1):130–149, January 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022677256.

Aspandiiarov:1996:PTM

[AIM96] S. Aspandiiarov, R. Iasnogorodski, and M. Menshikov. Passagetime moments for nonnegative stochastic processes and an application to reflected random walks in a quadrant. Annals of Probability, 24(2):932–960, April 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1039639371.

Alexander:1992:RSP

[AK92] Kenneth S. Alexander and Steven A. Kalikow. Random stationary processes. Annals of Probability, 20(3):1174–1198, July 1992. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989685.

Accardi:1993:CVF

[AL93] L. Accardi and Y. G. Lu. A continuous version of De Finetti’s theorem. Annals of Probability, 21(3):1478–1493, July 1993. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989127.

Aspandiiarov:1995:SNC

[AL95] Sanjar Aspandiiarov and Jean-François Le Gall. Some new classes of exceptional times of linear Brownian motion. Annals of Probability, 23(4):1605–1626, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176987795.

Albin:1999:ABC

[AL99] J. M. P. Albin and M. R. Leadbetter. Asymptotic behavior of conditional laws and moments of ∞-stable random vectors, with application to upcrossing intensities. Annals of Probability, 27 (3):1468–1500, July 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022677455.

Albin:1990:ETS

[Alb90] J. M. P. Albin. On extremal theory for stationary processes. Annals of Probability, 18(1):92–128, January 1990. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990940.

Albin:1994:ULC

[Alb94] J. M. P. Albin. On the upper and lower classes for a stationary Gaussian stochastic process. Annals of Probability, 22(1):77–93, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988848.

Albin:1996:MVG

[Alb96] J. M. P. Albin. Minima of H-valued Gaussian processes. Annals of Probability, 24(2):788–824, April 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1039639362.

Albin:1998:ETS

[Alb98] J. M. P. Albin. On extremal theory for self-similar processes. Annals of Probability, 26(2):743–793, April 1998. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1022855649.

Aldous:1991:CRT

[Ald91] David Aldous. The continuum random tree. I. Annals of Probability, 19(1):1–28, January 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990534.

Aldous:1993:CRT

[Ald93] David Aldous. The continuum random tree III. Annals of Probability, 21(1):248–289, January 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989404.

Aldous:1994:RSS

[Ald94] David Aldous. Recursive self-similarity for random trees, random triangulations and Brownian excursion. Annals of Probability, 22 (2):527–545, April 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988720.

Aldous:1997:BEC

[Ald97] David Aldous. Brownian excursions, critical random graphs and the multiplicative coalescent. Annals of Probability, 25(2):812– 854, April 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1024404421.

Alexander:1990:LBC

[Ale90] Kenneth S. Alexander. Lower bounds on the connectivity function in all directions for Bernoulli percolation in two and three dimensions. Annals of Probability, 18(4):1547–1562, October 1990. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990631.

Alexander:1995:PMS

[Ale95] Kenneth S. Alexander. Percolation and minimal spanning forests in infinite graphs. Annals of Probability, 23(1):87–104, January 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988378.

Alexander:1996:BLL

[Ale96] Kenneth S. Alexander. Boundedness of level lines for twodimensional random fields. Annals of Probability, 24(4):1653– 1674, October 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1041903201.

Alexander:1997:ASF

[Ale97] Kenneth S. Alexander. Approximation of subadditive functions and convergence rates in limiting-shape results. Annals of Probability, 25(1):30–55, January 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1024404277.

Algoet:1992:USP

[Alg92] Paul Algoet. Universal schemes for prediction, gambling and portfolio selection. Annals of Probability, 20(2):901–941, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.

aop/1176989811. See correction [Alg95].

Algoet:1995:CUP

[Alg95] Paul H. Algoet. Correction: Universal prediction schemes. Annals of Probability, 23(1):474–478, January 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988396. See [Alg92].

Ancona:1999:CEC

[ALP99] Alano Ancona, Russell Lyons, and Yuval Peres. Crossing estimates and convergence of Dirichlet functions along random walk and diffusion paths. Annals of Probability, 27(2):970–989, April 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022677392.

Alsmeyer:1992:GRM

[Als92] Gerold Alsmeyer. On generalized renewal measures and certain first passage times. Annals of Probability, 20(3):1229–1247, July 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989690.

Adler:1990:CLT

[AMZ90] Robert J. Adler, Michael B. Marcus, and Joel Zinn. Central Limit Theorems for the local times of certain Markov processes and the squares of Gaussian processes. Annals of Probability, 18 (3):1126–1140, July 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990738.

Adams:1998:DEE

[AN98] Terrence M. Adams and Andrew B. Nobel. On density estimation from ergodic processes. Annals of Probability, 26(2):794–804, April 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022855650.

Andjel:1993:CET

[And93] Enrique D. Andjel. Characteristic exponents for two-dimensional bootstrap percolation. Annals of Probability, 21(2):926–935, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989275.

Anonymous:1990:MBPa

[Ano90a] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 18, Number 1 (1990). Annals of Probability, 18(1):??, January 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990959.

Anonymous:1990:MBPb

[Ano90b] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 18, Number 2 (1990). Annals of Probability, 18(2):??, April 1990. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990867.

Anonymous:1990:MBPc

[Ano90c] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 18, Number 3 (1990). Annals of Probability, 18(3):??, July 1990. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990753.

Anonymous:1990:MBPd

[Ano90d] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 18, Number 4 (1990). Annals of Probability, 18(4):??, October 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990653.

Anonymous:1990:MFPa

[Ano90e] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 18, Number 1 (1990). Annals of Probability, 18(1):??, January 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990935.

Anonymous:1990:MFPb

[Ano90f] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 18, Number 2 (1990). Annals of Probability, 18(2):??, April 1990. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990839.

Anonymous:1990:MFPc

[Ano90g] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 18, Number 3 (1990). Annals of Probability, 18(3):??, July 1990. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990728.

Anonymous:1990:MFPd

[Ano90h] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 18, Number 4 (1990). Annals of Probability, 18(4):??, October 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990625.

Anonymous:1990:VI

[Ano90i] Anonymous. Volume information. Annals of Probability, 18(4):??, October 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990624.

Anonymous:1991:MBPa

[Ano91a] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 19, Number 1 (1991). Annals of Probability, 19(1):??, January 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990556.

Anonymous:1991:MBPb DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990458.

Anonymous:1991:MBPc

[Ano91c] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 19, Number 3 (1991). Annals of Probability, 19(3):??, July 1991. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990352.

Anonymous:1991:MBPd

[Ano91d] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 19, Number 4 (1991). Annals of Probability, 19(4):??, October 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990239.

Anonymous:1991:MFPa

[Ano91e] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 19, Number 1 (1991). Annals of Probability, 19(1):??, January 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990533.

Anonymous:1991:MFPb

[Ano91f] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 19, Number 2 (1991). Annals of Probability, 19(2):??, April 1991. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990434.

Anonymous:1991:MFPc

[Ano91g] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 19, Number 3 (1991). Annals of Probability, 19(3):??, July 1991. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990327.

Anonymous:1991:MFPd

[Ano91h] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 19, Number 4 (1991). Annals of Probability, 19(4):??, October 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990217.

Anonymous:1991:VI

[Ano91i] Anonymous. Volume information. Annals of Probability, 19(4):??, October 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990216.

Anonymous:1992:MBPa

[Ano92a] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 20, Number 1 (1992). Annals of Probability, 20(1):??, January 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989942.

Anonymous:1992:MBPb

[Ano92b] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 20, Number 2 (1992). Annals of Probability, 20(2):??, April 1992. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989823.

Anonymous:1992:MBPc

[Ano92c] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 20, Number 3 (1992). Annals of Probability, 20(3):??, July 1992. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989710.

Anonymous:1992:MBPd

[Ano92d] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 20, Number 4 (1992). Annals of Probability, 20(4):??, October 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989545.

Anonymous:1992:MFPa

[Ano92e] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 20, Number 1 (1992). Annals of Probability, 20(1):??, January 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989915.

Anonymous:1992:MFPb

[Ano92f] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 20, Number 2 (1992). Annals of Probability, 20(2):??, April 1992. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989789.

Anonymous:1992:MFPc

[Ano92g] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 20, Number 3 (1992). Annals of Probability, 20(3):??, July 1992. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989681.

Anonymous:1992:MFPd

[Ano92h] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 20, Number 4 (1992). Annals of Probability, 20(4):??, October 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989523.

Anonymous:1992:VI

[Ano92i] Anonymous. Volume information. Annals of Probability, 20(4):??, October 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989522.

Anonymous:1993:MBPa

[Ano93a] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 21, Number 1 (1993). Annals of Probability, 21(1):??, January 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989418.

Anonymous:1993:MBPb

[Ano93b] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 21, Number 2 (1993). Annals of Probability, 21(2):??, April 1993. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989288.

Anonymous:1993:MBPc

[Ano93c] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 21, Number 3 (1993). Annals of Probability, 21(3):??, July 1993. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989140.

Anonymous:1993:MBPd

[Ano93d] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 21, Number 4 (1993). Annals of Probability, 21(4):??, October 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989021.

Anonymous:1993:MFPa

[Ano93e] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 21, Number 1 (1993). Annals of Probability, 21(1):??, January 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989389.

Anonymous:1993:MFPb

[Ano93f] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 21, Number 2 (1993). Annals of Probability, 21(2):??, April 1993. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989255.

Anonymous:1993:MFPc

[Ano93g] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 21, Number 3 (1993). Annals of Probability, 21(3):??, July 1993. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989115.

Anonymous:1993:MFPd

[Ano93h] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 21, Number 4 (1993). Annals of Probability, 21(4):??, October 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988998.

Anonymous:1993:PFSa

[Ano93i] Anonymous. [photograph]: Frank Spitzer. Annals of Probability, 21(2):591, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989256.

Anonymous:1993:PFSb

[Ano93j] Anonymous. Publications of Frank Spitzer. Annals of Probability, 21(2):622–623, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989259.

Anonymous:1993:SFS

[Ano93k] Anonymous. Students of Frank Spitzer. Annals of Probability, 21(2):624, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989260.

Anonymous:1993:VI

[Ano93l] Anonymous. Volume information. Annals of Probability, 21(4):??, October 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988997.

Anonymous:1994:MBPa

[Ano94a] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 22, Number 1 (1994). Annals of Probability, 22(1):??, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988872.

Anonymous:1994:MBPb

[Ano94b] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 22, Number 2 (1994). Annals of Probability, 22(2):??, April 1994. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988745.

Anonymous:1994:MBPc

[Ano94c] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 22, Number 3 (1994). Annals of Probability, 22(3):??, July 1994. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988619.

Anonymous:1994:MBPd

[Ano94d] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 22, Number 4 (1994). Annals of Probability, 22(4):??, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988504.

Anonymous:1994:MFPa

[Ano94e] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 22, Number 1 (1994). Annals of Probability, 22(1):??, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988844.

Anonymous:1994:MFPb

[Ano94f] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 22, Number 2 (1994). Annals of Probability, 22(2):??, April 1994. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988719.

Anonymous:1994:MFPc

[Ano94g] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 22, Number 3 (1994). Annals of Probability, 22(3):??, July 1994. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988596.

Anonymous:1994:MFPd

[Ano94h] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 22, Number 4 (1994). Annals of Probability, 22(4):??, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988476.

Anonymous:1994:VI

[Ano94i] Anonymous. Volume information. Annals of Probability, 22(4):??, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988475.

Anonymous:1995:MBPa

[Ano95a] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 23, Number 1 (1995). Annals of Probability, 23(1):??, January 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988397.

Anonymous:1995:MBPb DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988299.

Anonymous:1995:MBPc

[Ano95c] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 23, Number 3 (1995). Annals of Probability, 23(3):??, July 1995. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988193.

Anonymous:1995:MBPd

[Ano95d] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 23, Number 4 (1995). Annals of Probability, 23(4):??, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176987816.

Anonymous:1995:MFPa

[Ano95e] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 23, Number 1 (1995). Annals of Probability, 23(1):??, January 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988373.

Anonymous:1995:MFPb

[Ano95f] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 23, Number 2 (1995). Annals of Probability, 23(2):??, April 1995. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988275.

Anonymous:1995:MFPc

[Ano95g] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 23, Number 3 (1995). Annals of Probability, 23(3):??, July 1995. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988170.

Anonymous:1995:MFPd

[Ano95h] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 23, Number 4 (1995). Annals of Probability, 23(4):??, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176987790.

Anonymous:1995:VI

[Ano95i] Anonymous. Volume information. Annals of Probability, 23(4):??, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176987789.

Anonymous:1996:CV

[Ano96] Anonymous. Contents of volume 24. Annals of Probability, 24(4):??, October 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1041903225.

Antal:1995:EOS

[Ant95] Peter Antal. Enlargement of obstacles for the simple random walk. Annals of Probability, 23(3):1061–1101, July 1995. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988174.

Amit:1991:NMP

[AP91] Yali Amit and Mauro Piccioni. A nonhomogeneous Markov process for the estimation of Gaussian random fields with nonlinear observations. Annals of Probability, 19(4):1664–1678, October 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990228.

Antal:1996:CDS

[AP96] Peter Antal and Agoston Pisztora. On the chemical distance for supercritical Bernoulli percolation. Annals of Probability, 24 (2):1036–1048, April 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1039639377.

Aldous:1998:SAC

[AP98] David Aldous and Jim Pitman. The standard additive coalescent. Annals of Probability, 26(4):1703–1726, October 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1022855879.

Adler:1993:ILT

[AR93] Robert J. Adler and Jay S. Rosen. Intersection local times of all orders for Brownian and stable density processes — construction, renormalisation and limit laws. Annals of Probability, 21 (2):1073–1123, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989283.

Arcones:1994:LTN

[Arc94] Miguel A. Arcones. Limit theorems for nonlinear functionals of a stationary Gaussian sequence of vectors. Annals of Probability, 22 (4):2242–2274, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988503.

Arnaud:1994:SPI

[Arn94] Jean-Pierre Arnaud. Stationary processes indexed by a homogeneous tree. Annals of Probability, 22(1):195–218, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988856.

Albeverio:1993:GTS

[ARZ93] S. Albeverio, M. Rockner, and T. S. Zhang. Girsanov transform for symmetric diffusions with infinite dimensional state space. Annals of Probability, 21(2):961–978, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989277.

Adler:1995:SFB

[AS95] Robert J. Adler and Gennady Samorodnitsky. Super fractional Brownian motion, fractional super Brownian motion and related self-similar (super) processes. Annals of Probability, 23(2):743– 766, April 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988287.

Asmussen:1992:PTR

[Asm92] Soren Asmussen. Phase-type representations in random walk and queueing problems. Annals of Probability, 20(2):772–789, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989805.

Arratia:1995:TVA

[AST95] Richard Arratia, Dudley Stark, and Simon Tavare. Total variation asymptotics for Poisson process approximations of logarithmic combinatorial assemblies. Annals of Probability, 23(3):1347–1388, July 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988188.

Arratia:1992:CSR

[AT92a] Richard Arratia and Simon Tavare. The cycle structure of random permutations. Annals of Probability, 20(3):1567–1591, July 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989707.

Avram:1992:WCS

[AT92b] Florin Avram and Murad S. Taqqu. Weak convergence of sums of moving averages in the α-stable domain of attraction. Annals of Probability, 20(1):483–503, January 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989938.

Arratia:1993:RDA

[AT93] Richard Arratia and Simon Tavare. Reviews: D. Aldous, Probability Approximations via the Poisson Clumping Heuristic; A. D. Barbour, L. Holst, S. Janson, Poisson Approximation. Annals of Probability, 21(4):2269–2279, October 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989020.

Atar:1998:ESN

[Ata98] Rami Atar. Exponential stability for nonlinear filtering of diffusion processes in a noncompact domain. Annals of Probability, 26(4): 1552–1574, October 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022855873.

Athreya:1992:MSC

[Ath92] K. B. Athreya. On a maximum sequence in a critical multitype branching process. Annals of Probability, 20(2):746–752, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989803.

Barsky:1991:PCE

[BA91] D. J. Barsky and M. Aizenman. Percolation critical exponents under the triangle condition. Annals of Probability, 19(4):1520–1536, October 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990221.

Baccelli:1992:ETS

[Bac92] François Baccelli. Ergodic theory of stochastic Petri networks. Annals of Probability, 20(1):375–396, January 1992. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989932.

Bai:1993:CREa

[Bai93a] Z. D. Bai. Convergence rate of expected spectral distributions of large random matrices. Part I. Wigner matrices. Annals of Probability, 21(2):625–648, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989261;
http://www.jstor.org/stable/2244669.

Bai:1993:CREb

[Bai93b] Z. D. Bai. Convergence rate of expected spectral distributions of large random matrices. Part II. Sample covariance matrices. Annals of Probability, 21(2):649–672, April 1993. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989262.

Bai:1997:CL

[Bai97] Z. D. Bai. Circular law. Annals of Probability, 25(1):494–529, January 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1024404298.

Baldi:1991:LDD

[Bal91] Paolo Baldi. Large deviations for diffusion processes with homogenization and applications. Annals of Probability, 19(2):509–524, April 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990438.

Baldi:1995:EAP

[Bal95] Paolo Baldi. Exact asymptotics for the probability of exit from a domain and applications to simulation. Annals of Probability, 23 (4):1644–1670, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176987797.

Bartoszek:1994:CFD

[Bar94] Wojciech Bartoszek. On concentration functions on discrete groups. Annals of Probability, 22(3):1596–1599, July 1994. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988614.

Barany:1999:SQP

[Bár99] Imre Bárány. Sylvester’s question: The probability that n points are in convex position. Annals of Probability, 27(4):2020–2034, October 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022874826.

Bass:1997:BRB

[Bas97] Richard F. Bass. Book review of “Positive Harmonic Functions and Diffusion: An Integrated Analytic and Probabilistic Approach” by Ross G. Pinsky. Annals of Probability, 25(3):1532–1535, July 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1024404524.

Barbour:1992:SCM

[BB92a] A. D. Barbour and Timothy C. Brown. The Stein–Chen method, point processes and compensators. Annals of Probability, 20 (3):1504–1527, July 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989704.

Basak:1992:SDC

[BB92b] Gopal K. Basak and Rabi N. Bhattacharya. Stability in distribution for a class of singular diffusions. Annals of Probability, 20 (1):312–321, January 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989928.

Bhatt:1996:OMC

[BB96] Abhay G. Bhatt and Vivek S. Borkar. Occupation measures for controlled Markov processes: characterization and optimality. Annals of Probability, 24(3):1531–1562, July 1996. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1065725192.

Bass:1999:SBM

[BB99] Richard F. Bass and Krzysztof Burdzy. Stochastic bifurcation models. Annals of Probability, 27(1):50–108, January 1999. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1022677254.

Babillot:1997:RDE

[BBE97] Martine Babillot, Philippe Bougerol, and Laure Elie. The random difference equation X_{n} = A_{n}X_{n−1}+B_{n} in the critical case. Annals of Probability, 25(1):478–493, January 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1024404297.

Bass:1994:ILT

[BBK94] Richard F. Bass, Krzysztof Burdzy, and Davar Khoshnevisan. Intersection local time for points of infinite multiplicity. Annals of Probability, 22(2):566–625, April 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988722.

Bramson:1996:SMS

[BCD96] Maury Bramson, J. Theodore Cox, and Richard Durrett. Spatial models for species area curves. Annals of Probability, 24(4): 1727–1751, October 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1041903204.

Bramson:1998:SMA

[BCD98] Maury Bramson, J. Theodore Cox, and Richard Durrett. A spatial model for the abundance of species. Annals of Probability, 26 (2):658–709, April 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022855647.

Bramson:1993:ECS

[BCG93] Maury Bramson, J. T. Cox, and Andreas Greven. Ergodicity of critical spatial branching processes in low dimensions. Annals of Probability, 21(4):1946–1957, October 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989006.

Bramson:1997:IMC

[BCG97] Maury Bramson, J. T. Cox, and Andreas Greven. Invariant measures of critical spatial branching processes in high dimensions. Annals of Probability, 25(1):56–70, January 1997. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1024404278.

Baillon:1993:VAB

[BCGH93] J.-B. Baillon, Ph. Clement, A. Greven, and F. Den Hollander. A variational approach to branching random walk in random environment. Annals of Probability, 21(1):290–317, January 1993. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989405.

Barbour:1992:CPA

[BCL92] A. D. Barbour, Louis H. Y. Chen, and Wei-Liem Loh. Compound Poisson approximation for nonnegative random variables via Stein’s method. Annals of Probability, 20(4):1843–1866, October 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989531.

Berkes:1993:SLT

[BD93a] István Berkes and H. Dehling. Some limit theorems in log density. Annals of Probability, 21(3):1640–1670, July 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989135.

Bolthausen:1993:CLD

[BD93b] Erwin Bolthausen and Jean-Dominique Deuschel. Critical large deviations for Gaussian fields in the phase transition regime, I. Annals of Probability, 21(4):1876–1920, October 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989003.

Bertoin:1994:CRW

[BD94a] J. Bertoin and R. A. Doney. On conditioning a random walk to stay nonnegative. Annals of Probability, 22(4):2152–2167, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988497.

Bose:1994:SAP

[BD94b] Arup Bose and Ratan Dasgupta. On some asymptotic properties of U statistics and one-sided estimates. Annals of Probability, 22 (4):1715–1724, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988479.

Boue:1998:VRC

[BD98a] Michelle Boué and Paul Dupuis. A variational representation for certain functionals of Brownian motion. Annals of Probability, 26 (4):1641–1659, October 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022855876.

Brown:1998:RWH

[BD98b] Kenneth S. Brown and Persi Diaconis. Random walks and hyperplane arrangements. Annals of Probability, 26(4):1813–1854, October 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022855884.

Bolthausen:1997:LTP

[BdH97] Erwin Bolthausen and Frank den Hollander. Localization transition for a polymer near an interface. Annals of Probability, 25 (3):1334–1366, July 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1024404516.

Bramson:1991:CPR

[BDS91] Maury Bramson, Rick Durrett, and Roberto H. Schonmann. The contact processes in a random environment. Annals of Probability, 19(3):960–983, July 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990331.

Bolthausen:1995:LAL

[BDT95] Erwin Bolthausen, Jean-Dominique Deuschel, and Yozo Tamura. Laplace approximations for large deviations of nonreversible Markov processes. The nondegenerate case. Annals of Probability, 23(1):236–267, January 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988385.

Brydges:1992:SAW

[BEI92] David Brydges, Steven N. Evans, and John Z. Imbrie. Selfavoiding walk on a hierarchical lattice in four dimensions. Annals of Probability, 20(1):82–124, January 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989919.

Benaim:1997:VRR

[Ben97] Michel Benaïm. Vertex-reinforced random walks and a conjecture of Pemantle. Annals of Probability, 25(1):361–392, January 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1024404292.

Berger:1991:MEI

[Ber91a] Erich Berger. Majorization, exponential inequalities and almost sure behavior of vector-valued random variables. Annals of Probability, 19(3):1206–1226, July 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990341.

Berman:1991:SIL

[Ber91b] Simeon M. Berman. Self-intersections and local nondeterminism of Gaussian processes. Annals of Probability, 19(1):160–191, January 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990539.

Berman:1992:CLT

[Ber92a] Simeon M. Berman. A Central Limit Theorem for the renormalized self-intersection local time of a stationary vector Gaussian process. Annals of Probability, 20(1):61–81, January 1992. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989918.

Bertoin:1992:EPT

[Ber92b] Jean Bertoin. An extension of Pitman’s theorem for spectrally positive Lévy processes. Annals of Probability, 20(3):1464–1483, July 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989701.

Bertoin:1998:DET

[Ber98] Jean Bertoin. Darling–Erdős theorems for normalized sums of i.i.d. variables close to a stable law. Annals of Probability, 26 (2):832–852, April 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022855652.

Bertoin:1999:RTE

[Ber99] Jean Bertoin. Renewal theory for embedded regenerative sets. Annals of Probability, 27(3):1523–1535, July 1999. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1022677457.

Boucher:1999:SRM

[BET99] Christopher Boucher, Richard S. Ellis, and Bruce Turkington. Spatializing random measures: Doubly indexed processes and the large deviation principle. Annals of Probability, 27(1):297–324, January 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.

aop/1022677264. See erratum [BET02].

Boucher:2002:ESR

[BET02] Christopher Boucher, Richard S. Ellis, and Bruce Turkington. Erratum: Spatializing random measures: Double indexed processes and the large deviation principle. Annals of Probability, 30(4):2113, October 2002. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1039548384. See [BET99].

Bezuidenhout:1992:RGG

[Bez92] Carol Bezuidenhout. Review: Geoffrey Grimmett, Percolation. Annals of Probability, 20(2):1099–1104, April 1992. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989820.

Burton:1990:PDP

[BF90] Robert M. Burton and Marie M. Franzosa. Positive dependence properties of point processes. Annals of Probability, 18(1):359–377, January 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990953.

Bezuidenhout:1990:CCP

[BG90] Carol Bezuidenhout and Geoffrey Grimmett. The critical contact process dies out. Annals of Probability, 18(4):1462–1482, October 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990627.

Bezuidenhout:1991:EDS

[BG91] Carol Bezuidenhout and Geoffrey Grimmett. Exponential decay for subcritical contact and percolation processes. Annals of Probability, 19(3):984–1009, July 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990332.

Betz:1994:HSE

[BG94a] C. Betz and H. Gzyl. Hitting spheres from the exterior. Annals of Probability, 22(1):177–179, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988854.

Bezuidenhout:1994:CAS

[BG94b] Carol Bezuidenhout and Lawrence Gray. Critical attractive spin systems. Annals of Probability, 22(3):1160–1194, July 1994. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988599.

Bentkus:1996:BEB

[BG96a] V. Bentkus and F. Götze. The Berry–Esseen bound for Student’s statistic. Annals of Probability, 24(1):491–503, January 1996. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1042644728.

Bentkus:1996:ORC

[BG96b] V. Bentkus and F. Götze. Optimal rates of convergence in the CLT for quadratic forms. Annals of Probability, 24(1):466–490, January 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1042644727.

Bovier:1996:ASL

[BG96c] Anton Bovier and Véronique Gayrard. An almost sure large deviation principle for the Hopfield model. Annals of Probability, 24(3):1444–1475, July 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1065725188.

Bentkus:1999:OBN

[BG99] V. Bentkus and F. Götze. Optimal bounds in non-Gaussian limit theorems for U-statistics. Annals of Probability, 27(1):454–521, January 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022677269.

Burton:1993:DPB

[BGM93] Robert M. Burton, Marc Goulet, and Ronald Meester. On 1dependent processes and k-block factors. Annals of Probability, 21 (4):2157–2168, October 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989014.

Bentkus:1994:LEC

[BGZ94] Vidmantas Bentkus, Friedrich Gotze, and Ricardas Zitikis. Lower estimates of the convergence rate for U-statistics. Annals of Probability, 22(4):1707–1714, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988478.

Bass:1991:SPT

[BH91] Richard F. Bass and Pei Hsu. Some potential theory for reflecting Brownian motion in holder and Lipschitz domains. Annals of Probability, 19(2):486–508, April 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990437.

Bolthausen:1994:SAB

[BH94] Erwin Bolthausen and Frank Den Hollander. Survival asymptotics for Brownian motion in a Poisson field of decaying traps. Annals of Probability, 22(1):160–176, January 1994. CODEN APBYAE.

ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988853.

Bobkov:1997:ICP

[BH97] S. G. Bobkov and C. Houdré. Isoperimetric constants for product probability measures. Annals of Probability, 25(1):184–205, January 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1024404284.

Biggins:1992:UCM

[Big92] J. D. Biggins. Uniform convergence of martingales in the branching random walk. Annals of Probability, 20(1):137–151, January 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989921.

Bishop:1992:BMD

[Bis92] Christopher J. Bishop. Brownian motion in Denjoy domains. Annals of Probability, 20(2):631–651, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989795.

Bhatt:1993:IME

[BK93] Abhay G. Bhatt and Rajeeva L. Karandikar. Invariant measures and evolution equations for Markov processes characterized via martingale problems. Annals of Probability, 21(4):2246–2268, October 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989019.

Bass:1995:LIL

[BK95a] Richard F. Bass and Davar Khoshnevisan. Laws of the iterated logarithm for local times of the empirical process. Annals of Probability, 23(1):388–399, January 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988391.

Benjamini:1995:PAW

[BK95b] Itai Benjamini and Harry Kesten. Percolation of arbitrary words in {0, 1}N. Annals of Probability, 23(3):1024–1060, July 1995. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988173.

Biggins:1997:SHN

[BK97] J. D. Biggins and A. E. Kyprianou. Seneta–Heyde norming in the branching random walk. Annals of Probability, 25(1):337–360, January 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1024404291.

Bhatt:1995:URS

[BKK95] Abhay G. Bhatt, G. Kallianpur, and Rajeeva L. Karandikar. Uniqueness and robustness of solution of measure-valued equations of nonlinear filtering. Annals of Probability, 23(4):1895–1938, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176987808.

Bhattacharya:1988:ACM

[BL88] Rabi N. Bhattacharya and Oesook Lee. Asymptotics of a class of Markov processes which are not in general irreducible. Annals of Probability, 16(3):1333–1347, July 1988. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991694. See cor-

rection [BL97].

Burdzy:1990:NEB

[BL90] Krzysztof Burdzy and Gregory F. Lawler. Nonintersection exponents for Brownian paths. II. estimates and applications to a random fractal. Annals of Probability, 18(3):981–1009, July 1990. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990733.

Baccelli:1992:CSR

[BL92a] François Baccelli and Zhen Liu. On a class of stochastic recursive sequences arising in queueing theory. Annals of Probability, 20 (1):350–374, January 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989931.

Bertoin:1992:RMB

[BL92b] J. Bertoin and Y. Le Jan. Representation of measures by balayage from a regular recurrent point. Annals of Probability, 20 (1):538–548, January 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989940.

Bhattacharya:1997:CBC

[BL97] Rabi N. Bhattacharya and Oesook Lee. Correction: “Asymptotics of a class of Markov processes which are not in general irreducible” [Ann. Probab. 16 (1988), no. 3, 1333–1347; MR 89m:60148]. Annals of Probability, 25(3):1541–1543, July 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1024404526. See

[BL88].

Brossard:1999:NRP

[BL99] Jean Brossard and Christophe Leuridan. Numérotation des records d’un processus de Poisson ponctuel (french) [the numbering of records of a Poisson point process]. Annals of Probability, 27(3):1304–1323, July 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022677448.

Bar-Lev:1991:VFM

[BLBE91] Shaul K. Bar-Lev, Daoud Bshouty, and Peter Enis. Variance functions with meromorphic means. Annals of Probability, 19 (3):1349–1366, July 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990348.

Blount:1991:CSD

[Blo91] Douglas Blount. Comparison of stochastic and deterministic models of a linear chemical reaction with diffusion. Annals of Probability, 19(4):1440–1462, October 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990219.

Blount:1994:DDL

[Blo94] Douglas Blount. Density-dependent limits for a nonlinear reactiondiffusion model. Annals of Probability, 22(4):2040–2070, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988494.

Blount:1996:DLN

[Blo96] Douglas Blount. Diffusion limits for a nonlinear density dependent space-time population model. Annals of Probability, 24 (2):639–659, April 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1039639357.

Bloznelis:1999:BEB

[Blo99] M. Bloznelis. A Berry–Esseen bound for finite population Student’s statistic. Annals of Probability, 27(4):2089–2108, October 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022874830.

Benjamini:1999:CPN

[BLPS99] Itai Benjamini, Russell Lyons, Yuval Peres, and Oded Schramm. Critical percolation on any nonamenable group has no infinite clusters. Annals of Probability, 27(3):1347–1356, July 1999. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1022677450.

Bell:1995:SDD

[BM95] Denis R. Bell and Salah-Eldin A. Mohammed. Smooth densities for degenerate stochastic delay equations with hereditary drift. Annals of Probability, 23(4):1875–1894, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176987807.

Bremaud:1996:SNH

[BM96] Pierre Brémaud and Laurent Massoulié. Stability of nonlinear Hawkes processes. Annals of Probability, 24(3):1563–1588, July 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1065725193.

Belisle:1997:WPM

[BMR97] Claude Bélisle, Jean-Claude Massé, and Thomas Ransford. When is a probability measure determined by infinitely many projections? Annals of Probability, 25(2):767–786, April 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1024404418.

Bally:1995:AST

[BMSS95] Vlad Bally, Annie Millet, and Marta Sanz-Sole. Approximation and support theorem in holder norm for parabolic stochastic partial differential equations. Annals of Probability, 23(1):178–222, January 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988383.

Bramson:1994:SOD

[BN94] Maury Bramson and Claudia Neuhauser. Survival of onedimensional cellular automata under random perturbations. Annals of Probability, 22(1):244–263, January 1994. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988858.

Bobkov:1994:IID

[Bob94] S. G. Bobkov. Isoperimetric inequalities for distributions of exponential type. Annals of Probability, 22(2):978–994, April 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988737.

Bobkov:1996:EPH

[Bob96] S. Bobkov. Extremal properties of half-spaces for log-concave distributions. Annals of Probability, 24(1):35–48, January 1996. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1042644706.

Bobkov:1997:IID

[Bob97] S. G. Bobkov. An isoperimetric inequality on the discrete cube, and an elementary proof of the isoperimetric inequality in Gauss space. Annals of Probability, 25(1):206–214, January 1997. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1024404285.

Bobkov:1999:IAI

[Bob99a] S. G. Bobkov. Isoperimetric and analytic inequalities for logconcave probability measures. Annals of Probability, 27(4): 1903–1921, October 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022874820.

Bobkov:1999:SSC

[Bob99b] S. G. Bobkov. The size of singular component and shift inequalities. Annals of Probability, 27(1):416–431, January 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1022677267.

Boivin:1993:WCR

[Boi93] Daniel Boivin. Weak convergence for reversible random walks in a random environment. Annals of Probability, 21(3):1427–1440, July 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989125.

Bolthausen:1990:VWS

[Bol90] E. Bolthausen. On the volume of the Wiener sausage. Annals of Probability, 18(4):1576–1582, October 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990633.

Bolthausen:1994:LTD

[Bol94] Erwin Bolthausen. Localization of a two-dimensional random walk with an attractive path interaction. Annals of Probability, 22 (2):875–918, April 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988734.

Borell:1993:GPS

[Bor93] Christer Borell. Geometric properties of some familiar diffusions in R^{n}. Annals of Probability, 21(1):482–489, January 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989412.

Bougerol:1992:SSG

[BP92] Philippe Bougerol and Nico Picard. Strict stationarity of generalized autoregressive processes. Annals of Probability, 20(4): 1714–1730, October 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989526.

Bovier:1993:LIL

[BP93a] Anton Bovier and Pierre Picco. A law of the iterated logarithm for random geometric series. Annals of Probability, 21(1):168–184, January 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989399.

Burton:1993:LCE

[BP93b] Robert Burton and Robin Pemantle. Local characteristics, entropy and limit theorems for spanning trees and domino tilings via transfer-impedances. Annals of Probability, 21(3):1329–1371, July 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989121.

Barlow:1994:FHB

[BP94a] Martin T. Barlow and Edwin A. Perkins. On the filtration of historical Brownian motion. Annals of Probability, 22(3):1273–1294, July 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988603.

Benjamini:1994:MCI

[BP94b] Itai Benjamini and Yuval Peres. Markov chains indexed by trees. Annals of Probability, 22(1):219–243, January 1994. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988857.

Bercovici:1996:LLN

[BP96] Hari Bercovici and Vittorino Pata. The law of large numbers for free identically distributed random variables. Annals of Probability, 24(1):453–465, January 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1042644726.

Berkes:1998:LTM

[BP98] István Berkes and Walter Philipp. Limit theorems for mixing sequences without rate assumptions. Annals of Probability, 26 (2):805–831, April 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022855651.

Benjamini:1995:MCM

[BPP95] Itai Benjamini, Robin Pemantle, and Yuval Peres. Martin capacity for Markov chains. Annals of Probability, 23(3):1332–1346, July 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988187.

Benjamini:1998:UPP

[BPP98] Itai Benjamini, Robin Pemantle, and Yuval Peres. Unpredictable paths and percolation. Annals of Probability, 26(3):1198–1211, July 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022855749.

Brassesco:1992:NPB

[Bra92] Stella Brassesco. A note on planar Brownian motion. Annals of Probability, 20(3):1498–1503, July 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989703.

Bradley:1993:EMC

[Bra93] Richard C. Bradley. Equivalent mixing conditions for random fields. Annals of Probability, 21(4):1921–1926, October 1993. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989004.

Bradley:1996:RBR

[Bra96] R. E. Bradley. Review of “Intersections of Random Walks: Probability and its Applications,” by Gregory F. Lawler. Annals of Probability, 24(3):1638–1642, July 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1065725196.

Brown:1990:EBE

[Bro90] Mark Brown. Error bounds for exponential approximations of geometric convolutions. Annals of Probability, 18(3):1388–1402, July 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990750.

Bryc:1992:LDP

[Bry92] Wlodzimierz Bryc. On the large deviation principle for stationary weakly dependent random fields. Annals of Probability, 20 (2):1004–1030, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989815.

Bruss:1990:CQS

[BS90] F. Thomas Bruss and Stephen M. Samuels. Conditions for quasistationarity of the Bayes rule in selection problems with an unknown number of rankable options. Annals of Probability, 18 (2):877–886, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990864.

Burdzy:1995:ILI

[BS95] Krzysztof Burdzy and Jaime San Martin. Iterated law of iterated logarithm. Annals of Probability, 23(4):1627–1643, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176987796.

Benjamini:1996:RWH

[BS96] Itai Benjamini and Oded Schramm. Random walks and harmonic functions on infinite planar graphs using square tilings. Annals of Probability, 24(3):1219–1238, July 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1065725179.

Bolthausen:1997:SAD

[BS97] Erwin Bolthausen and Uwe Schmock. On self-attracting ddimensional random walks. Annals of Probability, 25(2):531–572, April 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1024404411.

Bai:1998:NEO

[BS98] Z. D. Bai and Jack W. Silverstein. No eigenvalues outside the support of the limiting spectral distribution of large-dimensional sample covariance matrices. Annals of Probability, 26(1):316–345, January 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022855421.

Bai:1999:ESE

[BS99] Z. D. Bai and Jack W. Silverstein. Exact separation of eigenvalues of large dimensional sample covariance matrices. Annals of Probability, 27(3):1536–1555, July 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022677458.

Burdzy:1995:STL

[BT95] Krzysztof Burdzy and Ellen Toby. A Skorohod-type lemma and a decomposition of reflected Brownian motion. Annals of Probability, 23(2):586–604, April 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988280.

Bunge:1996:CSR

[Bun96] John Bunge. Composition semigroups and random stability. Annals of Probability, 24(3):1476–1489, July 1996. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1065725189.

Bunge:1997:NCD

[Bun97] John Bunge. Nested classes of C-decomposable laws. Annals of Probability, 25(1):215–229, January 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1024404286.

Burdzy:1990:NBM

[Bur90] Krzysztof Burdzy. On nonincrease of Brownian motion. Annals of Probability, 18(3):978–980, July 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990732.

Burkholder:1994:SDS

[Bur94] Donald L. Burkholder. Strong differential subordination and stochastic integration. Annals of Probability, 22(2):995–1025, April 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988738.

Busnello:1999:PAT

[Bus99] Barbara Busnello. A probabilistic approach to the twodimensional Navier–Stokes equations. Annals of Probability, 27(4): 1750–1780, October 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022874814.

Bertoin:1996:SW

[BW96a] Jean Bertoin and Wendelin Werner. Stable windings. Annals of Probability, 24(3):1269–1279, July 1996. CODEN APBYAE.

ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1065725181.

Burdzy:1996:NTP

[BW96b] Krzysztof Burdzy and Wendelin Werner. No triple point of planar Brownian motion is accessible. Annals of Probability, 24(1): 125–147, January 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1042644710.

Bai:1993:LSE

[BY93] Z. D. Bai and Y. Q. Yin. Limit of the smallest eigenvalue of a large dimensional sample covariance matrix. Annals of Probability, 21 (3):1275–1294, July 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989118.

Carmona:1997:MVB

[Car97] Philippe Carmona. The mean velocity of a Brownian motion in a random Lévy potential. Annals of Probability, 25(4):1774–1788, October 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1023481110.

Catoni:1992:RLD

[Cat92] Olivier Catoni. Rough large deviation estimates for simulated annealing: Application to exponential schedules. Annals of Probability, 20(3):1109–1146, July 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989682.

Chen:1992:SAL

[CC92a] Louis H. Y. Chen and K. P. Choi. Some asymptotic and large deviation results in Poisson approximation. Annals of Probability, 20 (4):1867–1876, October 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989532.

Csaki:1992:IIS

[CC92b] Endre Csaki and Miklós Csörgő. Inequalities for increments of stochastic processes and moduli of continuity. Annals of Probability, 20(2):1031–1052, April 1992. CODEN APBYAE.

ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989816.

Cerf:1999:FSS

[CC99] Raphaël Cerf and Emilio N. M. Cirillo. Finite size scaling in threedimensional bootstrap percolation. Annals of Probability, 27(4): 1837–1850, October 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022874817.

Cutler:1990:NNA

[CD90] Colleen D. Cutler and Donald A. Dawson. Nearest-neighbor analysis of a family of fractal distributions. Annals of Probability, 18 (1):256–271, January 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990948.

Cairoli:1995:OSB

[CD95] R. Cairoli and Robert C. Dalang. Optimal switching between two random walks. Annals of Probability, 23(4):1982–2013, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176987812.

Chow:1993:WEC

[CdlPT93] Y. S. Chow, V. H. de la Pena, and H. Teicher. Wald’s equation for a class of denormalized U-statistics. Annals of Probability, 21 (2):1151–1158, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989285.

Chen:1997:HDB

[CDM97] Zhen-Qing Chen, Richard Durrett, and Gang Ma. Holomorphic diffusions and boundary behavior of harmonic functions. Annals of Probability, 25(3):1103–1134, July 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1024404507.

Cerutti:1991:USD

[CEF91] M. Cristina Cerutti, Luis Escauriaza, and Eugene B. Fabes. Uniqueness for some diffusions with discontinuous coefficients. Annals of Probability, 19(2):525–537, April 1991. CODEN APBYAE.

ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990439.

Cepa:1998:PSM

[Cép98] Emmanuel Cépa. Problème de Skorohod multivoque. (French) [multivalent Skorohod problem]. Annals of Probability, 26(2):500– 532, April 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022855642.

Chavel:1991:FWS

[CFR91] Isaac Chavel, Edgar Feldman, and Jay Rosen. Fluctuations of the Wiener sausage for surfaces. Annals of Probability, 19(1):83–141, January 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990537.

Cox:1994:ETI

[CG94] J. T. Cox and Andreas Greven. Ergodic theorems for infinite systems of locally interacting diffusions. Annals of Probability, 22(2):833–853, April 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988732.

Carlson:1993:SDL

[CGST93] J. M. Carlson, E. R. Grannan, G. H. Swindle, and J. Tour. Singular diffusion limits of a class of reversible self-organizing particle systems. Annals of Probability, 21(3):1372–1393, July 1993. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989122.

Cuzick:1995:LLN

[CGZ95] Jack Cuzick, Evarist Gine, and Joel Zinn. Laws of large numbers for quadratic forms, maxima of products and truncated sums of i.i.d. random variables. Annals of Probability, 23(1):292–333, January 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988388.

Csorgo:1990:ATB

[CH90] Miklós Csörgő and Lajos Horváth. Asymptotic tail behavior of uniform multivariate empirical processes. Annals of Probability, 18 (4):1723–1738, October 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990643.

Courbage:1994:CKE

[CH94] M. Courbage and D. Hamdan. Chapman–Kolmogorov equation for non-Markovian shift-invariant measures. Annals of Probability, 22(3):1662–1677, July 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988618.

Chauvin:1991:PMS

[Cha91] Brigitte Chauvin. Product martingales and stopping lines for branching Brownian motion. Annals of Probability, 19(3): 1195–1205, July 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990340.

Chan:1994:DMV

[Cha94] Terence Chan. Dynamics of the McKean–Vlasov equation. Annals of Probability, 22(1):431–441, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988866.

Chen:1991:UCL

[Che91] Dongching Chen. A uniform Central Limit Theorem for nonuniform ϕ-mixing random fields. Annals of Probability, 19(2):636– 649, April 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990445.

Chen:1992:SPG

[Che92] Hwa-Nien Chen. On the stability of a population growth model with sexual reproduction on Z^{2}. Annals of Probability, 20(1): 232–285, January 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989926.

Chen:1993:LIL

[Che93a] Xia Chen. On the law of the iterated logarithm for independent Banach space valued random variables. Annals of Probability, 21 (4):1991–2011, October 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989008.

Cheng:1993:SMR

[Che93b] Ray Cheng. The spectral measure of a regular stationary random field with the weak or strong commutation property. Annals of Probability, 21(3):1263–1274, July 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989117.

Chen:1994:SPG

[Che94a] Hwa-Nien Chen. On the stability of a population growth model with sexual reproduction on Z^{d}, d ≥ 2. Annals of Probability, 22 (3):1195–1226, July 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988600.

Chen:1994:SLI

[Che94b] Xia Chen. On Strassen’s law of the iterated logarithm in Banach space. Annals of Probability, 22(2):1026–1043, April 1994. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988739.

Chen:1999:HOD

[Che99] Xia Chen. How often does a Harris recurrent Markov chain recur? Annals of Probability, 27(3):1324–1346, July 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1022677449.

Csorgo:1991:ADE

[CHM91] Sandor Csörgő, Erich Haeusler, and David M. Mason. The asymptotic distribution of extreme sums. Annals of Probability, 19 (2):783–811, April 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990451.

Carlen:1991:EIS

[CK91] Eric Carlen and Paul Kree. L^{P} estimates on iterated stochastic integrals. Annals of Probability, 19(1):354–368, January 1991. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990549.

Cvitanic:1996:BSD

[CK96] Jakša Cvitaniç and Ioannis Karatzas. Backward stochastic differential equations with reflection and Dynkin games. Annals of Probability, 24(4):2024–2056, October 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1041903216.

Choi:1997:SBP

[CK97] K. P. Choi and Michael J. Klass. Some best possible prophet inequalities for convex functions of sums of independent variates and unordered martingale difference sequences. Annals of Probability, 25(2):803–811, April 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1024404420.

Chang:1993:NLT

[CKS93] Fu-Chen Chang, J. H. B. Kemperman, and W. J. Studden. A normal limit theorem for moment sequences. Annals of Probability, 21(3):1295–1309, July 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989119.

Choulli:1998:MTA

[CKS98a] Tahir Choulli, Leszek Krawczyk, and Christophe Stricker. ⌉- martingales and their applications in mathematical finance. Annals of Probability, 26(2):853–876, April 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022855653.

Cvitanic:1998:BSD

[CKS98b] Jakša Cvitanić, Ioannis Karatzas, and H. Mete Soner. Backward stochastic differential equations with constraints on the gains-process. Annals of Probability, 26(4):1522–1551, October 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022855872.

Chen:1992:FRN

[CL92] Dayue Chen and Thomas M. Liggett. Finite reversible nearest particle systems in inhomogeneous and random environments. Annals of Probability, 20(1):152–173, January 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989922.

Cai:1993:LLN

[CL93] Haiyan Cai and Xiaolong Luo. Laws of large numbers for a cellular automaton. Annals of Probability, 21(3):1413–1426, July 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989124.

Chamayou:1994:TRW

[CL94] Jean-François Chamayou and Gerard Letac. A transient random walk on stochastic matrices with Dirichlet distributions. Annals of Probability, 22(1):424–430, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988865.

Chen:1991:SDF

[CM91] Hong Chen and Avi Mandelbaum. Stochastic discrete flow networks: Diffusion approximations and bottlenecks. Annals of Probability, 19(4):1463–1519, October 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990220.

Csorgo:1994:ADI

[CM94] Sandor Csörgő and David M. Mason. The asymptotic distribution of intermediate sums. Annals of Probability, 22(1):145–159, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988852.

Cranston:1996:SLL

[CM96] M. Cranston and T. S. Mountford. The Strong Law of Large Numbers for a Brownian polymer. Annals of Probability, 24 (3):1300–1323, July 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1065725183.

Comets:1998:LFR

[CMP98] Francis Comets, Mikhail Menshikov, and Serguei Popov. Lyapunov functions for random walks and strings in random environment. Annals of Probability, 26(4):1433–1445, October 1998. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1022855869.

Collet:1995:ALO

[CMS95] Pierre Collet, Servet Martinez, and Jaime San Martin. Asymptotic laws for one-dimensional diffusions conditioned to nonabsorption. Annals of Probability, 23(3):1300–1314, July 1995. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988185.

Collet:1999:RLT

[CMS99] Pierre Collet, Servet Martínez, and Jaime San Martín. Ratio limit theorems for a Brownian motion killed at the boundary of a Benedicks domain. Annals of Probability, 27(3):1160–1182, July 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022677443.

Cohn:1990:PNM

[CN90] Harry Cohn and Olle Nerman. On products of nonnegative matrices. Annals of Probability, 18(4):1806–1815, October 1990. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990650.

Cutland:1993:WSW

[CN93] Nigel Cutland and Siu-Ah Ng. The Wiener sphere and Wiener measure. Annals of Probability, 21(1):1–13, January 1993. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989390.

Cogburn:1990:DCP

[Cog90] Robert Cogburn. On direct convergence and periodicity for transition probabilities of Markov chains in random environments. Annals of Probability, 18(2):642–654, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990850.

Cogburn:1991:CLT

[Cog91] Robert Cogburn. On the Central Limit Theorem for Markov chains in random environments. Annals of Probability, 19(2):587– 604, April 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990442.

Collamore:1996:HPL

[Col96] Jeffrey F. Collamore. Hitting probabilities and large deviations. Annals of Probability, 24(4):2065–2078, October 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1041903218.

Cassandro:1999:TCO

[COP99] Marzio Cassandro, Enza Orlandi, and Pierre Picco. Typical configurations for one-dimensional random field Kac model. Annals of Probability, 27(3):1414–1467, July 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022677454.

Costantini:1991:DAC

[Cos91] C. Costantini. Diffusion approximation for a class of transport processes with physical reflection boundary conditions. Annals of Probability, 19(3):1071–1101, July 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990335.

Chang:1997:LHG

[CP97] Joseph T. Chang and Yuval Peres. Ladder heights, Gaussian random walks and the Riemann zeta function. Annals of Probability, 25(2):787–802, April 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1024404419.

Cacoullos:1994:VIE

[CPU94] T. Cacoullos, V. Papathanasiou, and S. A. Utev. Variational inequalities with examples and an application to the Central Limit Theorem. Annals of Probability, 22(3):1607–1618, July 1994. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988616.

Christofides:1990:MIM

[CS90] Tasos C. Christofides and Robert J. Serfling. Maximal inequalities for multidimensionally indexed submartingale arrays. Annals of Probability, 18(2):630–641, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990849.

Chaganty:1993:SLD

[CS93a] Narasinga Rao Chaganty and Jayaram Sethuraman. Strong large deviation and local limit theorems. Annals of Probability, 21 (3):1671–1690, July 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989136.

Csorgo:1993:SLT

[CS93b] Miklós Csörgő and Qi-Man Shao. Strong limit theorems for large and small increments of l^{p}-valued Gaussian processes. Annals of Probability, 21(4):1958–1990, October 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989007.

Chayes:1996:HLO

[CS96] L. Chayes and G. Swindle. Hydrodynamic limits for onedimensional particle systems with moving boundaries. Annals of Probability, 24(2):559–598, April 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1039639355.

Chen:1996:PTE

[CZ96] Z. Q. Chen and Z. Zhao. Potential theory for elliptic systems. Annals of Probability, 24(1):293–319, January 1996. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1042644718.

deAcosta:1996:RBT

[dA96] A. de Acosta. Review of “Probability Theory: An Analytic View,” by D. W. Stroock. Annals of Probability, 24(3):1643–1646, July 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1065725197.

deAcosta:1997:MDE

[dA97] A. de Acosta. Moderate deviations for empirical measures of Markov chains: lower bounds. Annals of Probability, 25(1): 259–284, January 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1024404288.

Dalang:1990:RTA

[Dal90] Robert C. Dalang. Randomization in the two-armed bandit problem. Annals of Probability, 18(1):218–225, January 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990946.

Daley:1999:HIL

[Dal99] D. J. Daley. The Hurst index of long-range dependent renewal processes. Annals of Probability, 27(4):2035–2041, October 1999. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1022874827.

Darling:1995:CGM

[Dar95] R. W. R. Darling. Constructing gamma-martingales with prescribed limit, using backwards SDE. Annals of Probability, 23 (3):1234–1261, July 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988182.

Davis:1996:WLP

[Dav96] Burgess Davis. Weak limits of perturbed random walks and the equation Y_{t} = B_{t} + α sup{Y_{s}: s ≤ t} + β inf{Y_{s}: s ≤ t}. Annals of Probability, 24(4):2007–2023, October 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1041903215.

Day:1992:CES

[Day92] Martin V. Day. Conditional exits for small noise diffusions with characteristic boundary. Annals of Probability, 20(3):1385–1419, July 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989696.

Delmas:1999:CRS

[DD99] Jean-François Delmas and Jean-Stéphane Dhersin. Characterization of G-regularity for super-Brownian motion and consequences for parabolic partial differential equations. Annals of Probability, 27(2):731–750, April 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022677384.

DeValk:1993:HSR

[De 93] Vincent De Valk. Hilbert space representations of m-dependent processes. Annals of Probability, 21(3):1550–1570, July 1993. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989130.

Deheuvels:1996:SLB

[DE96] Paul Deheuvels and John H. J. Einmahl. On the strong limiting behavior of local functionals of empirical processes based upon censored data. Annals of Probability, 24(1):504–525, January 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1042644729.

DeBlassie:1990:FET

[DeB90] R. Dante DeBlassie. The first exit time of a two-dimensional symmetric stable process from a wedge. Annals of Probability, 18 (3):1034–1070, July 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990735.

DeBlassie:1996:BMW

[DeB96] R. Dante DeBlassie. Brownian motion in a wedge with variable reflection: existence and uniqueness. Annals of Probability, 24 (1):148–181, January 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1042644711.

DeBlassie:1999:APK

[DeB99] R. Dante DeBlassie. The adjoint process of killed reflected Brownian motion in a cone and applications. Annals of Probability, 27 (4):1679–1737, October 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022874812.

Deheuvels:1997:SLL

[Deh97] Paul Deheuvels. Strong laws for local quantile processes. Annals of Probability, 25(4):2007–2054, October 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1023481119.

Delmas:1999:PPS

[Del99] Jean-François Delmas. Path properties of superprocesses with a general branching mechanism. Annals of Probability, 27(3): 1099–1134, July 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022677441.

Dembo:1997:IIC

[Dem97] Amir Dembo. Information inequalities and concentration of measure. Annals of Probability, 25(2):927–939, April 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1024404424.

Dermoune:1999:PIS

[Der99] Azzouz Dermoune. Probabilistic interpretation of sticky particle model. Annals of Probability, 27(3):1357–1367, July 1999. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1022677451.

Deuschel:1994:ADA

[Deu94] Jean-Dominique Deuschel. Algebraic L^{2} decay of attractive critical processes on the lattice. Annals of Probability, 22(1):264–283, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988859.

Dupuis:1991:LDM

[DEW91] Paul Dupuis, Richard S. Ellis, and Alan Weiss. Large deviations for Markov processes with discontinuous statistics, i: General upper bounds. Annals of Probability, 19(3):1280–1297, July 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990344.

Diaconis:1990:SST

[DF90] Persi Diaconis and James Allen Fill. Strong stationary times via a new form of duality. Annals of Probability, 18(4):1483–1522, October 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990628.

deFitte:1997:TEP

[dF97] Paul Raynaud de Fitte. Théorème ergodique ponctuel et lois fortes des grands nombres pour des points aléatoires d’un espace métrique à courbure négative. (French) [Pointwise ergodic theorem and Strong Laws of Large Numbers for random points in a metric space with negative curvature]. Annals of Probability, 25(2):738–766, April 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1024404417.

Dalang:1998:SWE

[DF98] Robert C. Dalang and N. E. Frangos. The stochastic wave equation in two spatial dimensions. Annals of Probability, 26(1): 187–212, January 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022855416.

Dawson:1998:CDC

[DFL98] Donald A. Dawson, Klaus Fleischmann, and Guillaume Leduc. Continuous dependence of a class of superprocesses on branching parameters and applications. Annals of Probability, 26(2):562–601, April 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022855644.

Dawson:1995:SSB

[DFLM95] Donald A. Dawson, Klaus Fleischmann, Yi Li, and Carl Mueller. Singularity of super-Brownian local time at a point catalyst. Annals of Probability, 23(1):37–55, January 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988375.

Dubins:1996:DSF

[DFST96] Lester Dubins, Jacob Feldman, Meir Smorodinsky, and Boris Tsirelson. Decreasing sequences of σ-fields and a measure change for Brownian motion. Annals of Probability, 24(2):882–904, April 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1039639367.

DeSantis:1999:BPF

[DG99] E. De Santis and A. Gandolfi. Bond percolation in frustrated systems. Annals of Probability, 27(4):1781–1808, October 1999. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1022874815.

Davis:1995:PPP

[DH95] Richard A. Davis and Tailen Hsing. Point process and partial sum convergence for weakly dependent random variables with infinite variance. Annals of Probability, 23(2):879–917, April 1995. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988294.

Dou:1996:ERR

[DH96] Carl Dou and Martin Hildebrand. Enumeration and random random walks on finite groups. Annals of Probability, 24(2):987–1000, April 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1039639374.

Dupuis:1993:SOR

[DI93] Paul Dupuis and Hitoshi Ishii. SDEs with oblique reflection on nonsmooth domains. Annals of Probability, 21(1):554–580, January 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989415.

Dinwoodie:1991:NUB

[Din91] I. H. Dinwoodie. A note on the upper bound for i.i.d. large deviations. Annals of Probability, 19(4):1732–1736, October 1991. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990231.

Dinwoodie:1993:ILD

[Din93] I. H. Dinwoodie. Identifying a large deviation rate function. Annals of Probability, 21(1):216–231, January 1993. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989402.

Dupuis:1990:VSA

[DIS90] Paul Dupuis, Hitoshi Ishii, and H. Mete Soner. A viscosity solution approach to the asymptotic analysis of queueing systems. Annals of Probability, 18(1):226–255, January 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990947.

Donnelly:1992:WCP

[DJ92] Peter Donnelly and Paul Joyce. Weak convergence of population genealogical processes to the coalescent with ages. Annals of Probability, 20(1):322–341, January 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989929.

Dabrowski:1994:SLA

[DJ94] Andre Robert Dabrowski and Adam Jakubowski. Stable limits for associated random variables. Annals of Probability, 22 (1):1–16, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988845.

Dembo:1991:SLTb

[DK91a] Amir Dembo and Samuel Karlin. Strong limit theorems of empirical distributions for large segmental exceedances of partial sums of Markov variables. Annals of Probability, 19(4):1756–1767, October 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990233.

Dembo:1991:SLTa

[DK91b] Amir Dembo and Samuel Karlin. Strong limit theorems of empirical functionals for large exceedances of partial sums of i.i.d. variables. Annals of Probability, 19(4):1737–1755, October 1991. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990232.

Djehiche:1995:RFS

[DK95] Boualem Djehiche and Ingemar Kaj. The rate function for some measure-valued jump processes. Annals of Probability, 23(3): 1414–1438, July 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988190.

Donnelly:1996:CRF

[DK96] Peter Donnelly and Thomas G. Kurtz. A countable representation of the Fleming–Viot measure-valued diffusion. Annals of Probability, 24(2):698–742, April 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1039639359.

Dynkin:1997:NLA

[DK97a] E. B. Dynkin and S. E. Kuznetsov. Natural linear additive functionals of superprocesses. Annals of Probability, 25(2):640–661, April 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1024404414.

Dynkin:1997:NPP

[DK97b] E. B. Dynkin and S. E. Kuznetsov. Nonlinear parabolic P.D.E. and additive functionals of superdiffusions. Annals of Probability, 25(2):662–701, April 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1024404415.

Donnelly:1999:PRM

[DK99] Peter Donnelly and Thomas G. Kurtz. Particle representations for measure-valued population models. Annals of Probability, 27 (1):166–205, January 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022677258.

Dembo:1994:CPS

[DKZ94a] Amir Dembo, Samuel Karlin, and Ofer Zeitouni. Critical phenomena for sequence matching with scoring. Annals of Probability, 22 (4):1993–2021, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988492.

Dembo:1994:LDM

[DKZ94b] Amir Dembo, Samuel Karlin, and Ofer Zeitouni. Limit distribution of maximal non-aligned two-sequence segmental score. Annals of Probability, 22(4):2022–2039, October 1994. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988493.

Deheuvels:1994:NSC

[DL94] Paul Deheuvels and Mikhail A. Lifshits. Necessary and sufficient conditions for the Strassen law of the iterated logarithm in nonuniform topologies. Annals of Probability, 22(4):1838–1856, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988486.

deLara:1995:GSP

[dL95] M. Cohen de Lara. Geometric and symmetry properties of a nondegenerate diffusion process. Annals of Probability, 23(4): 1557–1604, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176987794.

Dhersin:1998:KTS

[DL98] Jean-Stéphane Dhersin and Jean-François Le Gall. Kolmogorov’s test for super-Brownian motion. Annals of Probability, 26(3): 1041–1056, July 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022855744.

delaCal:1990:LDM

[dlC90] Jesus de la Cal. On the limit distribution of multiplicative functions with values in the interval [−1, 1]. Annals of Probability, 18(2):901–904, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990866.

Dawson:1995:SMV

[DLM95] D. Dawson, Y. Li, and C. Mueller. The support of measurevalued branching processes in a random environment. Annals of Probability, 23(4):1692–1718, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176987799.

delaPena:1992:DKI

[dlP92] Victor H. de la Pena. Decoupling and Khintchine’s inequalities for U-statistics. Annals of Probability, 20(4):1877–1892, October 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989533.

delaPena:1999:GCE

[dlP99] Victor H. de la Peña. A general class of exponential inequalities for martingales and ratios. Annals of Probability, 27(1):537–564, January 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022677271.

delaPena:1994:OMB

[dlPK94] Victor H. de la Pena and Michael J. Klass. Order-of-magnitude bounds for expectations involving quadratic forms. Annals of Probability, 22(2):1044–1077, April 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988740.

delaPena:1995:DIT

[dlPMS95] Victor H. de la Pena and S. J. Montgomery-Smith. Decoupling inequalities for the tail probabilities of multivariate U-statistics. Annals of Probability, 23(2):806–816, April 1995. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988291.

delaPena:1994:CDI

[dlPMSS94] V. H. de la Pena, S. J. Montgomery-Smith, and Jerzy Szulga. Contraction and decoupling inequalities for multilinear forms and U-statistics. Annals of Probability, 22(4):1745–1765, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988481.

Deheuvels:1990:BKT

[DM90a] Paul Deheuvels and David M. Mason. Bahadur–Kiefer-type processes. Annals of Probability, 18(2):669–697, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990852.

Deheuvels:1990:NFL

[DM90b] Paul Deheuvels and David M. Mason. Nonstandard functional laws of the iterated logarithm for tail empirical and quantile processes. Annals of Probability, 18(4):1693–1722, October 1990. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990642.

Deheuvels:1992:FLI

[DM92] Paul Deheuvels and David M. Mason. Functional laws of the iterated logarithm for the increments of empirical and quantile processes. Annals of Probability, 20(3):1248–1287, July 1992. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989691.

Deheuvels:1994:FLI

[DM94] Paul Deheuvels and David M. Mason. Functional laws of the iterated logarithm for local empirical processes indexed by sets. Annals of Probability, 22(3):1619–1661, July 1994. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988617.

Deheuvels:1995:FNE

[DM95] Paul Deheuvels and David M. Mason. On the fractal nature of empirical increments. Annals of Probability, 23(1):355–387, January 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988390.

Dalang:1996:NCB

[DM96] Robert C. Dalang and T. Mountford. Nondifferentiability of curves on the Brownian sheet. Annals of Probability, 24(1):182–195, January 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1042644712.

Davis:1999:MPN

[DM99] Richard A. Davis and Thomas Mikosch. The maximum of the periodogram of a non-Gaussian sequence. Annals of Probability, 27(1):522–536, January 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022677270.

deMonvel:1999:NED

[dMK99] Anne Boutet de Monvel and Alexei Khorunzhy. On the norm and eigenvalue distribution of large random matrices. Annals of Probability, 27(2):913–944, April 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022677390.

Dilworth:1993:DVV

[DMS93] S. J. Dilworth and S. J. Montgomery-Smith. The distribution of vector-valued Rademacher series. Annals of Probability, 21(4): 2046–2052, October 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989010.

Down:1995:EUE

[DMT95] D. Down, S. P. Meyn, and R. L. Tweedie. Exponential and uniform ergodicity of Markov processes. Annals of Probability, 23(4): 1671–1691, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176987798.

Donati-Martin:1997:SBF

[DMY97] C. Donati-Martin and M. Yor. Some Brownian functionals and their laws. Annals of Probability, 25(3):1011–1058, July 1997. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1024404505.

Durrett:1994:PSR

[DN94] R. Durrett and C. Neuhauser. Particle systems and reactiondiffusion equations. Annals of Probability, 22(1):289–333, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988861.

Donnelly:1991:WCM

[Don91] Peter Donnelly. Weak convergence to a Markov chain with an entrance boundary: Ancestral processes in population genetics. Annals of Probability, 19(3):1102–1117, July 1991. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990336.

Doney:1996:ILP

[Don96] R. A. Doney. Increase of Lévy processes. Annals of Probability, 24(2):961–970, April 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1039639372.

Diebolt:1996:DMS

[DP96] Jean Diebolt and Christian Posse. On the density of the maximum of smooth Gaussian processes. Annals of Probability, 24 (3):1104–1129, July 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1065725176.

Darling:1997:BSR

[DP97] R. W. R. Darling and Etienne Pardoux. Backwards SDE with random terminal time and applications to semilinear elliptic PDE.

Annals of Probability, 25(3):1135–1159, July 1997. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1024404508.

Dawson:1998:LTB

[DP98] Donald A. Dawson and Edwin A. Perkins. Long-time behavior and coexistence in a mutually catalytic branching model. Annals of Probability, 26(3):1088–1138, July 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022855746.

DeMasi:1994:SPW

[DPPV94] A. De Masi, A. Pellegrinotti, E. Presutti, and M. E. Vares. Spatial patterns when phases separate in an interacting particle system. Annals of Probability, 22(1):334–371, January 1994. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988862.

Davis:1991:EMA

[DR91] Richard A. Davis and Sidney I. Resnick. Extremes of moving averages of random variables with finite endpoint. Annals of Probability, 19(1):312–328, January 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990546.

Deuschel:1998:OTL

[DR98] Jean-Dominique Deuschel and Jay Rosen. Occupation time large deviations for critical branching Brownian motion, super-Brownian motion and related processes. Annals of Probability, 26(2):602–643, April 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022855645.

deFMarques:1992:QIP

[dS92] Mauro S. de F. Marques and Luiz San Martin. Quasi-invariance of product measures under Lie group perturbations: Fisher information and L^{2}-differentiability. Annals of Probability, 20(3): 1420–1435, July 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989697.

Durrett:1993:FRT

[DS93] Richard Durrett and Jeffrey E. Steif. Fixation results for threshold voter systems. Annals of Probability, 21(1):232–247, January 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989403.

Durrett:1995:IPC

[DS95] Rick Durrett and Rinaldo Schinazi. Intermediate phase for the contact process on a tree. Annals of Probability, 23(2):668–673, April 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988283.

Diaconis:1993:CTR

[DSC93] Persi Diaconis and Laurent Saloff-Coste. Comparison techniques for random walk on finite groups. Annals of Probability, 21(4): 2131–2156, October 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989013.

delTio:1991:CRM

[dTB91] P. Rodriguez del Tio and M. C. Valsero Blanco. A characterization of reversible Markov chains by a rotational representation. Annals of Probability, 19(2):605–608, April 1991. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990443.

Dudley:1992:FDV

[Dud92] R. M. Dudley. Frechet differentiability, p-variation and uniform Donsker classes. Annals of Probability, 20(4):1968–1982, October 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989537.

Duheille:1998:RVH

[Duh98] F. Duheille. On the range of R^{2} or R^{3}-valued harmonic morphisms. Annals of Probability, 26(1):308–315, January 1998. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1022855420.

Durrett:1993:RDR

[Dur93] Rick Durrett. Review: D. Revuz, M. Yor, Continuous Martingales and Brownian Motion. Annals of Probability, 21(1):588–589, January 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989417.

Dalang:1992:SMP

[DW92] Robert C. Dalang and John B. Walsh. The sharp Markov property of Lévy sheets. Annals of Probability, 20(2):591–626, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989793.

Dupuis:1994:LFS

[DW94] Paul Dupuis and Ruth J. Williams. Lyapunov functions for semimartingale reflecting Brownian motions. Annals of Probability, 22(2):680–702, April 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988725.

Dynkin:1991:BPS

[Dyn91] E. B. Dynkin. Branching particle systems and superprocesses. Annals of Probability, 19(3):1157–1194, July 1991. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990339.

Dynkin:1992:SPN

[Dyn92] E. B. Dynkin. Superdiffusions and parabolic nonlinear differential equations. Annals of Probability, 20(2):942–962, April 1992. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989812.

Dynkin:1993:SPD

[Dyn93] E. B. Dynkin. Superprocesses and partial differential equations. Annals of Probability, 21(3):1185–1262, July 1993. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989116.

Dinwoodie:1992:LDE

[DZ92] I. H. Dinwoodie and S. L. Zabell. Large deviations for exchangeable random vectors. Annals of Probability, 20(3):1147–1166, July 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989683.

Deuschel:1995:LCD

[DZ95] Jean-Dominique Deuschel and Ofer Zeitouni. Limiting curves for i.i.d. records. Annals of Probability, 23(2):852–878, April 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988293.

Eizenberg:1993:LDM

[EF93] Alexander Eizenberg and Mark Freidlin. Large deviations for Markov processes corresponding to PDE systems. Annals of Probability, 21(2):1015–1044, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989280.

Evans:1996:CFS

[EF96] Steven N. Evans and Klaus Fleischmann. Cluster formation in a stepping-stone model with continuous, hierarchically structured sites. Annals of Probability, 24(4):1926–1952, October 1996. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1041903211.

Ethier:1993:TFF

[EG93] S. N. Ethier and R. C. Griffiths. The transition function of a Fleming–Viot process. Annals of Probability, 21(3):1571–1590, July 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989131.

Elton:1992:FPD

[EH92] J. Elton and T. P. Hill. Fusions of a probability distribution. Annals of Probability, 20(1):421–454, January 1992. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989936.

Einmahl:1991:ASB

[Ein91] Uwe Einmahl. On the almost sure behavior of sums of IID random variables in Hilbert space. Annals of Probability, 19(3):1227–1263, July 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990342.

Einmahl:1992:BWE

[Ein92] John H. J. Einmahl. The A.S. behavior of the weighted empirical process and the LIL for the weighted tail empirical process. Annals of Probability, 20(2):681–695, April 1992. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989800.

Einmahl:1993:TGL

[Ein93] Uwe Einmahl. Toward a general law of the iterated logarithm in Banach space. Annals of Probability, 21(4):2012–2045, October 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989009.

Einmahl:1995:CSP

[Ein95] Uwe Einmahl. On the cluster set problem for the generalized law of the iterated logarithm in Euclidean space. Annals of Probability, 23(2):817–851, April 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988292.

Einmahl:1996:SEP

[Ein96] John H. J. Einmahl. A short and elementary proof of the main Bahadur–Kiefer theorem. Annals of Probability, 24(1):526–531, January 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1042644730.

Eisenbaum:1993:NSC

[EK93] Nathalie Eisenbaum and Haya Kaspi. A necessary and sufficient condition for the Markov property of the local time process. Annals of Probability, 21(3):1591–1598, July 1993. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989132.

Ethier:1998:CET

[EK98] S. N. Ethier and Thomas G. Kurtz. Coupling and ergodic theorems for Fleming–Viot processes. Annals of Probability, 26(2):533–561, April 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022855643.

ElKaroui:1997:RSB

[EKP^{+}97] N. El Karoui, C. Kapoudjian, E. Pardoux, S. Peng, and M. C. Quenez. Reflected solutions of backward SDE’s, and related obstacle problems for PDE’s. Annals of Probability, 25(2):702–737, April 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1024404416.

Eloranta:1990:CMP

[Elo90] Kari Eloranta. α-congruence for Markov processes. Annals of Probability, 18(4):1583–1601, October 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990634.

Einmahl:1994:UCT

[EM94] Uwe Einmahl and David M. Mason. A universal Chung-type law of the iterated logarithm. Annals of Probability, 22(4):1803–1825, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988484.

Einmahl:1996:SUR

[EM96] Uwe Einmahl and David M. Mason. Some universal results on the behavior of increments of partial sums. Annals of Probability, 24(3):1388–1407, July 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1065725186.

Enchev:1993:NTW

[Enc93a] Ognian Enchev. Nonlinear transformations on the Wiener space. Annals of Probability, 21(4):2169–2188, October 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989015.

Enchev:1993:PNF

[Enc93b] Ognian Enchev. Pathwise nonlinear filtering on abstract Wiener spaces. Annals of Probability, 21(3):1728–1754, July 1993. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989139.

Enchev:1998:WNI

[Enc98] Ognian B. Enchev. White noise indexed by loops. Annals of Probability, 26(3):985–999, July 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022855741.

Evans:1995:ESI

[EP95] Steven N. Evans and Edwin A. Perkins. Explicit stochastic integral representations for historical functionals. Annals of Probability, 23(4):1772–1815, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176987803.

Englander:1999:CSP

[EP99] János Engländer and Ross G. Pinsky. On the construction and support properties of measure-valued diffusions on D ⊆ R^{d} with spatially dependent branching. Annals of Probability, 27(2):684– 730, April 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022677383.

Erhardsson:1999:CPA

[Erh99] Torkel Erhardsson. Compound Poisson approximation for Markov chains using Stein’s method. Annals of Probability, 27(1):565–596, January 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022677272.

Enchev:1993:RTW

[ES93] O. Enchev and D. W. Stroock. Rademacher’s theorem for Wiener functionals. Annals of Probability, 21(1):25–33, January 1993. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989392.

Evans:1992:PNS

[Eva92] Steven N. Evans. Polar and nonpolar sets for a tree indexed process. Annals of Probability, 20(2):579–590, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989792.

Fang:1999:IPF

[Fan99] Shizan Fang. Integration by parts formula and logarithmic Sobolev inequality on the path space over loop groups. Annals of Probability, 27(2):664–683, April 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022677382.

Farahmand:1990:ANL

[Far90] Kambiz Farahmand. On the average number of level crossings of a random trigonometric polynomial. Annals of Probability, 18 (3):1403–1409, July 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990751.

Feldman:1996:CMP

[Fel96] J. Feldman. ϵ-close measures producing nonisomorphic filtrations. Annals of Probability, 24(2):912–915, April 1996. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1039639369.

Feng:1994:LDE

[Fen94] Shui Feng. Large deviations for empirical process of mean-field interacting particle system with unbounded jumps. Annals of Probability, 22(4):2122–2151, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988496.

Fernique:1990:RFA

[Fer90a] X. Fernique. Régularité de fonctions aléatoires gaussiennes à valeurs vectorielles. (French) [Regularity of random Gaussian functions with vector values]. Annals of Probability, 18(4):1739–1745, October 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990644.

Ferrari:1990:ESS

[Fer90b] Pablo A. Ferrari. Ergodicity for spin systems with stirrings. Annals of Probability, 18(4):1523–1538, October 1990. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990629.

Fernique:1992:RFA

[Fer92] X. Fernique. Sur la regularité des fonctions aléatoires d’Ornstein– Uhlenbeck à valeurs dans l_{p}, p ∈ [1, ∞[. (French) [On the regularity of random Ornstein–Uhlenbeck functions with values in l_{p}, p ∈ [1, ∞[]. Annals of Probability, 20(3):1441–1449, July 1992. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989699.

Ferrari:1994:CFA

[FF94] P. A. Ferrari and L. R. G. Fontes. Current fluctuations for the asymmetric simple exclusion process. Annals of Probability, 22 (2):820–832, April 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988731.

Fitzsimmons:1992:DHT

[FG92a] P. J. Fitzsimmons and R. K. Getoor. On the distribution of the Hilbert transform of the local time of a symmetric Lévy process. Annals of Probability, 20(3):1484–1497, July 1992. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989702.

Fleischmann:1992:LSM

[FG92b] Klaus Fleischmann and Andreas Greven. Localization and selection in a mean field branching random walk in a random environment. Annals of Probability, 20(4):2141–2163, October 1992. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989543.

Follmer:1997:EMS

[FG97] Hans Föllmer and Nina Gantert. Entropy minimization and Schrödinger processes in infinite dimensions. Annals of Probability, 25(2):901–926, April 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1024404423.

Ferrari:1994:EWT

[FGL94] P. A. Ferrari, A. Galves, and C. Landim. Exponential waiting time for a big gap in a one-dimensional zero-range process. Annals of Probability, 22(1):284–288, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988860.

Funaki:1991:HLO

[FHU91] T. Funaki, K. Handa, and K. Uchiyama. Hydrodynamic limit of one-dimensional exclusion processes with speed change. Annals of Probability, 19(1):245–265, January 1991. CODEN APBYAE.

ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990543.

Finner:1992:GHI

[Fin92] Helmut Finner. A generalization of Holder’s inequality and some probability inequalities. Annals of Probability, 20(4):1893–1901, October 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989534.

Fisch:1992:COD

[Fis92a] Robert Fisch. Clustering in the one-dimensional three-color cyclic cellular automaton. Annals of Probability, 20(3):1528–1548, July 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989705.

Fisher:1992:LIL

[Fis92b] Evan Fisher. On the law of the iterated logarithm for martingales. Annals of Probability, 20(2):675–680, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989799.

Fitzsimmons:1990:OST

[Fit90] P. J. Fitzsimmons. On ordered stopping times of a Markov process. Annals of Probability, 18(4):1619–1622, October 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990636.

Fitzsimmons:1997:ACS

[Fit97] P. J. Fitzsimmons. Absolute continuity of symmetric diffusions. Annals of Probability, 25(1):230–258, January 1997. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1024404287.

Fleming:1992:ASE

[FJ92] W. H. Fleming and M. R. James. Asymptotic series and exit time probabilities. Annals of Probability, 20(3):1369–1384, July 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989695.

Fitzsimmons:1992:CDC

[FK92] P. J. Fitzsimmons and Mamoru Kanda. On Choquet’s dichotomy of capacity for Markov processes. Annals of Probability, 20(1): 342–349, January 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989930.

Fannjiang:1997:MAH

[FK97] Albert Fannjiang and Tomasz Komorowski. A martingale approach to homogenization of unbounded random flows. Annals of Probability, 25(4):1872–1894, October 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1023481115.

Fannjiang:1999:IPD

[FK99] Albert Fannjiang and Tomasz Komorowski. An invariance principle for diffusion in turbulence. Annals of Probability, 27(2):751– 781, April 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.

aop/1022677385. See correction [FK02].

Fannjiang:2002:CIP

[FK02] Albert C. Fannjiang and Tomasz Komorowski. Correction: An invariance principle for diffusion in turbulence. Annals of Probability, 30(1):480–482, January 2002. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1020107777. See [FK99].

Ferrari:1995:EQS

[FKMP95] P. A. Ferrari, H. Kesten, S. Martinez, and P. Picco. Existence of quasi-stationary distributions. A renewal dynamical approach. Annals of Probability, 23(2):501–521, April 1995. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988277.

Ferrari:1991:MST

[FKS91] P. A. Ferrari, C. Kipnis, and E. Saada. Microscopic structure of travelling waves in the asymmetric simple exclusion process. Annals of Probability, 19(1):226–244, January 1991. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990542.

Flandoli:1996:SFN

[Fla96] Franco Flandoli. Stochastic flows for nonlinear second-order parabolic SPDE. Annals of Probability, 24(2):547–558, April 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1039639354.

Falk:1993:MCR

[FM93] Michael Falk and Frank Marohn. Von Mises conditions revisited. Annals of Probability, 21(3):1310–1328, July 1993. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989120.

Fontes:1993:OPS

[Fon93] Luiz Renato G. Fontes. An ordered phase with slow decay of correlations in continuum 1/r^{2} Ising models. Annals of Probability, 21(3):1394–1412, July 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989123.

Fernandez:1997:GSN

[FP97] R. Fernández and C.-E. Pfister. Global specifications and nonquasilocality of projections of Gibbs measures. Annals of Probability, 25(3):1284–1315, July 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1024404514.

Feldman:1993:SRS

[FR93] Raisa Epstein Feldman and Svetlozar Rachev. U-statistics of random-size samples and limit theorems for systems of Markovian particles with non-Poisson initial distributions. Annals of Probability, 21(4):1927–1945, October 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989005.

Freidlin:1991:CRD

[Fre91] Mark Freidlin. Coupled reaction-diffusion equations. Annals of Probability, 19(1):29–57, January 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990535.

Fleming:1997:APE

[FS97] Wendell H. Fleming and Shuenn-Jyi Sheu. Asymptotics for the principal eigenvalue and eigenfunction of a nearly first-order operator with large potential. Annals of Probability, 25(4):1953–1994, October 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1023481117.

Feldman:1996:DSF

[FT96] Jacob Feldman and Boris Tsirelson. Decreasing sequences of σfields and a measure change for Brownian motion. II. Annals of Probability, 24(2):905–911, April 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1039639368.

Fukai:1996:PKT

[FU96] Yasunari Fukai and Kôhei Uchiyama. Potential kernel for twodimensional random walk. Annals of Probability, 24(4):1979–1992, October 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1041903213.

Freidlin:1992:RDE

[FW92] Mark I. Freidlin and Alexander D. Wentzell. Reaction-diffusion equations with randomly perturbed boundary conditions. Annals of Probability, 20(2):963–986, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989813.

Freidlin:1993:DPG

[FW93] Mark I. Freidlin and Alexander D. Wentzell. Diffusion processes on graphs and the averaging principle. Annals of Probability, 21(4): 2215–2245, October 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989018.

Freidlin:1998:RPN

[FW98] Mark Freidlin and Matthias Weber. Random perturbations of nonlinear oscillators. Annals of Probability, 26(3):925–967, July 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022855739.

Gadidov:1998:SLL

[Gad98] Anda Gadidov. Strong Law of Large Numbers for multilinear forms. Annals of Probability, 26(2):902–923, April 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1022855655.

Gantert:1993:ISL

[Gan93] Nina Gantert. An inversion of Strassen’s law of the iterated logarithm for small time. Annals of Probability, 21(2):1045–1049, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989281.

Gantert:1998:FER

[Gan98] Nina Gantert. Functional Erdős–Rényi laws for semiexponential random variables. Annals of Probability, 26(3):1356–1369, July 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022855755.

Greven:1994:LDR

[GdH94] Andreas Greven and Frank den Hollander. Large deviations for a random walk in random environment. Annals of Probability, 22 (3):1381–1428, July 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988607.

Gentz:1996:CLT

[Gen96] Barbara Gentz. A central limit theorem for the overlap in the Hopfield model. Annals of Probability, 24(4):1809–1841, October 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1041903207.

Georgii:1993:LDM

[Geo93] Hans-Otto Georgii. Large deviations and maximum entropy principle for interacting random fields on Z^{d}. Annals of Probability, 21 (4):1845–1875, October 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989002.

Gravner:1996:FPT

[GG96] Janko Gravner and David Griffeath. First passage times for threshold growth dynamics on Z^{2}. Annals of Probability, 24(4): 1752–1778, October 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1041903205.

Gine:1999:LIL

[GG99] Evarist Giné and Armelle Guillou. Laws of the iterated logarithm for censored data. Annals of Probability, 27(4):2042–2067, October 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022874828.

Gine:1997:WSS

[GGM97] Evarist Giné, Friedrich Götze, and David M. Mason. When is the Student t-statistic asymptotically standard normal? Annals of Probability, 25(3):1514–1531, July 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1024404523.

Gotze:1990:LLT

[GH90] F. Gotze and C. Hipp. Local limit theorems for sums of finite range potentials of a Gibbsian random fieldo. Annals of Probability, 18 (2):810–828, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990859.

Gilat:1992:OSR

[GH92] David Gilat and T. P. Hill. One-sided refinements of the Strong Law of Large Numbers and the Glivenko–Cantelli theorem. Annals of Probability, 20(3):1213–1221, July 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989688.

Goel:1994:ADB

[GH94] Prem K. Goel and Peter Hall. On the average difference between concomitants and order statistics. Annals of Probability, 22(1): 126–144, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988851.

Groeneboom:1999:IBM

[GJW99] Piet Groeneboom, Geurt Jongbloed, and Jon A. Wellner. Integrated Brownian motion, conditioned to be positive. Annals of Probability, 27(3):1283–1303, July 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022677447.

Griffin:1991:SEL

[GK91] Philip Griffin and James Kuelbs. Some extensions of the LIL via self-normalizations. Annals of Probability, 19(1):380–395, January 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990551.

Grigoryan:1998:RFB

[GK98] Alexander Grigor’yan and Mark Kelbert. Range of fluctuation of Brownian motion on a complete Riemannian manifold. Annals of Probability, 26(1):78–111, January 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022855412.

Gutmann:1991:EPM

[GKRS91] Sam Gutmann, J. H. B. Kemperman, J. A. Reeds, and L. A. Shepp. Existence of probability measures with given marginals. Annals of Probability, 19(4):1781–1797, October 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990236.

Gu:1990:FLI

[GL90] Ming Gao Gu and Tze Leung Lai. Functional laws of the iterated logarithm for the product-limit estimator of a distribution function under random censorship or truncation. Annals of Probability, 18 (1):160–189, January 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990943.

Glover:1990:RMJ

[Glo90] Joseph Glover. Review: M. J. Sharpe, General Theory of Markov Processes. Annals of Probability, 18(4):1823–1827, October 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990652.

Glover:1991:SGT

[Glo91] Joseph Glover. Symmetry groups and translation invariant representations of Markov processes. Annals of Probability, 19(2):562– 586, April 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990441.

Glover:1990:SFM

[GM90] Joseph Glover and Joanna Mitro. Symmetries and functions of Markov processes. Annals of Probability, 18(2):655–668, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990851.

Griffin:1992:PRW

[GM92] Philip S. Griffin and Terry R. McConnell. On the position of a random walk at the time of first exit from a sphere. Annals of Probability, 20(2):825–854, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989808.

Griffin:1994:GRF

[GM94] Philip S. Griffin and Terry R. McConnell. Gambler’s ruin and the first exit position of random walk from large spheres. Annals of Probability, 22(3):1429–1472, July 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988608.

Griffin:1995:BOM

[GM95] Philip S. Griffin and Terry R. McConnell. L^{p}-boundedness of the overshoot in multidimensional renewal theory. Annals of Probability, 23(4):2022–2056, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176987814.

Glynn:1996:LBS

[GM96] Peter W. Glynn and Sean P. Meyn. A Liapounov bound for solutions of the Poisson equation. Annals of Probability, 24(2):916– 931, April 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1039639370.

Graham:1997:SPA

[GM97] Carl Graham and Sylvie Méléard. Stochastic particle approximations for generalized Boltzmann models and convergence estimates. Annals of Probability, 25(1):115–132, January 1997. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1024404281.

Gine:1999:CLT

[GMdB99] Evarist Giné, Carlos Matrán, and Eustasio del Barrio. Central Limit Theorems for the Wasserstein distance between the empirical and the true distributions. Annals of Probability, 27(2): 1009–1071, April 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022677394. See correction [GMdB03].

Gine:2003:CCLn

[GMdB03] Evarist Giné, Carlos Matrán, and Eustasio del Barrio. Correction: Central limit theorems for the Wasserstein distance between the empirical and the true distributions. Annals of Probability, 31 (2):1142–1143, April 2003. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1048516548. See [GMdB99].

Gyongy:1999:SBE

[GN99] István Gyöngy and David Nualart. On the stochastic Burgers’ equation in the real line. Annals of Probability, 27(2):782–802, April 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022677386.

Gnedin:1994:SGG

[Gne94] Alexander V. Gnedin. A solution to the game of googol. Annals of Probability, 22(3):1588–1595, July 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988613.

Gnedin:1997:RCS

[Gne97] Alexander V. Gnedin. The representation of composition structures. Annals of Probability, 25(3):1437–1450, July 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1024404519.

Goggin:1994:CDC

[Gog94] Eimear M. Goggin. Convergence in distribution of conditional expectations. Annals of Probability, 22(2):1097–1114, April 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988743.

Goldman:1998:CDG

[Gol98] André Goldman. Sur une conjecture de D. G. Kendall concernant la cellule de Crofton du plan et sur sa contrepartie brownienne. (French) [On a conjecture of D. G. Kendall concerning the Crofton cell of the plan and its Brownian counterpart]. Annals of Probability, 26(4):1727–1750, October 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022855880.

Gotze:1991:RCM

[Got91] F. Gotze. On the rate of convergence in the multivariate CLT. Annals of Probability, 19(2):724–739, April 1991. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990448.

Gouet:1993:MFC

[Gou93] Raul Gouet. Martingale functional Central Limit Theorems for a generalized Pólya urn. Annals of Probability, 21(3):1624–1639, July 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989134.

Grubel:1992:FAS

[GP92] Rudolf Grubel and Susan M. Pitts. A functional approach to the stationary waiting time and idle period distributions of the GI/G/1 queue. Annals of Probability, 20(4):1754–1778, October 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989528.

Grama:1997:MDM

[Gra97] I. G. Grama. On moderate deviations for martingales. Annals of Probability, 25(1):152–183, January 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1024404283.

Griffeath:1993:FSP

[Gri93] David Griffeath. Frank Spitzer’s pioneering work on interacting particle systems. Annals of Probability, 21(2):608–621, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989258.

Grimmett:1995:SRC

[Gri95] Geoffrey Grimmett. The stochastic random-cluster process and the uniqueness of random-cluster measures. Annals of Probability, 23(4):1461–1510, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176987791.

Grigorescu:1999:SDB

[Gri99a] Ilie Grigorescu. Self-diffusion for Brownian motions with local interaction. Annals of Probability, 27(3):1208–1267, July 1999. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1022677445.

Grigorescu:1999:UTP

[Gri99b] Ilie Grigorescu. Uniqueness of the tagged particle process in a system with local interactions. Annals of Probability, 27(3): 1268–1282, July 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022677446.

Grimmett:1998:CPS

[GS98] G. R. Grimmett and A. M. Stacey. Critical probabilities for site and bond percolation models. Annals of Probability, 26(4): 1788–1812, October 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022855883.

Giraitis:1998:CLT

[GT98] Liudas Giraitis and Murad S. Taqqu. Central limit theorems for quadratic forms with time-domain conditions. Annals of Probability, 26(1):377–398, January 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022855425.

Gotze:1999:ADQ

[GT99] F. Götze and A. N. Tikhomirov. Asymptotic distribution of quadratic forms. Annals of Probability, 27(2):1072–1098, April 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022677395.

Gorostiza:1991:PCC

[GW91] Luis G. Gorostiza and Anton Wakolbinger. Persistence criteria for a class of critical branching particle systems in continuous time. Annals of Probability, 19(1):266–288, January 1991. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990544.

Gine:1990:BGE

[GZ90] Evarist Gine and Joel Zinn. Bootstrapping general empirical measures. Annals of Probability, 18(2):851–869, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990862.

Gine:1991:GCU

[GZ91] Evarist Gine and Joel Zinn. Gaussian characterization of uniform Donsker classes of functions. Annals of Probability, 19(2):758–782, April 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990450.

Gine:1994:RCD

[GZ94] Evarist Gine and Joel Zinn. A remark on convergence in distribution of U-statistics. Annals of Probability, 22(1):117–125, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988850.

Haeusler:1993:NLI

[Hae93] Erich Haeusler. A nonstandard law of the iterated logarithm for trimmed sums. Annals of Probability, 21(2):831–860, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989270.

Haggstrom:1997:ICD

[Häg97] Olle Häggström. Infinite clusters in dependent automorphism invariant percolation on trees. Annals of Probability, 25(3): 1423–1436, July 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1024404518.

Hahn:1994:RML

[Hah94] Marjorie G. Hahn. Review: Michel Ledoux, Michel Talagrand, Probability in Banach Spaces: Isoperimetry and Processes. Annals of Probability, 22(2):1115–1120, April 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988744.

Hall:1990:APB

[Hal90a] Peter Hall. Asymptotic properties of the bootstrap for heavytailed distributions. Annals of Probability, 18(3):1342–1360, July 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990748.

Hall:1990:OCR

[Hal90b] Peter Hall. Optimal convergence rates in signal recovery. Annals of Probability, 18(2):887–900, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990865.

Hall:1991:ILL

[Hal91] Peter Hall. On iterated logarithm laws for linear arrays and nonparametric regression estimators. Annals of Probability, 19 (2):740–757, April 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990449.

Hammack:1995:SID

[Ham95] William Hammack. Sharp inequalities for the distribution of a stochastic integral in which the integrator is a bounded submartingale. Annals of Probability, 23(1):223–235, January 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988384.

Hamana:1997:FRM

[Ham97a] Yuji Hamana. The fluctuation result for the multiple point range of two-dimensional recurrent random walks. Annals of Probability, 25(2):598–639, April 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1024404413.

Hambly:1997:BMR

[Ham97b] B. M. Hambly. Brownian motion on a random recursive Sierpinski gasket. Annals of Probability, 25(3):1059–1102, July 1997. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1024404506.

Hansen:1989:FCL

[Han89] Jennie C. Hansen. A functional Central Limit Theorem for random mappings. Annals of Probability, 17(1):317–332, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.

aop/1176991511. See correction [Han91].

Hansen:1991:CFC

[Han91] Jennie C. Hansen. Correction: A functional Central Limit Theorem for random mappings. Annals of Probability, 19(3):1393–1396, July 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990350. See [Han89].

Handjani:1999:CCT

[Han99] Shirin J. Handjani. The complete convergence theorem for coexistent threshold voter models. Annals of Probability, 27(1): 226–245, January 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022677260.

Harris:1993:RHK

[Har93] T. E. Harris. Review: H. Kunita, Stochastic Flows and Stochastic Differential Equations. Annals of Probability, 21(1):581–587, January 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989416.

Hara:1996:WFA

[Har96] Keisuke Hara. Wiener functionals associated with joint distributions of exit time and position from small geodesic balls. Annals of Probability, 24(2):825–837, April 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1039639363.

Harge:1999:PCC

[Har99] Gilles Hargé. A particular case of correlation inequality for the Gaussian measure. Annals of Probability, 27(4):1939–1951, October 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022874822.

He:1993:CLP

[He93] Yuanjiang He. Central limit properties of GZH-semigroups and their applications in probability theory. Annals of Probability, 21 (1):185–201, January 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989400.

Hennion:1997:LTP

[Hen97] H. Hennion. Limit theorems for products of positive random matrices. Annals of Probability, 25(4):1545–1587, October 1997. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1023481103.

Ho:1997:LTF

[HH97] Hwai-Chung Ho and Tailen Hsing. Limit theorems for functionals of moving averages. Annals of Probability, 25(4):1636–1669, October 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1023481106.

Hijab:1990:POC

[Hij90] Omar Hijab. Partially observed control of Markov processes. III. Annals of Probability, 18(3):1099–1125, July 1990. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990737.

Hildebrand:1992:NCS

[Hil92] Adolf Hildebrand. A note on the convergence of sums of independent random variables. Annals of Probability, 20(3):1204–1212, July 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989687.

Hildebrand:1993:RPF

[Hil93] Martin Hildebrand. Random processes of the form X_{n+1} = a_{n}X_{n} + b_{n} (mod p). Annals of Probability, 21(2):710–720, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL http://www.jstor.org/stable/2244672;
http://projecteuclid.org/euclid.aop/1176989264.

Hitczenko:1990:BCM

[Hit90] Pawel Hitczenko. Best constants in martingale version of Rosenthal’s inequality. Annals of Probability, 18(4):1656–1668, October 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990639.

Hitczenko:1994:DSR

[Hit94] Pawel Hitczenko. On a domination of sums of random variables by sums of conditionally independent ones. Annals of Probability, 22 (1):453–468, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988868.

Holst:1990:PAU

[HJ90] Lars Holst and Svante Janson. Poisson approximation using the Stein–Chen method and coupling: Number of exceedances of Gaussian random variables. Annals of Probability, 18(2):713–723, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990854.

Hill:1991:MOS

[HK91] Theodore P. Hill and Ulrich Krengel. Minimax-optimal stop rules and distributions in secretary problems. Annals of Probability, 19 (1):342–353, January 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990548.

Hahn:1995:ULP

[HK95] Marjorie G. Hahn and Michael J. Klass. Uniform local probability approximations: Improvements on Berry–Esseen. Annals of Probability, 23(1):446–463, January 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988394.

Hahn:1997:APS

[HK97] Marjorie G. Hahn and Michael J. Klass. Approximation of partial sums of arbitrary i.i.d. random variables and the precision of the usual exponential upper bound. Annals of Probability, 25 (3):1451–1470, July 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1024404520.

Hahn:1990:AJD

[HKW90] Marjorie G. Hahn, Jim Kuelbs, and Daniel C. Weiner. The asymptotic joint distribution of self-normalized censored sums and sums of squares. Annals of Probability, 18(3):1284–1341, July 1990. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990747.

Houdre:1996:SCS

[HL96] C. Houdré and M. T. Lacey. Spectral criteria, SLLN’s and a.s. convergence of series of stationary variables. Annals of Probability, 24(2):838–856, April 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1039639364.

Hambly:1998:SAB

[HL98a] B. M. Hambly and T. J. Lyons. Stochastic area for Brownian motion on the Sierpinski gasket. Annals of Probability, 26(1): 132–148, January 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022855414.

Hsing:1998:ERM

[HL98b] Tailen Hsing and M. R. Leadbetter. On the excursion random measure of stationary processes. Annals of Probability, 26(2):710– 742, April 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022855648.

Haggstrom:1995:ASS

[HM95] Olle Haggstrom and Ronald Meester. Asymptotic shapes for stationary first passage percolation. Annals of Probability, 23(4):

1511–1522, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176987792.

Haggstrom:1998:NNW

[HM98] Olle Häggström and Elchanan Mossel. Nearest-neighbor walks with low predictability profile and percolation in 2+ϵ dimensions. Annals of Probability, 26(3):1212–1231, July 1998. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1022855750.

Hoffman:1998:UNN

[Hof98] Christopher Hoffman. Unpredictable nearest neighbor processes. Annals of Probability, 26(4):1781–1787, October 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1022855882.

Hoglund:1990:AEP

[Hog90] Thomas Hoglund. An asymptotic expression for the probability of ruin within finite time. Annals of Probability, 18(1):378–389, January 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990954.

Hoglund:1991:RPF

[Hog91] Thomas Hoglund. The ruin problem for finite Markov chains. Annals of Probability, 19(3):1298–1310, July 1991. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990345.

Horvath:1990:NRP

[Hor90] Lajos Horváth. A note on the rate of Poisson approximation of empirical processes. Annals of Probability, 18(2):724–726, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990855.

Houdre:1991:LPM

[Hou91] Christian Houdre. On the linear prediction of multivariate (2, p)-bounded processes. Annals of Probability, 19(2):843–867, April 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990454.

Houdre:1992:SSP

[Hou92] Christian Houdre. On the spectral SLLN and pointwise ergodic theorem in L^{α}. Annals of Probability, 20(4):1731–1753, October 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989527.

Houdre:1995:ASC

[Hou95a] Christian Houdre. On the almost sure convergence of series of stationary and related nonstationary variables. Annals of Probability, 23(3):1204–1218, July 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988180.

Houdre:1995:RBL

[Hou95b] Christian Houdre. Reconstruction of band limited processes from irregular samples. Annals of Probability, 23(2):674–696, April 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988284.

Harel:1990:WCS

[HP90] Michel Harel and Madan L. Puri. Weak convergence of serial rank statistics under dependence with applications in time series and Markov processes. Annals of Probability, 18(3):1361–1387, July 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990749.

Haiman:1993:SIP

[HP93] George Haiman and Madan L. Puri. A strong invariance principle concerning the J-upper order statistics for stationary Gaussian sequences. Annals of Probability, 21(1):86–135, January 1993. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989395.

Houdre:1995:CII

[HPA95] Christian Houdre and Victor Perez-Abreu. Covariance identities and inequalities for functionals on Wiener and Poisson spaces. Annals of Probability, 23(1):400–419, January 1995. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988392.

Ho:1990:LDN

[HS90] Hwai-Chung Ho and Tze-Chien Sun. Limiting distributions of nonlinear vector functions of stationary Gaussian processes. Annals of Probability, 18(3):1159–1173, July 1990. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990740.

Horvath:1996:LDL

[HS96] Lajos Horváth and Qi-Man Shao. Large deviations and law of the iterated logarithm for partial sums normalized by the largest absolute observation. Annals of Probability, 24(3):1368–1387, July 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1065725185.

Hu:1998:LSS

[HS98] Yueyun Hu and Zhan Shi. The limits of Sinai’s simple random walk in random environment. Annals of Probability, 26(4): 1477–1521, October 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022855871.

Hebisch:1993:GEM

[HSC93] W. Hebisch and L. Saloff-Coste. Gaussian estimates for Markov chains and random walks on groups. Annals of Probability, 21 (2):673–709, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989263.

Hsing:1995:NAI

[Hsi95] Tailen Hsing. A note on the asymptotic independence of the sum and maximum of strongly mixing stationary random variables. Annals of Probability, 23(2):938–947, April 1995. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988296.

Hsing:1999:ADP

[Hsi99] Tailen Hsing. On the asymptotic distributions of partial sums of functionals of infinite-variance moving averages. Annals of Probability, 27(3):1579–1599, July 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022677460.

Hsu:1998:BRB

[Hsu98] Elton P. Hsu. Book review of “Probabilistic Techniques in Analysis” by Richard F. Bass. Annals of Probability, 26(3):1403–1405, July 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022855758.

Hu:1991:NRT

[Hu91] Inchi Hu. Nonlinear renewal theory for conditional random walks. Annals of Probability, 19(1):401–422, January 1991. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990553.

Hughes:1992:BED

[Hug92] H. R. Hughes. Brownian exit distributions from normal balls in S^{3} × H^{3}. Annals of Probability, 20(2):655–659, April 1992. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989797.

Husler:1990:EVH

[Hus90] J. Husler. Extreme values and high boundary crossings of locally stationary Gaussian processes. Annals of Probability, 18 (3):1141–1158, July 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990739.

Henze:1992:ASC

[HV92] N. Henze and B. Voigt. Almost sure convergence of certain slowly changing symmetric one- and multi-sample statistics. Annals of Probability, 20(2):1086–1098, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989819.

Harrison:1990:QMB

[HW90] J. M. Harrison and R. J. Williams. On the quasireversibility of a multiclass Brownian service station. Annals of Probability, 18 (3):1249–1268, July 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990745.

Hahn:1992:ABS

[HW92] Marjorie G. Hahn and Daniel C. Weiner. Asymptotic behavior of self-normalized trimmed sums: Nonnormal limits. Annals of Probability, 20(1):455–482, January 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989937.

Inglot:1992:SMD

[IKL92] Tadeusz Inglot, Wilbert C. M. Kallenberg, and Teresa Ledwina. Strong moderate deviation theorems. Annals of Probability, 20 (2):987–1003, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989814.

Inglot:1993:MLD

[IL93] Tadeusz Inglot and Teresa Ledwina. Moderately large deviations and expansions of large deviations for some functionals of weighted empirical processes. Annals of Probability, 21(3): 1691–1705, July 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989137.

Iscoe:1990:IDF

[IM90] I. Iscoe and D. McDonald. Induced Dirichlet forms and capacitary inequalities. Annals of Probability, 18(3):1195–1221, July 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990742.

Iscoe:1994:AETa

[IM94a] I. Iscoe and D. McDonald. Asymptotics of exit times for Markov jump processes I. Annals of Probability, 22(1):372–397, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988863.

Iscoe:1994:AETb

[IM94b] I. Iscoe and D. McDonald. Asymptotics of exit times for Markov jump processes. II: Applications to Jackson networks. Annals of Probability, 22(4):2168–2182, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988498.

Imkeller:1993:ECO

[Imk93] Peter Imkeller. Existence and continuity of occupation densities of stochastic integral processes. Annals of Probability, 21 (2):1050–1072, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989282.

Iscoe:1990:CVO

[IMM^{+}90] I. Iscoe, M. B. Marcus, D. McDonald, M. Talagrand, and J. Zinn. Continuity of l^{2}-valued Ornstein–Uhlenbeck processes. Annals of Probability, 18(1):68–84, January 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990938.

Imkeller:1994:IPW

[IN94] Peter Imkeller and David Nualart. Integration by parts on Wiener space and the existence of occupation densities. Annals of Probability, 22(1):469–493, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988869.

Ioffe:1991:SAV

[Iof91] Dimitry Ioffe. On some applicable versions of abstract large deviations theorems. Annals of Probability, 19(4):1629–1639, October 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990226.

Ioffe:1993:RPC

[Iof93] D. Ioffe. Recurrent perturbations of certain transient radially symmetric diffusions. Annals of Probability, 21(2):1124–1150, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989284.

Imkeller:1999:SAB

[IS99] Peter Imkeller and Michael Scheutzow. On the spatial asymptotic behavior of stochastic flows in Euclidean space. Annals of Probability, 27(1):109–129, January 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022677255.

Jacka:1991:OSB

[Jac91] S. D. Jacka. Optimal stopping and best constants for Doob-like inequalities I: The case p = 1. Annals of Probability, 19(4): 1798–1821, October 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990237.

Jacka:1993:LTO

[Jac93] S. D. Jacka. Local times, optimal stopping and semimartingales. Annals of Probability, 21(1):329–339, January 1993. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989407.

Jain:1990:LDL

[Jai90] Naresh C. Jain. Large deviation lower bounds for additive functionals of Markov processes. Annals of Probability, 18(3): 1071–1098, July 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990736.

Jakubowski:1993:AIR

[Jak93] Adam Jakubowski. An asymptotic independent representation in limit theorems for maxima of nonstationary random sequences. Annals of Probability, 21(2):819–830, April 1993. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989269.

Jakubowski:1996:CVT

[Jak96] Adam Jakubowski. Convergence in various topologies for stochastic integrals driven by semimartingales. Annals of Probability, 24 (4):2141–2153, October 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1041903222.

Janssen:1994:SIT

[Jan94] Arnold Janssen. Sums of independent triangular arrays and extreme order statistics. Annals of Probability, 22(4):1766–1793, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988482.

Janvresse:1998:FOC

[Jan98] Elise Janvresse. First order correction for the hydrodynamic limit of symmetric simple exclusion processes with speed change in dimension d ≥ 3. Annals of Probability, 26(4):1874–1912, October 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.

aop/1022855886. See note [Jan02].

Janvresse:2002:NAB

[Jan02] Elise Janvresse. Note about “First order correction for the hydrodynamic limit of symmetric simple exclusion processes with speed change in dimension d ≥ 3”. Annals of Probability, 30 (2):1000–1001, April 2002. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1023481015. See [Jan98].

Jerrum:1992:RGC

[Jer92] Mark Jerrum. Review: E. G. Coffman, George S. Lueker, Probabilistic Analysis of Packing and Partitioning Algorithms. Annals of Probability, 20(4):2164–2167, October 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989544.

Janvresse:1999:REC

[JLQY99] E. Janvresse, C. Landim, J. Quastel, and H. T. Yau. Relaxation to equilibrium of conservative dynamics. I: Zero-range processes. Annals of Probability, 27(1):325–360, January 1999. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1022677265.

Jacod:1998:AED

[JP98] Jean Jacod and Philip Protter. Asymptotic error distributions for the Euler method for stochastic differential equations. Annals of Probability, 26(1):267–307, January 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022855419.

Jurek:1992:OEP

[Jur92] Zbigniew J. Jurek. Operator exponents of probability measures and Lie semigroups. Annals of Probability, 20(2):1053–1062, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989817.

Jing:1999:ENB

[JW99] Bing-Yi Jing and Qiying Wang. An exponential nonuniform Berry–Esseen bound for self-normalized sums. Annals of Probability, 27(4):2068–2088, October 1999. CODEN APBYAE.

ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022874829.

Kallenberg:1991:ICM

[Kal91] Olav Kallenberg. On an independence criterion for multiple Wiener integrals. Annals of Probability, 19(2):483–485, April 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990436.

Kalpazidou:1994:RRT

[Kal94] S. Kalpazidou. Rotational representations of transition matrix functions. Annals of Probability, 22(2):703–712, April 1994. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988726.

Kalpazidou:1995:RDS

[Kal95] S. Kalpazidou. On the rotational dimension of stochastic matrices. Annals of Probability, 23(2):966–975, April 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988298.

Kallenberg:1996:EUF

[Kal96] Olav Kallenberg. On the existence of universal functional solutions to classical SDE’s. Annals of Probability, 24(1):196–205, January 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1042644713.

Kallenberg:1999:BTS

[Kal99a] Olav Kallenberg. Ballot theorems and sojourn laws for stationary processes. Annals of Probability, 27(4):2011–2019, October 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022874825.

Kallenberg:1999:PMD

[Kal99b] Olav Kallenberg. Palm measure duality and conditioning in regenerative sets. Annals of Probability, 27(2):945–969, April 1999. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1022677391.

Katzenberger:1991:SSD

[Kat91] G. S. Katzenberger. Solutions of a stochastic differential equation forced onto a manifold by a large drift. Annals of Probability, 19 (4):1587–1628, October 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990225.

Konig:1997:CLT

[KdHvdH97] W. König, F. den Hollander, and R. van der Hofstad. Central limit theorem for the Edwards model. Annals of Probability, 25 (2):573–597, April 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1024404412.

Keener:1992:LTR

[Kee92] Robert W. Keener. Limit theorems for random walks conditioned to stay positive. Annals of Probability, 20(2):801–824, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989807.

Kendall:1995:RPM

[Ken95] Wilfrid S. Kendall. The radial part of a Γ-martingale and a non-implosion theorem. Annals of Probability, 23(2):479–500, April 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988276.

Kersting:1996:HCD

[Ker96] G. Kersting. Harmonic coordinates for diffusions in the plane. Annals of Probability, 24(3):1239–1268, July 1996. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1065725180.

Kesten:1993:FSW

[Kes93] Harry Kesten. Frank Spitzer’s work on random walk and Brownian motion. Annals of Probability, 21(2):593–607, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989257.

Kesten:1997:UFC

[Kes97] Harry Kesten. A universal form of the Chung-type law of the iterated logarithm. Annals of Probability, 25(4):1588–1620, October 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1023481104.

Kouritzin:1994:LIL

[KH94] M. A. Kouritzin and A. J. Heunis. A law of the iterated logarithm for stochastic processes defined by differential equations with a small parameter. Annals of Probability, 22(2):659–679, April 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988724.

Khoshnevisan:1992:MIF

[Kho92] Davar Khoshnevisan. Moment inequalities for functionals of the Brownian convex Hull. Annals of Probability, 20(2):627–630, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989794.

Khoshnevisan:1993:ECC

[Kho93] Davar Khoshnevisan. An embedding of compensated compound Poisson processes with applications to local times. Annals of Probability, 21(1):340–361, January 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989408.

Khoshnevisan:1994:ERC

[Kho94] Davar Khoshnevisan. Exact rates of convergence to Brownian local time. Annals of Probability, 22(3):1295–1330, July 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988604.

Khoshnevisan:1995:DBB

[Kho95] Davar Khoshnevisan. On the distribution of bubbles of the Brownian sheet. Annals of Probability, 23(2):786–805, April 1995. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988290.

Kifer:1990:DTV

[Kif90] Yuri Kifer. A discrete-time version of the Wentzell–Friedlin theory. Annals of Probability, 18(4):1676–1692, October 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990641.

Kennedy:1991:ABR

[KK91] Douglas P. Kennedy and Robert P. Kertz. The asymptotic behavior of the reward sequence in the optimal stopping of i.i.d. random variables. Annals of Probability, 19(1):329–341, January 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990547.

Khalfin:1994:SCT

[KK94] L. A. Khalfin and L. B. Klebanov. A solution of the computer tomography paradox and estimating the distances between the densities of measures with the same marginals. Annals of Probability, 22(4):2235–2241, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988502.

Kersting:1995:SCN

[KK95] G. Kersting and F. C. Klebaner. Sharp conditions for nonexplosions and explosions in Markov jump processes. Annals of Probability, 23(1):268–272, January 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988386.

Kesten:1992:NCM

[KL92] Harry Kesten and Gregory F. Lawler. A necessary condition for making money from fair games. Annals of Probability, 20 (2):855–882, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989809.

Klass:1990:ULB

[Kla90] Michael J. Klass. Uniform lower bounds for randomly stopped Banach space valued random sums. Annals of Probability, 18 (2):790–809, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990858.

Klein:1994:ECP

[Kle94] Abel Klein. Extinction of contact and percolation processes in a random environment. Annals of Probability, 22(3):1227–1251, July 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988601.

Klenke:1996:DCR

[Kle96] Achim Klenke. Different clustering regimes in systems of hierarchically interacting diffusions. Annals of Probability, 24(2):660–697, April 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1039639358.

Klenke:1997:MSA

[Kle97] Achim Klenke. Multiple scale analysis of clusters in spatial branching models. Annals of Probability, 25(4):1670–1711, October 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1023481107.

Klenke:1998:CIM

[Kle98] Achim Klenke. Clustering and invariant measures for spatial branching models with infinite variance. Annals of Probability, 26(3):1057–1087, July 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022855745.

Khoshnevisan:1996:PET

[KLS96] Davar Khoshnevisan, Thomas M. Lewis, and Zhan Shi. On a problem of Erdős and Taylor. Annals of Probability, 24(2):761–787, April 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1039639361.

Kuelbs:1994:LIR

[KLT94] James Kuelbs, Wenbo V. Li, and Michel Talagrand. Lim inf results for Gaussian samples and Chung’s functional LIL. Annals of Probability, 22(4):1879–1903, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988488.

Kesten:1992:RTS

[KM92] Harry Kesten and R. A. Maller. Ratios of trimmed sums and order statistics. Annals of Probability, 20(4):1805–1842, October 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989530.

Kesten:1994:ILI

[KM94a] Harry Kesten and R. A. Maller. Infinite limits and infinite limit points of random walks and trimmed sums. Annals of Probability, 22(3):1473–1513, July 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988609.

Knight:1994:CST

[KM94b] Frank B. Knight and Bernard Maisonneuve. A characterization of stopping times. Annals of Probability, 22(3):1600–1606, July 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988615.

Klass:1997:OMB

[KN97] Michael J. Klass and Krzysztof Nowicki. Order of magnitude bounds for expectations of Δ_{2}-functions of nonnegative random bilinear forms and generalized U-statistics. Annals of Probability, 25(3):1471–1501, July 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1024404521.

Komorowski:1996:DAA

[Kom96] Tomasz Komorowski. Diffusion approximation for the advection of particles in a strongly turbulent random environment. Annals of Probability, 24(1):346–376, January 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1042644720.

Konig:1996:CLT

[Kön96] Wolfgang König. A central limit theorem for a one-dimensional polymer measure. Annals of Probability, 24(2):1012–1035, April 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1039639376.

Kurtz:1991:WLT

[KP91] Thomas G. Kurtz and Philip Protter. Weak limit theorems for stochastic integrals and stochastic differential equations. Annals of Probability, 19(3):1035–1070, July 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990334.

Kinoshita:1991:CSR

[KR91] K. Kinoshita and Sidney I. Resnick. Convergence of scaled random samples in R^{d}. Annals of Probability, 19(4):1640–1663, October 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990227.

Kristiansen:1994:PSC

[Kri94] Gundorph K. Kristiansen. A proof of Steutel’s conjecture. Annals of Probability, 22(1):442–452, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988867.

Krone:1993:LTS

[Kro93] Stephen M. Krone. Local times for superdiffusions. Annals of Probability, 21(3):1599–1623, July 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989133.

Krylov:1997:SS

[Kry97] N. V. Krylov. On SPDE’s and superdiffusions. Annals of Probability, 25(4):1789–1809, October 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1023481111.

Kwapien:1991:HMM

[KS91] Stanislaw Kwapien and Jerzy Szulga. Hypercontraction methods in moment inequalities for series of independent random variables in normed spaces. Annals of Probability, 19(1):369–379, January 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990550.

Klein:1996:QDV

[KS96] Irene Klein and Walter Schachermayer. A quantitative and a dual version of the Halmos–Savage theorem with applications to mathematical finance. Annals of Probability, 24(2):867–881, April 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1039639366.

Khoshnevisan:1999:BSC

[KS99] Davar Khoshnevisan and Zhan Shi. Brownian sheet and capacity. Annals of Probability, 27(3):1135–1159, July 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1022677442.

Kurtz:1991:RTC

[Kur91] Thomas G. Kurtz. Random time changes and convergence in distribution under the Meyer–Zheng conditions. Annals of Probability, 19(3):1010–1034, July 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990333.

Kokic:1990:EES

[KW90] P. N. Kokic and N. C. Weber. An Edgeworth expansion for Ustatistics based on samples from finite populations. Annals of Probability, 18(1):390–404, January 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990955.

Kalikow:1992:ECS

[KW92] Steven Kalikow and Benjamin Weiss. Explicit codes for some infinite entropy Bernoulli shifts. Annals of Probability, 20(1):397–402, January 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989933.

Kallianpur:1996:LDC

[KX96] Gopinath Kallianpur and Jie Xiong. Large deviations for a class of stochastic partial differential equations. Annals of Probability, 24 (1):320–345, January 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1042644719.

Kesten:1990:PLF

[KZ90] Harry Kesten and Yu Zhang. The probability of a large finite cluster in supercritical Bernoulli percolation. Annals of Probability, 18(2):537–555, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990844.

Klass:1994:ASM

[KZ94] Michael J. Klass and Cun-Hui Zhang. On the almost sure minimal growth rate of partial sum maxima. Annals of Probability, 22(4): 1857–1878, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988487.

Lacey:1990:LDM

[Lac90] Michael Lacey. Large deviations for the maximum local time of stable Lévy processes. Annals of Probability, 18(4):1669–1675, October 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990640.

Lachal:1997:LAC

[Lac97] Aimé Lachal. Local asymptotic classes for the successive primitives of Brownian motion. Annals of Probability, 25(4):1712–1734, October 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1023481108.

Lahiri:1993:RAE

[Lah93] Soumendra Nath Lahiri. Refinements in asymptotic expansions for sums of weakly dependent random vectors. Annals of Probability, 21(2):791–799, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989267.

Lalley:1992:BME

[Lal92] Steven P. Lalley. Brownian motion and the equilibrium measure on the Julia set of a rational mapping. Annals of Probability, 20(4): 1932–1967, October 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989536.

Lalley:1993:FRR

[Lal93] Steven P. Lalley. Finite range random walk on free groups and homogeneous trees. Annals of Probability, 21(4):2087–2130, October 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989012.

Lalley:1996:CSR

[Lal96] Steven P. Lalley. Cycle structure of riffle shuffles. Annals of Probability, 24(1):49–73, January 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1042644707.

Lalley:1999:GPI

[Lal99] Steven P. Lalley. Growth profile and invariant measures for the weakly supercritical contact process on a homogeneous tree. Annals of Probability, 27(1):206–225, January 1999. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1022677259. See cor-

rection [Lal02].

Lalley:2002:CGP

[Lal02] Steven P. Lalley. Correction: Growth profile and invariant measures for the weakly supercritical contact process on a homogeneous tree. Annals of Probability, 30(4):2108–2112, October 2002. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1039548383. See [Lal99].

Landim:1991:HEA

[Lan91] C. Landim. Hydrodynamical equation for attractive particle systems on Z^{d}. Annals of Probability, 19(4):1537–1558, October 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990222.

Landim:1992:OTL

[Lan92] C. Landim. Occupation time large deviations for the symmetric simple exclusion process. Annals of Probability, 20(1):206–231, January 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989925.

Landim:1993:CLE

[Lan93] C. Landim. Conservation of local equilibrium for attractive particle systems on Z^{d}. Annals of Probability, 21(4):1782–1808, October 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989000.

Landim:1996:HLS

[Lan96] C. Landim. Hydrodynamical limit for space inhomogeneous onedimensional totally asymmetric zero-range processes. Annals of Probability, 24(2):599–638, April 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1039639356.

Latala:1997:EMS

[Lat97] Rafał Latała. Estimation of moments of sums of independent real random variables. Annals of Probability, 25(3):1502–1513, July 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1024404522.

Lawler:1995:SFI

[Law95] Gregory F. Lawler. Subdiffusive fluctuations for internal diffusion limited aggregation. Annals of Probability, 23(1):71–86, January 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988377.

Lawler:1992:IDL

[LBG92] Gregory F. Lawler, Maury Bramson, and David Griffeath. Internal diffusion limited aggregation. Annals of Probability, 20(4): 2117–2140, October 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989542.

Lai:1997:MRS

[LdlP97] Tze Leung Lai and Victor H. de la Peña. Moments of randomly stopped U-statistics. Annals of Probability, 25(4):2055–2081, October 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1023481120.

LeGall:1991:BET

[Le 91] Jean-François Le Gall. Brownian excursions, trees and measurevalued branching processes. Annals of Probability, 19(4):1399– 1439, October 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990218.

Ledrappier:1995:CLT

[Led95] François Ledrappier. Central Limit Theorem in negative curvature. Annals of Probability, 23(3):1219–1233, July 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988181.

Lee:1990:CLT

[Lee90] Tzong-Yow Lee. Conditioned limit theorems of stopped critical branching Bessel processes. Annals of Probability, 18(1):272–289, January 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990949.

Lee:1991:CLD

[Lee91] Tzong-Yow Lee. Conditional limit distributions of critical branching Brownian motions. Annals of Probability, 19(1):289–311, January 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990545.

Lee:1992:SLD

[Lee92] Tzong-Yow Lee. Some large-deviation theorems for branching diffusions. Annals of Probability, 20(3):1288–1309, July 1992. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989692.

Lee:1993:SLT

[Lee93] Tzong-Yow Lee. Some limit theorems for super-Brownian motion and semilinear differential equations. Annals of Probability, 21 (2):979–995, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989278.

Lee:1995:LDI

[Lee95] Tzong-Yow Lee. Large deviations for independent random walks on the line. Annals of Probability, 23(3):1315–1331, July 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988186.

Leuridan:1995:CNM

[Leu95] Christophe Leuridan. Contrôle de la norme H^{p} d’une martingale par des maximums de temps locaux. (French) [Control of the H^{p} norm of a martingale by maxima of local times]. Annals of Probability, 23(3):1289–1299, July 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988184.

Leuridan:1997:PFP

[Leu97] Christophe Leuridan. Le point d’un fermé le plus visité par le mouvement brownien. (French) [The closed point most visited by Brownian motion]. Annals of Probability, 25(2):953–996, April 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1024404426.

Liao:1992:SGM

[Lia92] Ming Liao. Symmetry groups of Markov processes. Annals of Probability, 20(2):563–578, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989791.

Liao:1997:LES

[Lia97] Ming Liao. Liapounov exponents of stochastic flows. Annals of Probability, 25(3):1241–1256, July 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1024404512.

Lifshits:1997:LTP

[Lif97] M. A. Lifshits. On the lower tail probabilities of some random series. Annals of Probability, 25(1):424–442, January 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1024404294.

Liggett:1991:RCA

[Lig91] Thomas M. Liggett. L_{2} rates of convergence for attractive reversible nearest particle systems: The critical case. Annals of Probability, 19(3):935–959, July 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990330.

Liggett:1992:SOD

[Lig92] Thomas M. Liggett. The survival of one-dimensional contact processes in random environments. Annals of Probability, 20(2):696– 723, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989801.

Liggett:1994:CTV

[Lig94] Thomas M. Liggett. Coexistence in threshold voter models. Annals of Probability, 22(2):764–802, April 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:/

/projecteuclid.org/euclid.aop/1176988729. See correction

[Lig99].

Liggett:1995:IUB

[Lig95] Thomas M. Liggett. Improved upper bounds for the contact process critical value. Annals of Probability, 23(2):697–723, April 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988285.

Liggett:1996:MTP

[Lig96] Thomas M. Liggett. Multiple transition points for the contact process on the binary tree. Annals of Probability, 24(4):1675–1710, October 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1041903202.

Liggett:1997:SMI

[Lig97] Thomas M. Liggett. Stochastic models of interacting systems. Annals of Probability, 25(1):1–29, January 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1024404276.

Liggett:1999:CBT

[Lig99] Thomas M. Liggett. Correction: “Coexistence in threshold voter models” [Ann. Probab. 22 (1994), no. 2, 764–802; MR 95m:60164]. Annals of Probability, 27(4):2120–2121, October 1999. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1022874831. See

[Lig94].

Linde:1991:GML

[Lin91a] Werner Linde. Gaussian measure of large balls in l_{p}. Annals of Probability, 19(3):1264–1279, July 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990343.

Lindvall:1991:WD

[Lin91b] Torgny Lindvall. W. Doeblin, 1915–1940. Annals of Probability, 19(3):929–934, July 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990329.

Liu:1995:STP

[Liu95] Xijian Liu. Symmetric two-particle exclusion-eating processes. Annals of Probability, 23(3):1439–1455, July 1995. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988191.

LeGall:1998:BPLb

[LL98a] Jean-François Le Gall and Yves Le Jan. Branching processes in Lévy processes: Laplace functionals of snakes and superprocesses. Annals of Probability, 26(4):1407–1432, October 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1022855868.

LeGall:1998:BPLa

[LL98b] Jean-François Le Gall and Yves Le Jan. Branching processes in Lévy processes: the exploration process. Annals of Probability, 26 (1):213–252, January 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022855417.

Li:1999:AME

[LL99] Wenbo V. Li and Werner Linde. Approximation, metric entropy and small ball estimates for Gaussian measures. Annals of Probability, 27(3):1556–1578, July 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022677459.

Licea:1996:GTD

[LN96] Cristina Licea and Charles M. Newman. Geodesics in twodimensional first-passage percolation. Annals of Probability, 24 (1):399–410, January 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1042644722.

Leon:1998:SEE

[LN98] Jorge A. León and David Nualart. Stochastic evolution equations with random generators. Annals of Probability, 26(1):149–186, January 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022855415.

Li:1998:CSV

[LO98] Shoumei Li and Yukio Ogura. Convergence of set valued suband supermartingales in the Kuratowski–Mosco sense. Annals of Probability, 26(3):1384–1402, July 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022855757.

Latala:1999:GMD

[LO99] Rafał Latała and Krzysztof Oleszkiewicz. Gaussian measures of dilatations of convex symmetric sets. Annals of Probability, 27(4): 1922–1938, October 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022874821.

Lotov:1996:SBC

[Lot96] V. I. Lotov. On some boundary crossing problems for Gaussian random walks. Annals of Probability, 24(4):2154–2171, October 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1041903223.

Landim:1996:SPD

[LOY96] C. Landim, S. Olla, and H. T. Yau. Some properties of the diffusion coefficient for asymmetric simple exclusion processes. Annals of Probability, 24(4):1779–1808, October 1996. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1041903206.

Lyons:1992:RWR

[LP92] Russell Lyons and Robin Pemantle. Random walk in a random environment and first-passage percolation on trees. Annals of Probability, 20(1):125–136, January 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:/

/projecteuclid.org/euclid.aop/1176989920. See correction

[LP03].

LeGall:1995:HMS

[LP95] Jean-François Le Gall and Edwin A. Perkins. The Hausdorff measure of the support of two-dimensional super-Brownian motion.

Annals of Probability, 23(4):1719–1747, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176987800.

Lyons:2003:CRW

[LP03] Russell Lyons and Robin Pemantle. Correction: Random walk in a random environment and first-passage percolation on trees. Annals of Probability, 31(1):528–529, January 2003. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1046294319. See

[LP92].

Lyons:1995:CPL

[LPP95] Russell Lyons, Robin Pemantle, and Yuval Peres. Conceptual proofs of L log L criteria for mean behavior of branching processes. Annals of Probability, 23(3):1125–1138, July 1995. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988176.

Lyons:1996:RWL

[LPP96] Russell Lyons, Robin Pemantle, and Yuval Peres. Random walks on the lamplighter group. Annals of Probability, 24(4):1993–2006, October 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1041903214.

LeGall:1991:RSR

[LR91] Jean-François Le Gall and Jay Rosen. The range of stable random walks. Annals of Probability, 19(2):650–705, April 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990446.

Lee:1995:LDT

[LR95] Tzong-Yow Lee and Bruno Remillard. Large deviations for the three-dimensional super-Brownian motion. Annals of Probability, 23(4):1755–1771, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176987802.

Lee:1990:ASR

[LRS90] Mei-Ling Ting Lee, Svetlozar T. Rachev, and Gennady Samorodnitsky. Association of stable random variables. Annals of Probability, 18(4):1759–1764, October 1990. CODEN APBYAE.

ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990646.

Li:1992:LIL

[LRW92] Deli Li, M. Bhaskara Rao, and Xiangchen Wang. The law of the iterated logarithm for independent random variables with multidimensional indices. Annals of Probability, 20(2):660–674, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989798.

Lalley:1992:LTF

[LS92] S. Lalley and T. Sellke. Limit theorems for the frontier of a one-dimensional branching diffusion. Annals of Probability, 20 (3):1310–1340, July 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989693.

Lloyd:1997:QOC

[LS97] Peter Lloyd and Aidan Sudbury. Quantum operators in classical probability theory. IV. quasi-duality and thinnings of interacting particle systems. Annals of Probability, 25(1):96–114, January 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1024404280.

Lalley:1998:LSW

[LS98] Steven P. Lalley and Tom Sellke. Limit set of a weakly supercritical contact process on a homogeneous tree. Annals of Probability, 26(2):644–657, April 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022855646.

Lyons:1999:IPC

[LS99a] Russell Lyons and Oded Schramm. Indistinguishability of percolation clusters. Annals of Probability, 27(4):1809–1836, October 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022874816.

Lyons:1999:LSI

[LS99b] Terry Lyons and Lucretiu Stoica. The limits of stochastic integrals of differential forms. Annals of Probability, 27(1):1–49, January 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022677253.

Liggett:1997:DPM

[LSS97] T. M. Liggett, R. H. Schonmann, and A. M. Stacey. Domination by product measures. Annals of Probability, 25(1):71–95, January 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1024404279.

Landim:1996:SGZ

[LSV96] C. Landim, S. Sethuraman, and S. Varadhan. Spectral gap for zero-range dynamics. Annals of Probability, 24(4):1871–1902, October 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1041903209.

Ledoux:1990:SAI

[LT90] M. Ledoux and M. Talagrand. Some applications of isoperimetric methods to strong limit theorems for sums of independent random variables. Annals of Probability, 18(2):754–789, April 1990. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990857.

Lee:1998:WPL

[LT98] Tzong-Yow Lee and Fred Torcaso. Wave propagation in a lattice KPP equation in random media. Annals of Probability, 26 (3):1179–1197, July 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022855748.

Lu:1994:EFO

[Lu94] Shenglin Lu. Equilibrium fluctuations of a one-dimensional nongradient Ginzburg–Landau model. Annals of Probability, 22(3): 1252–1272, July 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988602.

Lefevre:1995:PAF

[LU95a] Claude Lefevre and Sergey Utev. Poisson approximation for the final state of a generalized epidemic process. Annals of Probability, 23(3):1139–1162, July 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988177.

Lu:1995:HSL

[Lu95b] Shenglin Lu. Hydrodynamic scaling limits with deterministic initial configurations. Annals of Probability, 23(4):1831–1852, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176987805.

Lai:1994:AED

[LW94] Tze Leung Lai and Julia Qizhi Wang. Asymptotic expansions for the distributions of stopped random walks and first passage times. Annals of Probability, 22(4):1957–1992, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988491.

Lawler:1999:IEP

[LW99] Gregory F. Lawler and Wendelin Werner. Intersection exponents for planar Brownian motion. Annals of Probability, 27(4): 1601–1642, October 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022874810.

Lee:1998:LSI

[LY98] Tzong-Yow Lee and Horng-Tzer Yau. Logarithmic Sobolev inequality for some models of random walks. Annals of Probability, 26(4):1855–1873, October 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022855885.

Lyons:1990:RWP

[Lyo90] Russell Lyons. Random walks and percolation on trees. Annals of Probability, 18(3):931–958, July 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990730.

Lyons:1992:RWC

[Lyo92] Russell Lyons. Random walks, capacity and percolation on trees. Annals of Probability, 20(4):2043–2088, October 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989540.

Lyons:1994:DDP

[LZ94] T. J. Lyons and T. S. Zhang. Decomposition of Dirichlet processes and its application. Annals of Probability, 22(1):494–524, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988870.

Liao:1995:RPB

[LZ95] M. Liao and W. A. Zheng. Radial part of Brownian motion on a Riemannian manifold. Annals of Probability, 23(1):173–177, January 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988382.

Lyons:1999:CEM

[LZ99] Terry Lyons and Ofer Zeitouni. Conditional exponential moments for iterated Wiener integrals. Annals of Probability, 27(4): 1738–1749, October 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022874813.

Major:1990:NAU

[Maj90] Peter Major. A note on the approximation of the uniform empirical process. Annals of Probability, 18(1):129–139, January 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990941.

Major:1994:ADW

[Maj94] Peter Major. Asymptotic distributions for weighted U-statistics. Annals of Probability, 22(3):1514–1535, July 1994. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988610.

Major:1999:LBE

[Maj99] Péter Major. The limit behavior of elementary symmetric polynomials of i.i.d. random variables when their order tends to infinity. Annals of Probability, 27(4):1980–2010, October 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1022874824.

Marton:1996:BDI

[Mar96] K. Marton. Bounding d̄-distance by informational divergence: a method to prove measure concentration. Annals of Probability, 24(2):857–866, April 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1039639365.

Marsalle:1999:SPF

[Mar99] Laurence Marsalle. Slow points and fast points of local times. Annals of Probability, 27(1):150–165, January 1999. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1022677257.

Massart:1990:TCD

[Mas90] P. Massart. The tight constant in the Dvoretzky–Kiefer–Wolfowitz inequality. Annals of Probability, 18(3):1269–1283, July 1990. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990746.

Mason:1994:UOS

[Mas94] David M. Mason. A universal one-sided law of the iterated logarithm. Annals of Probability, 22(4):1826–1837, October 1994. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988485.

Matthews:1991:GRL

[Mat91] Peter Matthews. Generating a random linear extension of a partial order. Annals of Probability, 19(3):1367–1392, July 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990349.

Meerschaert:1990:MRV

[Mee90] Mark M. Meerschaert. Moments of random vectors which belong to some domain of normal attraction. Annals of Probability, 18 (2):870–876, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990863.

Meerschaert:1991:SDG

[Mee91] Mark M. Meerschaert. Spectral decomposition for generalized domains of attraction. Annals of Probability, 19(2):875–892, April 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990456.

Melfi:1992:NMR

[Mel92] Vincent F. Melfi. Nonlinear Markov renewal theory with statistical applications. Annals of Probability, 20(2):753–771, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989804.

Mikami:1990:AAI

[Mik90] Toshio Mikami. Asymptotic analysis of invariant density of randomly perturbed dynamical systems. Annals of Probability, 18 (2):524–536, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990843.

Mikami:1991:LDT

[Mik91] Toshio Mikami. Large deviations theorems for empirical measures in Freidlin–Wentzell exit problems. Annals of Probability, 19(1):58–82, January 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990536.

Merzbach:1990:MPP

[MN90] Ely Merzbach and David Nualart. Markov properties for point processes on the plane. Annals of Probability, 18(1):342–358, January 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990952.

Millet:1992:LDC

[MNS92] A. Millet, D. Nualart, and M. Sanz. Large deviations for a class of anticipating stochastic differential equations. Annals of Probability, 20(4):1902–1931, October 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989535.

Mogulskii:1993:LDP

[Mog93] A. A. Mogulskii. Large deviations for processes with independent increments. Annals of Probability, 21(1):202–215, January 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989401.

Mountford:1992:CVU

[Mou92] T. S. Mountford. The critical value for the uniform nearest particle process. Annals of Probability, 20(4):2031–2042, October 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989539.

Mountford:1991:RLP

[MP91] T. S. Mountford and S. C. Port. The range of a Lévy process. Annals of Probability, 19(1):221–225, January 1991. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990541.

Mountford:1996:SGR

[MP96] T. S. Mountford and S. C. Port. Small gaps in the range of stable processes. Annals of Probability, 24(1):438–452, January 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1042644725.

Maitra:1991:BMV

[MPS91] A. Maitra, R. Purves, and W. Sudderth. A Borel measurable version of Konig’s lemma for random paths. Annals of Probability, 19(1):423–451, January 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990554.

Marcus:1992:VLT

[MR92a] Michael B. Marcus and Jay Rosen. p-variation of the local times of symmetric stable processes and of Gaussian processes with stationary increments. Annals of Probability, 20(4):1685–1713, October 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989525.

Marcus:1992:SPP

[MR92b] Michael B. Marcus and Jay Rosen. Sample path properties of the local times of strongly symmetric Markov processes via Gaussian processes. Annals of Probability, 20(4):1603–1684, October 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989524.

Marcus:1994:LIL

[MR94] Michael B. Marcus and Jay Rosen. Laws of the iterated logarithm for the local times of symmetric Lévy processes and recurrent random walks. Annals of Probability, 22(2):626–658, April 1994. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988723.

Marcus:1996:GCS

[MR96a] Michael B. Marcus and Jay Rosen. Gaussian chaos and sample path properties of additive functionals of symmetric Markov processes. Annals of Probability, 24(3):1130–1177, July 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1065725177.

Marcus:1996:RFS

[MR96b] Michael B. Marcus and Jay Rosen. Random Fourier series and continuous additive functionals of Lévy processes on the torus. Annals of Probability, 24(3):1178–1218, July 1996. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1065725178.

Marcus:1999:AFS

[MR99] Michael B. Marcus and Jay Rosen. Additive functionals of several Lévy processes and intersection local times. Annals of Probability, 27(4):1643–1678, October 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022874811.

Mason:1992:NSC

[MS92] David M. Mason and Galen R. Shorack. Necessary and sufficient conditions for asymptotic normality of L-statistics. Annals of Probability, 20(4):1779–1804, October 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989529.

Maitra:1993:BSG

[MS93] A. Maitra and W. Sudderth. Borel stochastic games with lim sup payoff. Annals of Probability, 21(2):861–885, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989271.

Marton:1994:ECE

[MS94] Katalin Marton and Paul C. Shields. Entropy and the consistent estimation of joint distributions. Annals of Probability, 22 (2):960–977, April 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988736. See correction [MS96].

Monat:1995:FSD

[MS95] Pascale Monat and Christophe Stricker. Follmer–Schweizer decomposition and mean-variance hedging for general claims. Annals of Probability, 23(2):605–628, April 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988281.

Marton:1996:CBC

[MS96] Katalin Marton and Paul C. Shields. Correction: “Entropy and the consistent estimation of joint distributions”. Annals of Probability, 24(1):541–545, January 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1042644732. See [MS94].

Mohammed:1997:LEL

[MS97] Salah-Eldin A. Mohammed and Michael K. R. Scheutzow. Lyapunov exponents of linear stochastic functional-differential equations. II. examples and case studies. Annals of Probability, 25 (3):1210–1240, July 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1024404511.

Mohammed:1998:SES

[MS98a] Salah-Eldin A. Mohammed and Michael K. R. Scheutzow. Spatial estimates for stochastic flows in Euclidean space. Annals of Probability, 26(1):56–77, January 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022855411.

Mountford:1998:FAC

[MS98b] Thomas Mountford and Ted Sweet. Finite approximations to the critical reversible nearest particle system. Annals of Probability, 26 (4):1751–1780, October 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022855881.

Mohammed:1999:SMT

[MS99] Salah-Eldin A. Mohammed and Michael K. R. Scheutzow. The stable manifold theorem for stochastic differential equations. Annals of Probability, 27(2):615–652, April 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022677380.

Madras:1994:CBC

[MSS94] Neal Madras, Rinaldo Schinazi, and Roberto H. Schonmann. On the critical behavior of the contact process in deterministic inhomogeneous environments. Annals of Probability, 22(3):1140–1159, July 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988598.

Millet:1999:SWE

[MSS99] Annie Millet and Marta Sanz-Solé. A stochastic wave equation in two space dimension: Smoothness of the law. Annals of Probability, 27(2):803–844, April 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022677387.

Mandelbaum:1990:OSB

[MSV90] Avi Mandelbaum, Larry A. Shepp, and Robert J. Vanderbei. Optimal switching between a pair of Brownian motions. Annals of Probability, 18(3):1010–1033, July 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990734.

Mueller:1993:CIM

[Mue93] Carl Mueller. Coupling and invariant measures for the heat equation with noise. Annals of Probability, 21(4):2189–2199, October 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989016.

Mueller:1996:SIC

[Mue96] Carl Mueller. Singular initial conditions for the heat equation with a noise term. Annals of Probability, 24(1):377–398, January 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1042644721.

Mueller:1997:LTE

[Mue97] Carl Mueller. Long time existence for the wave equation with a noise term. Annals of Probability, 25(1):133–151, January 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1024404282.

Mukherjea:1994:TPD

[Muk94] Arunava Mukherjea. Tightness of products of i.i.d. random matrices. II. Annals of Probability, 22(4):2223–2234, October 1994. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988501.

Mauldin:1991:SOP

[MvW91] R. Daniel Mauldin and H. v. Weizsacker. Some orthogonality preserving kernels which are not completely orthogonal. Annals of Probability, 19(1):396–400, January 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990552.

Mayer-Wolf:1990:CRF

[MW90] Eduardo Mayer-Wolf. The Cramér–Rao functional and limiting laws. Annals of Probability, 18(2):840–850, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990861.

Mayer-Wolf:1993:PSG

[MWZ93] Eddy Mayer-Wolf and Ofer Zeitouni. The probability of small Gaussian ellipsoids and associated conditional moments. Annals of Probability, 21(1):14–24, January 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989391.

Mykland:1993:AEM

[Myk93] Per Aslak Mykland. Asymptotic expansions for martingales. Annals of Probability, 21(2):800–818, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989268.

Mytnik:1996:SRE

[Myt96] Leonid Mytnik. Superprocesses in random environments. Annals of Probability, 24(4):1953–1978, October 1996. CODEN APBYAE.

ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1041903212.

Mytnik:1998:WUH

[Myt98] Leonid Mytnik. Weak uniqueness for the heat equation with noise. Annals of Probability, 26(3):968–984, July 1998. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1022855740.

Nair:1990:RSC

[Nai90] M. Gopalan Nair. Random space change for multiparameter point processes. Annals of Probability, 18(3):1222–1231, July 1990. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990743.

Ney:1998:LDL

[NdA98] Peter Ney and A. de Acosta. Large deviation lower bounds for arbitrary additive functionals of a Markov chain. Annals of Probability, 26(4):1660–1682, October 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022855877.

Neuhauser:1990:ODS

[Neu90] Claudia Neuhauser. One dimensional stochastic Ising models with small migration. Annals of Probability, 18(4):1539–1546, October 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990630.

Noble:1992:EBS

[Nob92] Chris Noble. Equilibrium behavior of the sexual reproduction process with rapid diffusion. Annals of Probability, 20(2):724–745, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989802.

Nolan:1988:PPI

[Nol88] John P. Nolan. Path properties of index-β stable fields. Annals of Probability, 16(4):1596–1607, October 1988. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991586. See cor-

rection [Nol92].

Nolan:1992:CPP

[Nol92] John P. Nolan. Correction: Path properties of index-β stable fields. Annals of Probability, 20(3):1601–1602, July 1992. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989709. See

[Nol88].

Norris:1998:OUP

[Nor98] J. R. Norris. Ornstein–Uhlenbeck processes indexed by the circle. Annals of Probability, 26(2):465–478, April 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1022855640.

Nualart:1991:BVP

[NP91] D. Nualart and E. Pardoux. Boundary value problems for stochastic differential equations. Annals of Probability, 19(3):1118–1144, July 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990337.

Newman:1995:CDS

[NP95a] Charles M. Newman and Marcelo S. T. Piza. Correction: Divergence of shape fluctuations in two dimensions. Annals of Probability, 23(4):2057, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176987815. See [NP95b].

Newman:1995:DSF

[NP95b] Charles M. Newman and Marcelo S. T. Piza. Divergence of shape fluctuations in two dimensions. Annals of Probability, 23 (3):977–1005, July 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988171. See correction [NP95a].

Norvaisa:1994:SPS

[NS94] Rimas Norvaisa and Gennady Samorodnitsky. Stable processes with sample paths in Orlicz spaces. Annals of Probability, 22(4): 1904–1929, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988489.

Nualart:1992:RSP

[Nua92] David Nualart. Randomized stopping points and optimal stopping on the plane. Annals of Probability, 20(2):883–900, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989810.

Nummelin:1998:RBC

[Num98] Esa Nummelin. Review of “Topics in the Constructive Theory of Countable Markov Chains” by G. Fayolle, V. A. Malyshev and M. V. Menshikov. Annals of Probability, 26(1):425–427, January 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022855428.

Newman:1996:PSO

[NV96] Charles M. Newman and Sergio B. Volchan. Persistent survival of one-dimensional contact processes in random environments. Annals of Probability, 24(1):411–421, January 1996. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1042644723.

Newman:1997:PCP

[NW97] Charles M. Newman and C. Chris Wu. Percolation and contact processes with low-dimensional inhomogeneity. Annals of Probability, 25(4):1832–1845, October 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1023481113.

Nguyen:1993:TCO

[NY93] Bao Gia Nguyen and Wei-Shih Yang. Triangle condition for oriented percolation in high dimensions. Annals of Probability, 21(4): 1809–1844, October 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989001.

OBrien:1999:ULV

[O’B99] George L. O’Brien. Unusually large values for spectrally positive stable and related processes. Annals of Probability, 27 (2):990–1008, April 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022677393.

OConnell:1995:RTL

[O’C95] Neil O’Connell. Review: Torgny Lindvall, Lectures on the Coupling Method. Annals of Probability, 23(3):1456–1460, July 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988192.

Ocone:1990:TBM

[Oco90] Daniel L. Ocone. Two books on the Malliavin calculus. Annals of Probability, 18(1):431–437, January 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990957.

Oelschlager:1990:LTA

[Oel90] Karl Oelschlager. Limit theorems for age-structured populations. Annals of Probability, 18(1):290–318, January 1990. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990950.

ONeil:1993:ADW

[OR93] Kevin A. O’Neil and Richard A. Redner. Asymptotic distributions of weighted U-statistics of degree 2. Annals of Probability, 21 (2):1159–1169, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989286.

Orey:1991:MCS

[Ore91] Steven Orey. Markov chains with stochastically stationary transition probabilities. Annals of Probability, 19(3):907–928, July 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990328.

Overbeck:1995:AAF

[ORS95] Ludger Overbeck, Michael Rockner, and Byron Schmuland. An analytic approach to Fleming–Viot processes with interactive selection. Annals of Probability, 23(1):1–36, January 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988374.

Ornstein:1990:UAS

[OS90] Donald S. Ornstein and Paul C. Shields. Universal almost sure data compression. Annals of Probability, 18(2):441–452, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990840.

Okazaki:1994:DSR

[OS94] Yoshiaki Okazaki and Hiroshi Sato. Distinguishing a sequence of random variables from a random translate of itself. Annals of Probability, 22(2):1092–1096, April 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988742.

Osada:1999:GMR

[OS99] Hirofumi Osada and Herbert Spohn. Gibbs measures relative to Brownian motion. Annals of Probability, 27(3):1183–1207, July 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022677444.

Overbeck:1996:NS

[Ove96] L. Overbeck. Nonlinear superprocesses. Annals of Probability, 24(2):743–760, April 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1039639360.

Ornstein:1990:HSR

[OW90] Donald S. Ornstein and Benjamin Weiss. How sampling reveals a process. Annals of Probability, 18(3):905–930, July 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990729.

Peligrad:1998:BBE

[Pel98] Magda Peligrad. On the blockwise bootstrap for empirical processes for stationary sequences. Annals of Probability, 26(2):877– 901, April 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022855654.

Pemantle:1990:NUP

[Pem90] Robin Pemantle. Nonconvergence to unstable points in urn models and stochastic approximations. Annals of Probability, 18(2):698– 712, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990853.

Pemantle:1991:CST

[Pem91] Robin Pemantle. Choosing a spanning tree for the integer lattice uniformly. Annals of Probability, 19(4):1559–1574, October 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990223.

Pemantle:1992:AIM

[Pem92a] Robin Pemantle. Automorphism invariant measures on trees. Annals of Probability, 20(3):1549–1566, July 1992. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989706.

Pemantle:1992:CPT

[Pem92b] Robin Pemantle. The contact process on trees. Annals of Probability, 20(4):2089–2116, October 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989541.

Penrose:1992:SMS

[Pen92] Mathew D. Penrose. Semi-min-stable processes. Annals of Probability, 20(3):1450–1463, July 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989700.

Penrose:1996:RMS

[Pen96] Mathew D. Penrose. The random minimal spanning tree in high dimensions. Annals of Probability, 24(4):1903–1925, October 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1041903210.

Penrose:1999:SLL

[Pen99] Mathew D. Penrose. A Strong Law for the longest edge of the minimal spanning tree. Annals of Probability, 27(1):246–260, January 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022677261.

Perkins:1990:PSM

[Per90] Edwin Perkins. Polar sets and multiple points for super-Brownian motion. Annals of Probability, 18(2):453–491, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990841.

Peskir:1998:OSM

[Pes98] Goran Peskir. Optimal stopping of the maximum process: the maximality principle. Annals of Probability, 26(4):1614–1640, October 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022855875.

Papamarcou:1991:UCE

[PF91] Adrian Papamarcou and Terrence L. Fine. Unstable collectives and envelopes of probability measures. Annals of Probability, 19 (2):893–906, April 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990457.

Piau:1998:TCL

[Pia98] Didier Piau. Théoràme central limite fonctionnel pour une marche au hasard en environnement aléatoire. (French) [Functional central limit theorem for a random walk in a stochastic environment]. Annals of Probability, 26(3):1016–1040, July 1998. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1022855743.

Pinsky:1987:RTB

[Pin87] Ross G. Pinsky. Recurrence, transience and bounded harmonic functions for diffusions in the plane. Annals of Probability, 15 (3):954–984, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992075. See correction [Pin90b].

Pinsky:1990:ABS

[Pin90a] Ross G. Pinsky. The asymptotic behavior of the solution of the exterior Dirichlet problem for Brownian motion perturbed by a small parameter drift. Annals of Probability, 18(4):1602–1618, October 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990635.

Pinsky:1990:CRT

[Pin90b] Ross G. Pinsky. Correction: Recurrence, transience and bounded harmonic functions for diffusions in the plane. Annals of Probability, 18(1):438–440, January 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990958. See [Pin87].

Pinsky:1993:NAM

[Pin93] Ross G. Pinsky. A new approach to the Martin boundary via diffusions conditioned to hit a compact set. Annals of Probability, 21(1):453–481, January 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989411.

Pinelis:1994:OBD

[Pin94] Iosif Pinelis. Optimum bounds for the distributions of martingales in Banach spaces. Annals of Probability, 22(4):1679–1706, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.

aop/1176988477. See correction [Pin99].

Pinsky:1995:LTG

[Pin95] Ross G. Pinsky. On the large time growth rate of the support of supercritical super- Brownian motion. Annals of Probability, 23 (4):1748–1754, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176987801.

Pinsky:1996:TRL

[Pin96] Ross G. Pinsky. Transience, recurrence and local extinction properties of the support for supercritical finite measure-valued diffusions. Annals of Probability, 24(1):237–267, January 1996. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1042644715.

Pinelis:1999:CBB

[Pin99] Iosif Pinelis. Correction: “Optimum bounds for the distributions of martingales in Banach spaces” [Ann. Probab. 22 (1994), no. 4, 1679–1706; MR 96b:60010]. Annals of Probability, 27(4):2119, October 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022874833. See [Pin94].

Pittel:1993:SRP

[Pit93] Boris Pittel. The “Stable Roommates” problem with random preferences. Annals of Probability, 21(3):1441–1477, July 1993. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989126.

Piterbarg:1998:ECI

[Pit98] Vladimir V. Piterbarg. Expansions and contractions of isotropic stochastic flows of homeomorphisms. Annals of Probability, 26 (2):479–499, April 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022855641.

Pitman:1999:CMC

[Pit99a] Jim Pitman. Coalescents with multiple collisions. Annals of Probability, 27(4):1870–1902, October 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022874819.

Pitman:1999:SSL

[Pit99b] Jim Pitman. The SDE solved by local times of a Brownian excursion or bridge derived from the height profile of a random tree or forest. Annals of Probability, 27(1):261–283, January 1999. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1022677262.

Pollett:1992:RXQ

[Pol92] P. K. Pollett. Review: Xiang-Qun Yang, The Construction Theory of Denumerable Markov Processes. Annals of Probability, 20 (1):549–552, January 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989941.

Port:1990:AEE

[Por90] Sidney C. Port. Asymptotic expansions for the expected volume of a stable sausage. Annals of Probability, 18(2):492–523, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990842.

Porod:1996:CPR

[Por96] Ursula Porod. The cut-off phenomenon for random reflections. Annals of Probability, 24(1):74–96, January 1996. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1042644708.

Povel:1997:CLD

[Pov97] Tobias Povel. Critical large deviations of one-dimensional annealed Brownian motion in a Poissonian potential. Annals of Probability, 25(4):1735–1773, October 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1023481109.

Pardoux:1990:SVE

[PP90] Etienne Pardoux and Philip Protter. Stochastic Volterra equations with anticipating coefficients. Annals of Probability, 18(4): 1635–1655, October 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990638.

Pinsky:1993:TRC

[PP93] Mark Pinsky and Ross G. Pinsky. Transience/recurrence and Central Limit Theorem behavior for diffusions in random temporal environments. Annals of Probability, 21(1):433–452, January 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989410.

Pemantle:1994:DBT

[PP94] Robin Pemantle and Yuval Peres. Domination between trees and application to an explosion problem. Annals of Probability, 22 (1):180–194, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988855.

Pemantle:1995:CRW

[PP95a] Robin Pemantle and Yuval Peres. Critical random walk in random environment on trees. Annals of Probability, 23(1):105–140, January 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988379.

Pemantle:1995:GWT

[PP95b] Robin Pemantle and Yuval Peres. Galton–Watson trees with the same mean have the same polar sets. Annals of Probability, 23 (3):1102–1124, July 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988175.

Pisztora:1999:LDP

[PP99] Agoston Pisztora and Tobias Povel. Large deviation principle for random walk in a quenched random environment in the low speed regime. Annals of Probability, 27(3):1389–1413, July 1999. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1022677453.

Peddada:1991:PCM

[PR91] Shyamal Das Peddada and Donald St. P. Richards. Proof of a conjecture of M. L. Eaton on the characteristic function of the Wishart distribution. Annals of Probability, 19(2):868–874, April 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990455.

Peddada:1992:APP

[PR92] Shyamal D. Peddada and Donald St. P. Richards. Acknowledgment of priority: Proof of a conjecture of M. L. Eaton on the characteristic function of the Wishart distribution. Annals of Probability, 20(2):1107, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989822.

Pradeilles:1998:WPR

[Pra98] Frédéric Pradeilles. Wavefront propagation for reaction-diffusion systems and backward SDEs. Annals of Probability, 26(4): 1575–1613, October 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022855874.

Pritchard:1996:CSR

[Pri96] Geoffrey Pritchard. On the convergence of scaled random samples. Annals of Probability, 24(3):1490–1506, July 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1065725190.

Pruitt:1990:RER

[Pru90] William E. Pruitt. The rate of escape of random walk. Annals of Probability, 18(4):1417–1461, October 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990626.

Pemantle:1999:RPT

[PS99] Robin Pemantle and Jeffrey E. Steif. Robust phase transitions for Heisenberg and other models on general trees. Annals of Probability, 27(2):876–912, April 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022677389.

Pruitt:1996:PCI

[PT96] William E. Pruitt and S. James Taylor. Packing and covering indices for a general Lévy process. Annals of Probability, 24 (2):971–986, April 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1039639373.

Protter:1997:ESL

[PT97] Philip Protter and Denis Talay. The Euler scheme for Lévy driven stochastic differential equations. Annals of Probability, 25(1): 393–423, January 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1024404293.

Peligrad:1997:CLT

[PU97] Magda Peligrad and Sergey Utev. Central limit theorem for linear processes. Annals of Probability, 25(1):443–456, January 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1024404295.

Pemantle:1999:VRR

[PV99] Robin Pemantle and Stanislav Volkov. Vertex-reinforced random walk on Z has finite range. Annals of Probability, 27(3):1368–1388, July 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022677452.

Praestgaard:1993:EWB

[PW93] Jens Praestgaard and Jon A. Wellner. Exchangeably weighted bootstraps of the general empirical process. Annals of Probability, 21(4):2053–2086, October 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989011.

Picardello:1994:FMB

[PW94] Massimo A. Picardello and Wolfgang Woess. The full Martin boundary of the bi-tree. Annals of Probability, 22(4):2203–2222, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988500.

Perera:1998:COM

[PW98] Gonzalo Perera and Mario Wschebor. Crossings and occupation measures for a class of semimartingales. Annals of Probability, 26 (1):253–266, January 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022855418.

Pittel:1990:RTT

[PWM90] B. Pittel, W. A. Woyczynski, and J. A. Mann. Random tree-type partitions as a model for acyclic polymerization: Holtsmark (3/2stable) distribution of the supercritical gel. Annals of Probability, 18(1):319–341, January 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990951. See correction [PWM92].

Pittel:1992:CRT

[PWM92] B. Pittel, W. A. Woyczynski, and J. A. Mann. Correction: Random tree-type partitions as a model for acyclic polymerization: Holtsmark (3/2-stable) distribution of the supercritical gel. Annals of Probability, 20(2):1105–1106, April 1992. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989821. See

[PWM90].

Pitman:1997:TPP

[PY97] Jim Pitman and Marc Yor. The two-parameter Poisson–Dirichlet distribution derived from a stable subordinator. Annals of Probability, 25(2):855–900, April 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1024404422.

Pitman:1998:RBS

[PY98] Jim Pitman and Marc Yor. Random Brownian scaling identities and splicing of Bessel processes. Annals of Probability, 26(4): 1683–1702, October 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022855878.

Peszat:1995:SFP

[PZ95] Szymon Peszat and Jerzy Zabczyk. Strong Feller property and irreducibility for diffusions on Hilbert spaces. Annals of Probability, 23(1):157–172, January 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988381.

Qian:1996:CPF

[Qia96] Zhongmin Qian. On conservation of probability and the Feller property. Annals of Probability, 24(1):280–292, January 1996. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1042644717.

Qian:1998:VPS

[Qia98] Jinghua Qian. The p-variation of partial sum processes and the empirical process. Annals of Probability, 26(3):1370–1383, July 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022855756.

Quastel:1995:LDH

[Qua95] Jeremy Quastel. Large deviations from a hydrodynamic scaling limit for a nongradient system. Annals of Probability, 23 (2):724–742, April 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988286.

Rajput:1993:SCI

[Raj93] Balram S. Rajput. Supports of certain infinitely divisible probability measures on locally convex spaces. Annals of Probability, 21(2):886–897, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989272. See correction [Raj94].

Rajput:1994:CSC

[Raj94] Balram S. Rajput. Correction: Supports of certain infinitely divisible probability measures on locally convex spaces. Annals of Probability, 22(1):525, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988871. See [Raj93].

Resnick:1996:SOR

[RdH96] Sidney Resnick and Laurens de Haan. Second-order regular variation and rates of convergence in extreme-value theory. Annals of Probability, 24(1):97–124, January 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1042644709.

Reinert:1995:WLL

[Rei95] Gesine Reinert. A weak Law of Large Numbers for empirical measures via Stein’s method. Annals of Probability, 23(1):334–354, January 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988389.

Remillard:1994:CLI

[Rem94] Bruno Remillard. On Chung’s law of the iterated logarithm for some stochastic integrals. Annals of Probability, 22(4):1794–1802, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988483.

Renz:1996:NEC

[Ren96] Joachim Renz. A note on exact convergence rates in some martingale central limit theorems. Annals of Probability, 24(3): 1616–1637, July 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1065725195.

Revesz:1990:ELD

[Rev90] P. Revesz. Estimates of the largest disc covered by a random walk. Annals of Probability, 18(4):1784–1789, October 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990648.

Revesz:1993:CRW

[Rev93] P. Revesz. Clusters of a random walk on the plane. Annals of Probability, 21(1):318–328, January 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989406.

Rezakhanlou:1998:LDK

[Rez98] Fraydoun Rezakhanlou. Large deviations from a kinetic limit. Annals of Probability, 26(3):1259–1340, July 1998. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1022855753.

Robinson:1990:APS

[RHHQ90] J. Robinson, T. Hoglund, L. Holst, and M. P. Quine. On approximating probabilities for small and large deviations in R^{d}. Annals of Probability, 18(2):727–753, April 1990. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990856.

Rio:1993:SASa

[Rio93a] Emmanuel Rio. Strong approximation for set-indexed partial sum processes via KMT constructions I. Annals of Probability, 21 (2):759–790, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989266.

Rio:1993:SASb

[Rio93b] Emmanuel Rio. Strong approximation for set-indexed partial-sum processes, via KMT constructions II. Annals of Probability, 21 (3):1706–1727, July 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989138.

Rio:1995:FLI

[Rio95a] Emmanuel Rio. The functional law of the iterated logarithm for stationary strongly mixing sequences. Annals of Probability, 23 (3):1188–1203, July 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988179.

Rio:1995:MID

[Rio95b] Emmanuel Rio. A maximal inequality and dependent Marcinkiewicz– Zygmund Strong Laws. Annals of Probability, 23(2):918–937, April 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988295.

Roberts:1991:AAB

[Rob91] G. O. Roberts. Asymptotic approximations for Brownian motion boundary hitting times. Annals of Probability, 19(4):1689–1731, October 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990230.

Rockner:1992:PMB

[Roc92] Michael Rockner. On the parabolic Martin boundary of the Ornstein–Uhlenbeck operator on Wiener space. Annals of Probability, 20(2):1063–1085, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989818.

Rocha:1993:ISP

[Roc93] Amy L. Rocha. The infinite secretary problem with recall. Annals of Probability, 21(2):898–916, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989273.

Roichman:1996:RRW

[Roi96] Yuval Roichman. On random random walks. Annals of Probability, 24(2):1001–1011, April 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1039639375.

Rosen:1990:RWI

[Ros90a] Jay Rosen. Random walks and intersection local time. Annals of Probability, 18(3):959–977, July 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990731.

Rosinski:1990:SRI

[Ros90b] Jan Rosinski. On series representations of infinitely divisible random vectors. Annals of Probability, 18(1):405–430, January 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990956.

Rosen:1992:ASS

[Ros92a] Jay Rosen. The asymptotics of stable sausages in the plane. Annals of Probability, 20(1):29–60, January 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989917.

Rosen:1992:RLT

[Ros92b] Jay Rosen. Renormalization and limit theorems for selfintersections of superprocesses. Annals of Probability, 20(3): 1341–1368, July 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989694.

Rosenthal:1994:RRC

[Ros94] Jeffrey S. Rosenthal. Random rotations: Characters and random walks on SO(N). Annals of Probability, 22(1):398–423, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988864.

Rosinski:1995:SSS

[Ros95a] Jan Rosinski. On the structure of stationary stable processes. Annals of Probability, 23(3):1163–1187, July 1995. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988178.

Rosinski:1995:RSE

[Ros95b] Jan Rosinski. Remarks on strong exponential integrability of vector-valued random series and triangular arrays. Annals of Probability, 23(1):464–473, January 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988395.

Roy:1990:RSW

[Roy90] Rahul Roy. The Russo–Seymour–Welsh theorem and the equality of critical densities and the “dual” critical densities for continuum percolation on R^{2}. Annals of Probability, 18(4):1563–1575, October 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990632.

Rachev:1991:AID

[RR91] S. T. Rachev and L. Rüschendorf. Approximate independence of distributions on spheres and their stability properties. Annals of Probability, 19(3):1311–1337, July 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990346.

Rosinski:1993:DSF

[RS93] Jan Rosinski and Gennady Samorodnitsky. Distributions of subadditive functionals of sample paths of infinitely divisible processes. Annals of Probability, 21(2):996–1014, April 1993. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989279.

Rosinski:1996:SCI

[RS96] Jan Rosiński and Gennady Samorodnitsky. Symmetrization and concentration inequalities for multilinear forms with applications to zero-one laws for Lévy chaos. Annals of Probability, 24(1): 422–437, January 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1042644724.

Rheinlander:1997:PSS

[RS97] Thorsten Rheinländer and Martin Schweizer. On L^{2}-projections on a space of stochastic integrals. Annals of Probability, 25(4): 1810–1831, October 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1023481112.

Rogers:1991:LTS

[RW91] L. C. G. Rogers and J. B. Walsh. Local time and stochastic area integrals. Annals of Probability, 19(2):457–482, April 1991. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990435.

Rota:1997:SIC

[RW97] Gian-Carlo Rota and Timothy C. Wallstrom. Stochastic integrals: a combinatorial approach. Annals of Probability, 25(3):1257–1283, July 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1024404513.

Rosen:1991:TFR

[RY91] Jay Rosen and Marc Yor. Tanaka formulae and renormalization for triple intersections of Brownian motion in the plane. Annals of Probability, 19(1):142–159, January 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990538.

SanMartin:1993:ODS

[San93] Jaime San Martin. One-dimensional Stratonovich differential equations. Annals of Probability, 21(1):509–553, January 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989414.

Sato:1993:CUD

[Sat93] Ken-Iti Sato. Convolution of unimodal distributions can produce any number of modes. Annals of Probability, 21(3):1543–1549, July 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989129.

Samuel-Cahn:1992:DPI

[SC92] Ester Samuel-Cahn. A difference prophet inequality for bounded i.i.d. variables, with cost for observations. Annals of Probability, 20(3):1222–1228, July 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989689.

Schurger:1991:ASE

[Sch91] Klaus Schurger. Almost subadditive extensions of Kingman’s ergodic theorem. Annals of Probability, 19(4):1575–1586, October 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990224.

Schinazi:1992:BFE

[Sch92a] Rinaldo Schinazi. Brownian fluctuations of the edge for critical reversible nearest-particle systems. Annals of Probability, 20 (1):194–205, January 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989924.

Schonmann:1992:BSC

[Sch92b] Roberto H. Schonmann. On the behavior of some cellular automata related to bootstrap percolation. Annals of Probability, 20 (1):174–193, January 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989923.

Schneemeier:1993:WCG

[Sch93a] Wilhelm Schneemeier. Weak convergence and Glivenko–Cantelli results for weighted empirical U-processes. Annals of Probability, 21(2):1170–1184, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989287.

Schroeder:1993:PCL

[Sch93b] Carolyn Schroeder. I-projection and conditional limit theorems for discrete parameter Markov processes. Annals of Probability, 21(2):721–758, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989265.

Schweizer:1994:ARV

[Sch94] Martin Schweizer. Approximating random variables by stochastic integrals. Annals of Probability, 22(3):1536–1575, July 1994. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988611.

Schweizer:1996:APV

[Sch96] Martin Schweizer. Approximation pricing and the varianceoptimal martingale measure. Annals of Probability, 24(1):206–236, January 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1042644714.

Schied:1997:GAF

[Sch97] Alexander Schied. Geometric aspects of Fleming–Viot and Dawson–Watanabe processes. Annals of Probability, 25(3): 1160–1179, July 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1024404509.

Sepanski:1993:ASB

[Sep93] Steven J. Sepanski. Almost sure bootstrap of the mean under random normalization. Annals of Probability, 21(2):917–925, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989274.

Seppalainen:1994:LDM

[Sep94] Timo Seppalainen. Large deviations for Markov chains with random transitions. Annals of Probability, 22(2):713–748, April 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988727.

Seppalainen:1998:ETI

[Sep98a] Timo Seppäläinen. Entropy for translation-invariant randomcluster measures. Annals of Probability, 26(3):1139–1178, July 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022855747.

Seppalainen:1998:ELS

[Sep98b] Timo Seppäläinen. Exact limiting shape for a simplified model of first-passage percolation on the plane. Annals of Probability, 26 (3):1232–1250, July 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022855751.

Seppalainen:1999:EHT

[Sep99] Timo Seppäläinen. Existence of hydrodynamics for the totally asymmetric simple K-exclusion process. Annals of Probability, 27 (1):361–415, January 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022677266.

Shao:1995:MIP

[Sha95] Qi-Man Shao. Maximal inequalities for partial sums of ρ-mixing sequences. Annals of Probability, 23(2):948–965, April 1995. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988297.

Shao:1997:SNL

[Sha97] Qi-Man Shao. Self-normalized large deviations. Annals of Probability, 25(1):285–328, January 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1024404289.

Shafer:1998:RBM

[Sha98] Glenn Shafer. Review of “Creating Modern Probability: Its Mathematics, Physics, and Philosophy in Historical Perspective. Cambridge Studies in Probability, Induction, and Decision Theory” by Jan Von Plato. Annals of Probability, 26(1):416–424, January 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022855427.

Shepp:1991:API

[She91a] L. Shepp. Acknowledgment of priority: On the integral of the absolute value of the pinned Wiener process. Annals of Probability, 19(3):1397, July 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990351.

Sheu:1991:SET

[She91b] Shuenn-Jyi Sheu. Some estimates of the transition density of a nondegenerate diffusion Markov process. Annals of Probability, 19(2):538–561, April 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990440.

Sheu:1996:SEM

[She96] Yuan-Chung Sheu. On states of exit measures for superdiffusions. Annals of Probability, 24(1):268–279, January 1996. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1042644716.

Shih:1990:OSS

[Shi90] C. T. Shih. Ordered Skorokhod stopping for a sequence of measures. Annals of Probability, 18(4):1623–1634, October 1990. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990637.

Shieh:1992:MPS

[Shi92a] Narn-Rueih Shieh. Multiple points of sample paths of Markov processes. Annals of Probability, 20(2):553–562, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989790.

Shields:1992:EP

[Shi92b] Paul C. Shields. Entropy and prefixes. Annals of Probability, 20 (1):403–409, January 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989934.

Shields:1992:SME

[Shi92c] Paul C. Shields. String matching: The ergodic case. Annals of Probability, 20(3):1199–1203, July 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989686.

Shields:1997:SMB

[Shi97] Paul C. Shields. String matching bounds via coding. Annals of Probability, 25(1):329–336, January 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1024404290.

Shi:1998:WBM

[Shi98a] Zhan Shi. Windings of Brownian motion and random walks in the plane. Annals of Probability, 26(1):112–131, January 1998. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1022855413.

Shieh:1998:MPD

[Shi98b] Narn-Rueih Shieh. Multiple points of dilation-stable Lévy processes. Annals of Probability, 26(3):1341–1355, July 1998. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1022855754.

Shorack:1991:EFS

[Sho91] Galen R. Shorack. Embedding the finite sampling process at a rate. Annals of Probability, 19(2):826–842, April 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990453.

Silverstein:1990:WCR

[Sil90] Jack W. Silverstein. Weak convergence of random functions defined by the eigenvectors of sample covariance matrices. Annals of Probability, 18(3):1174–1194, July 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990741.

Simonelli:1995:LTC

[Sim95] Italo Simonelli. A limit theorem for a class of interacting particle systems. Annals of Probability, 23(1):141–156, January 1995. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988380.

Szczotka:1990:ASQ

[SK90] W. Szczotka and F. P. Kelly. Asymptotic stationarity of queues in series and the heavy traffic approximation. Annals of Probability, 18(3):1232–1248, July 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990744.

Skala:1993:EPM

[Ska93] Heinz J. Skala. The existence of probability measures with given marginals. Annals of Probability, 21(1):136–142, January 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989396.

Sudbury:1995:QOC

[SL95] Aidan Sudbury and Peter Lloyd. Quantum operators in classical probability theory: II. the concept of duality in interacting particle systems. Annals of Probability, 23(4):1816–1830, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176987804.

Slud:1993:MPP

[Slu93] Eric V. Slud. The moment problem for polynomial forms in normal random variables. Annals of Probability, 21(4):2200–2214, October 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989017.

Slud:1994:MRN

[Slu94] Eric V. Slud. MWI representation of the number of curve-crossings by a differentiable Gaussian process, with applications. Annals of Probability, 22(3):1355–1380, July 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988606.

Sowers:1992:LDR

[Sow92] Richard B. Sowers. Large deviations for a reaction-diffusion equation with non-Gaussian perturbations. Annals of Probability, 20 (1):504–537, January 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176989939.

Sowers:1994:MRD

[Sow94] R. B. Sowers. Multidimensional reaction-diffusion equations with white noise boundary perturbations. Annals of Probability, 22(4): 2071–2121, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988495.

Sowers:1995:ITE

[Sow95] Richard B. Sowers. Intermittency-type estimates for some nondegenerate SPDE’s. Annals of Probability, 23(4):1853–1874, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176987806.

Salzano:1997:SLE

[SS97] Marcia Salzano and Roberto H. Schonmann. The second lowest extremal invariant measure of the contact process. Annals of Probability, 25(4):1846–1871, October 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1023481114.

Salzano:1998:NPC

[SS98] Marcia Salzano and Roberto H. Schonmann. A new proof that for the contact process on homogeneous trees local survival implies complete convergence. Annals of Probability, 26(3):1251–1258, July 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022855752.

Salzano:1999:SLE

[SS99] Marcia Salzano and Roberto H. Schonmann. The second lowest extremal invariant measure of the contact process II. Annals of Probability, 27(2):845–875, April 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022677388.

Schechtman:1998:GMI

[SSZ98] G. Schechtman, Th. Schlumprecht, and J. Zinn. On the Gaussian measure of the intersection. Annals of Probability, 26(1): 346–357, January 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022855422.

Samorodnitsky:1991:PLS

[ST91] Gennady Samorodnitsky and Murad S. Taqqu. Probability laws with 1-stable marginals are 1-stable. Annals of Probability, 19(4): 1777–1780, October 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990235.

Samorodnitsky:1993:SMS

[ST93] Gennady Samorodnitsky and Murad S. Taqqu. Stochastic monotonicity and Slepian-type inequalities for infinitely divisible and stable random vectors. Annals of Probability, 21(1):143–160, January 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989397.

Samorodnitsky:1994:LMI

[ST94] Gennady Samorodnitsky and Murad S. Taqqu. Lévy measures of infinitely divisible random vectors and Slepian inequalities. Annals of Probability, 22(4):1930–1956, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988490.

Stafford:1990:PPC

[Sta90] Seth Stafford. A probabilistic proof of S.-Y. Cheng’s Liouville Theorem. Annals of Probability, 18(4):1816–1822, October 1990. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990651.

Stacey:1996:EIP

[Sta96] A. M. Stacey. The existence of an intermediate phase for the contact process on trees. Annals of Probability, 24(4):1711–1726, October 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1041903203.

Steif:1991:STB

[Ste91a] Jeffrey E. Steif. Space-time Bernoullicity of the lower and upper stationary processes for attractive spin systems. Annals of Probability, 19(2):609–635, April 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990444.

Steinebach:1991:SLS

[Ste91b] Josef Steinebach. Strong laws for small increments of renewal processes. Annals of Probability, 19(4):1768–1776, October 1991. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990234.

Steif:1994:TVA

[Ste94] Jeffrey E. Steif. The threshold voter automaton at a critical point. Annals of Probability, 22(3):1121–1139, July 1994. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988597.

Steinsaltz:1999:LCI

[Ste99] David Steinsaltz. Locally contractive iterated function systems. Annals of Probability, 27(4):1952–1979, October 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1022874823.

Stockbridge:1990:TACb

[Sto90a] Richard H. Stockbridge. Time-average control of martingale problems: A linear programming formulation. Annals of Probability, 18 (1):206–217, January 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990945.

Stockbridge:1990:TACa

[Sto90b] Richard H. Stockbridge. Time-average control of martingale problems: Existence of a stationary solution. Annals of Probability, 18 (1):190–205, January 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990944.

Stong:1995:ERW

[Sto95a] Richard Stong. Eigenvalues of random walks on groups. Annals of Probability, 23(4):1961–1981, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176987811.

Stong:1995:ENR

[Sto95b] Richard Stong. Eigenvalues of the natural random walk on the Burnside group B(3, n). Annals of Probability, 23(4):1950–1960, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176987810.

Stong:1995:RWG

[Sto95c] Richard Stong. Random walks on the groups of upper triangular matrices. Annals of Probability, 23(4):1939–1949, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176987809.

Stricker:1990:VPP

[Str90] C. Stricker. Valeurs prises par les martingales locales positives continues á un instant donné. (French) [Values taken by local positive continuous martingales at a given instant]. Annals of Probability, 18(2):626–629, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990848.

Stute:1991:CS

[Stu91] Winfried Stute. Conditional U-statistics. Annals of Probability, 19(2):812–825, April 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990452.

Stute:1994:SPM

[Stu94] Winfried Stute. U-statistic processes: A martingale approach. Annals of Probability, 22(4):1725–1744, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988480.

Sturm:1998:DPH

[Stu98] K. T. Sturm. Diffusion processes and heat kernels on metric spaces. Annals of Probability, 26(1):1–55, January 1998. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1022855410.

Sudbury:1990:BAP

[Sud90] Aidan Sudbury. The branching annihilating process: An interacting particle system. Annals of Probability, 18(2):581–601, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990846.

Sun:1993:TPM

[Sun93] Jiayang Sun. Tail probabilities of the maxima of Gaussian random fields. Annals of Probability, 21(1):34–71, January 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989393.

Steif:1999:ENF

[SvdB99] J. E. Steif and J. van den Berg. On the existence and nonexistence of finitary codings for a class of random fields. Annals of Probability, 27(3):1501–1522, July 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022677456.

Sheehy:1992:UDC

[SW92] Anne Sheehy and Jon A. Wellner. Uniform Donsker classes of functions. Annals of Probability, 20(4):1983–2030, October 1992. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989538.

Sethuraman:1996:CLT

[SX96] Sunder Sethuraman and Lin Xu. A central limit theorem for reversible exclusion and zero-range particle systems. Annals of Probability, 24(4):1842–1870, October 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1041903208.

Shor:1991:MGM

[SY91] P. W. Shor and J. E. Yukich. Minimax Grid matching and empirical measures. Annals of Probability, 19(3):1338–1348, July 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990347.

Shao:1996:WCW

[SY96] Qi-Man Shao and Hao Yu. Weak convergence for weighted empirical processes of dependent sequences. Annals of Probability, 24(4): 2098–2127, October 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1041903220.

Shepp:1992:NCE

[SZ92] Larry A. Shepp and Ofer Zeitouni. A note on conditional exponential moments and Onsager–Machlup functionals. Annals of Probability, 20(2):652–654, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989796.

Sznitman:1999:LLN

[SZ99] Alain-Sol Sznitman and Martin Zerner. A Law of Large Numbers for random walks in random environment. Annals of Probability, 27(4):1851–1869, October 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1022874818.

Szczotka:1992:DFL

[Szc92] Wladyslaw Szczotka. A distributional form of Little’s law in heavy traffic. Annals of Probability, 20(2):790–800, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989806.

Sznitman:1993:BSA

[Szn93] Alain-Sol Sznitman. Brownian survival among Gibbsian traps. Annals of Probability, 21(1):490–508, January 1993. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989413.

Sznitman:1995:CVR

[Szn95] Alain-Sol Sznitman. Crossing velocities and random lattice animals. Annals of Probability, 23(3):1006–1023, July 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988172.

Sznitman:1996:DFL

[Szn96] Alain-Sol Sznitman. Distance fluctuations and Lyapounov exponents. Annals of Probability, 24(3):1507–1530, July 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1065725191.

Sznitman:1997:CPE

[Szn97] Alain-Sol Sznitman. Capacity and principal eigenvalues: the method of enlargement of obstacles revisited. Annals of Probability, 25(3):1180–1209, July 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1024404510.

Szulga:1990:NHS

[Szu90] Jerzy Szulga. A note on hypercontractivity of stable random variables. Annals of Probability, 18(4):1746–1758, October 1990. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176990645.

Szulga:1991:MSI

[Szu91] Jerzy Szulga. Multiple stochastic integrals with respect to symmetric infinitely divisible random measures. Annals of Probability, 19(3):1145–1156, July 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990338.

Takanobu:1997:EEI

[Tak97] Satoshi Takanobu. On the error estimate of the integral kernel for the Trotter product formula for Schrödinger operators. Annals of Probability, 25(4):1895–1952, October 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1023481116.

Talagrand:1990:CAS

[Tal90a] Michel Talagrand. Characterization of almost surely continuous 1-stable random Fourier series and strongly stationary processes. Annals of Probability, 18(1):85–91, January 1990. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990939.

Talagrand:1990:SBS

[Tal90b] Michel Talagrand. Sample boundedness of stochastic processes under increment conditions. Annals of Probability, 18(1):1–49, January 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990936.

Talagrand:1992:NSC

[Tal92] M. Talagrand. Necessary and sufficient conditions for sample continuity of random Fourier series and of harmonic infinitely divisible processes. Annals of Probability, 20(1):1–28, January 1992. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989916.

Talagrand:1993:RID

[Tal93] Michel Talagrand. Regularity of infinitely divisible processes. Annals of Probability, 21(1):362–432, January 1993. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989409.

Talagrand:1994:SBG

[Tal94a] M. Talagrand. Sharper bounds for Gaussian and empirical processes. Annals of Probability, 22(1):28–76, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988847.

Talagrand:1994:TCU

[Tal94b] M. Talagrand. The transportation cost from the uniform measure to the empirical measure in dimension ≥ 3. Annals of Probability, 22(2):919–959, April 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988735.

Talagrand:1994:RAZ

[Tal94c] Michel Talagrand. On Russo’s approximate zero-one law. Annals of Probability, 22(3):1576–1587, July 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988612.

Talagrand:1994:SBP

[Tal94d] Michel Talagrand. The small ball problem for the Brownian sheet. Annals of Probability, 22(3):1331–1354, July 1994. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988605.

Talagrand:1995:BRF

[Tal95a] Michel Talagrand. A borderline random Fourier series. Annals of Probability, 23(2):776–785, April 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988289.

Talagrand:1995:HMT

[Tal95b] Michel Talagrand. Hausdorff measure of trajectories of multiparameter fractional Brownian motion. Annals of Probability, 23 (2):767–775, April 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176988288.

Talagrand:1996:MMG

[Tal96a] Michel Talagrand. Majorizing measures: the generic chaining. Annals of Probability, 24(3):1049–1103, July 1996. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1065725175.

Talagrand:1996:NLI

[Tal96b] Michel Talagrand. A new look at independence. Annals of Probability, 24(1):1–34, January 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1042644705.

Tanemura:1993:CLT

[Tan93] Hideki Tanemura. Central Limit Theorem for a random walk with random obstacles in R^{d}. Annals of Probability, 21(2):936–960, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989276.

Thomassen:1992:IIT

[Tho92] Carsten Thomassen. Isoperimetric inequalities and transient random walks on graphs. Annals of Probability, 20(3):1592–1600, July 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989708.

Thorisson:1996:TRE

[Tho96] Hermann Thorisson. Transforming random elements and shifting random fields. Annals of Probability, 24(4):2057–2064, October 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1041903217.

Toth:1995:TSA

[Tot95] Balint Toth. The “True” self-avoiding walk with bond repulsion on Z: Limit theorems. Annals of Probability, 23(4):1523–1556, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176987793.

Toth:1996:GRK

[Tót96] Bálint Tóth. Generalized Ray–Knight theory and limit theorems for self-interacting random walks on Z^{1}. Annals of Probability, 24 (3):1324–1367, July 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1065725184.

Touati:1990:LFD

[Tou90] Abderrahmen Touati. Loi fonctionnelle du logarithme itéré pour les processus de Markov recurrents. (French) [Functional law of the iterated logarithm for recurrent Markov processes]. Annals of Probability, 18(1):140–159, January 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990942.

Towghi:1995:SIP

[Tow95] Nasser Towghi. Stochastic integration of processes with finite generalized variations. I. Annals of Probability, 23(2):629–667, April 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988282.

Tribe:1992:BSN

[Tri92] Roger Tribe. The behavior of superprocesses near extinction. Annals of Probability, 20(1):286–311, January 1992. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176989927.

Tuero:1993:SCR

[Tue93] Araceli Tuero. On the stochastic convergence of representations based on Wasserstein metrics. Annals of Probability, 21(1):72–85, January 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989394.

Turkman:1990:SRP

[TW90] K. F. Turkman and A. M. Walker. A stability result for the periodogram. Annals of Probability, 18(4):1765–1783, October 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990647.

Takeda:1997:APA

[TZ97] Masayoshi Takeda and Tusheng Zhang. Asymptotic properties of additive functionals of Brownian motion. Annals of Probability, 25(2):940–952, April 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1024404425.

Uchiyama:1998:WTR

[Uch98] Kôhei Uchiyama. Wiener’s test for random walks with mean zero and finite variance. Annals of Probability, 26(1):368–376, January 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022855424.

Vuolle-Apiala:1994:IET

[VA94] J. Vuolle-Apiala. Itô excursion theory for self-similar Markov processes. Annals of Probability, 22(2):546–565, April 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176988721.

VanZandt:1993:HMF

[Van93] Timothy Van Zandt. The Hausdorff metric of σ-fields and the value of information. Annals of Probability, 21(1):161–167, January 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989398.

Vanderbei:1995:PFC

[Van95] Robert J. Vanderbei. A probabilistic formula for the concave Hull of a function. Annals of Probability, 23(4):2014–2021, October 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176987813.

vandenBerg:1997:CMC

[vdB97] J. van den Berg. A constructive mixing condition for 2-D Gibbs measures with random interactions. Annals of Probability, 25 (3):1316–1333, July 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1024404515.

vanderVaart:1996:NDC

[vdV96] Aad van der Vaart. New Donsker classes. Annals of Probability, 24 (4):2128–2140, October 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1041903221.

Verzani:1995:SPS

[Ver95] John Verzani. Slow points in the support of historical Brownian motion. Annals of Probability, 23(1):56–70, January 1995. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988376.

Verzani:1996:CHP

[Ver96] John Verzani. On the convex hull of planar Brownian snake. Annals of Probability, 24(3):1280–1299, July 1996. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1065725182.

Veretennikov:1999:LDA

[Ver99] A. Yu. Veretennikov. On large deviations in the averaging principle for SDEs with a “full dependence”. Annals of Probability, 27 (1):284–296, January 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022677263.

Vitale:1996:WFG

[Vit96] Richard A. Vitale. The Wills functional and Gaussian processes. Annals of Probability, 24(4):2172–2178, October 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1041903224.

VanDenBerg:1994:DPS

[VM94] J. Van Den Berg and C. Maes. Disagreement percolation in the study of Markov fields. Annals of Probability, 22(2):749–763, April 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988728.

Voit:1997:RCG

[Voi97] Michael Voit. Rate of convergence to Gaussian measures on nspheres and Jacobi hypergroups. Annals of Probability, 25(1): 457–477, January 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1024404296.

Wang:1991:SIC

[Wan91a] Gang Wang. Sharp inequalities for the conditional square function of a martingale. Annals of Probability, 19(4):1679–1688, October 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990229.

Wang:1991:CPC

[Wan91b] Y. H. Wang. A compound Poisson convergence theorem. Annals of Probability, 19(1):452–455, January 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990555.

Wang:1995:DSS

[Wan95a] Gang Wang. Differential subordination and strong differential subordination for continuous-time martingales and related sharp inequalities. Annals of Probability, 23(2):522–551, April 1995. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988278.

Wang:1995:ABL

[Wan95b] Jia-Gang Wang. The asymptotic behavior of locally square integrable martingales. Annals of Probability, 23(2):552–585, April 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988279.

Wang:1997:SEL

[Wan97] Feng-Yu Wang. Sharp explicit lower bounds of heat kernels. Annals of Probability, 25(4):1995–2006, October 1997. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1023481118.

Wang:1999:HIL

[Wan99] Feng-Yu Wang. Harnack inequalities for log-Sobolev functions and estimates of log-Sobolev constants. Annals of Probability, 27(2):653–663, April 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022677381.

Watkins:1996:RBR

[Wat96] Joseph C. Watkins. Review of “Lectures on Random Evolutions,” by Mark A. Pinsky. Annals of Probability, 24(3):1647–1652, July 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1065725198.

Weintraub:1991:SEP

[Wei91] Keith Steven Weintraub. Sample and ergodic properties of some min-stable processes. Annals of Probability, 19(2):706–723, April 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990447.

Wen:1990:RED

[Wen90] Liu Wen. Relative entropy densities and a class of limit theorems of the sequence of m-valued random variables. Annals of Probability, 18(2):829–839, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176990860.

Wick:1990:DSR

[Wic90] W. David Wick. Diffusion in a singular random environment. Annals of Probability, 18(1):50–67, January 1990. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990937.

Wihstutz:1997:BRB

[Wih97] Volker Wihstutz. Book review of “Markov Chains and Stochastic Stability” by S. P. Meyn and R. L. Tweedie. Annals of Probability, 25(3):1536–1540, July 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1024404525.

Winkler:1990:DIF

[Win90] William E. Winkler. On Dykstra’s iterative fitting procedure. Annals of Probability, 18(3):1410–1415, July 1990. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176990752.

Woodroofe:1990:NRT

[Woo90] Michael Woodroofe. On the nonlinear renewal theorem. Annals of Probability, 18(4):1790–1805, October 1990. CODEN APBYAE.

ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990649.

Wood:1992:BMF

[Woo92] G. R. Wood. Binomial mixtures and finite exchangeability. Annals of Probability, 20(3):1167–1173, July 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989684.

Wu:1994:LDM

[Wu94a] Liming Wu. Large deviations, moderate deviations and LIL for empirical processes. Annals of Probability, 22(1):17–27, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988846.

Wu:1994:ABT

[Wu94b] Yadong Wu. Asymptotic behavior of the two level measure branching process. Annals of Probability, 22(2):854–874, April 1994. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176988733.

Wu:1995:MDD

[Wu95] Liming Wu. Moderate deviations of dependent random variables related to CLT. Annals of Probability, 23(1):420–445, January 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.

aop/1176988393. See correction [Wu99].

Wu:1999:CBD

[Wu99] L. Wu. Correction: “Moderate deviations of dependent random variables related to the CLT” [Ann. Probab. 23 (1995), no. 1, 420– 445; MR 96f:60047]. Annals of Probability, 27(4):2122–2123, October 1999. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022874832. See [Wu95].

Wuthrich:1998:SBT

[Wüt98] Mario V. Wüthrich. Superdiffusive behavior of two-dimensional Brownian motion in a Poissonian potential. Annals of Probability, 26(3):1000–1015, July 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1022855742.

Wehr:1998:AGF

[WW98] Jan Wehr and Jung Woo. Absence of geodesics in first-passage percolation on a half-plane. Annals of Probability, 26(1):358–367, January 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1022855423.

Xia:1994:WCM

[Xia94a] Aihua Xia. Weak convergence of Markov processes with extended generators. Annals of Probability, 22(4):2183–2202, October 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988499.

Xiang:1994:LLK

[Xia94b] Xiaojing Xiang. A law of the logarithm for kernel quantile density estimators. Annals of Probability, 22(2):1078–1091, April 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988741.

Xiao:1995:DRG

[Xia95] Yimin Xiao. Dimension results for Gaussian vector fields and Index-α stable fields. Annals of Probability, 23(1):273–291, January 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988387.

Yaguchi:1990:EAN

[Yag90] Hirotake Yaguchi. Entropy analysis of a nearest-neighbor attractive/repulsive exclusion process on one-dimensional lattices. Annals of Probability, 18(2):556–580, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990845.

Yaguchi:1991:APE

[Yag91] Hirotake Yaguchi. Acknowledgment of priority: Entropy analysis of a nearest-neighbor attractive/repulsive exclusion process on one-dimensional lattices. Annals of Probability, 19(4):1822, October 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990238.

Yu:1994:RCE

[Yu94] Bin Yu. Rates of convergence for empirical processes of stationary mixing sequences. Annals of Probability, 22(1):94–116, January 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988849.

Yu:1996:SIP

[Yu96] Hao Yu. A strong invariance principle for associated sequences. Annals of Probability, 24(4):2079–2097, October 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1041903219.

Zabell:1992:NTC

[Zab92] S. L. Zabell. A note on translation continuity of probability measures. Annals of Probability, 20(1):410–420, January 1992. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989935.

Zerner:1998:LEQ

[Zer98] Martin P. W. Zerner. Lyapounov exponents and quenched large deviations for multidimensional random walk in random environment. Annals of Probability, 26(4):1446–1476, October 1998. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1022855870.

Zhang:1993:STE

[Zha93] Yu Zhang. A shape theorem for epidemics and forest fires with finite range interactions. Annals of Probability, 21(4):1755–1781, October 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988999.

Zhang:1994:NIP

[Zha94] Yu Zhang. A note on inhomogeneous percolation. Annals of Probability, 22(2):803–819, April 1994. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176988730.

Zhang:1996:SLL

[Zha96a] Cun-Hui Zhang. Strong Law of Large Numbers for sums of products. Annals of Probability, 24(3):1589–1615, July 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1065725194.

Zhang:1996:CCT

[Zha96b] Yu Zhang. The complete convergence theorem of the contact process on trees. Annals of Probability, 24(3):1408–1443, July 1996. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1065725187.

Zhang:1997:SAT

[Zha97] Li-Xin Zhang. Strong approximation theorems for geometrically weighted random series and their applications. Annals of Probability, 25(4):1621–1635, October 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1023481105.

Zhang:1999:SBF

[Zha99] Cun-Hui Zhang. Sub-Bernoulli functions, moment inequalities and Strong Laws for nonnegative and symmetrized U-statistics. Annals of Probability, 27(1):432–453, January 1999. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1022677268.

Zheng:1995:CPC

[Zhe95] Weian Zheng. Conditional propagation of chaos and a class of quasilinear PDE’s. Annals of Probability, 23(3):1389–1413, July 1995. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176988189.

Zakai:1992:WDR

[ZZ92] M. Zakai and O. Zeitouni. When does the Ramer formula look like the Girsanov formula? Annals of Probability, 20(3):1436–1440, July 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176989698.